### v3

# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

Abstract
We compute the twist-2 matching of transverse momentum dependent (TMD) helicity parton distribution and fragmentation functions at next-to-next-to-next-to-leading order (N3LO) in QCD. This calculation entails the complete set of next-to-next-to-leading order (NNLO) Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) splitting functions govering the evolution of helicity-dependent parton distribution functions (PDFs) and fragmentation functions (FFs). Within TMD factorization framework, we quantify the impact of radiative corrections by completing the next-to-next-to-next-to-leading logarithmic (N3LL) prediction for lepton-hadron transverse momentum imbalance in semi-inclusive deep inelastic scattering (SIDIS). Our results provide the most precise theoretical input for probing the helicity structure and confined motion of quarks and gluons at future electron-ion collider (EIC).
preprint: MPP-2025-159
1 Introduction
Transverse-Momentum-Dependent (TMD) Parton Distribution Functions (PDFs) parameterize the confined motion of parton constituents inside a nucleon target. TMD Fragmentation Functions (FFs) describe the hadronization process of parton decay into color-singlet hadrons. On top of that, TMDs also carry novel information about the nucleon’s internal flavor and spin structures, small-x TMDs encodes Regge asymptotics of QCD governed by Balitsky-Fadin-Kuraev-Lipatov (BFKL) evolution Balitsky:1978ic ; Kuraev:1977fs and beyond. TMDs are key objects in the next QCD frontier, and are essential for understanding the glue that binds us all Accardi:2012qut .
A longstanding objective in QCD is to uncover the fundamental mechanisms of color confinement. On the other hand, QCD is known for the converse of confinement– the celebrated property of asymptotic freedom Gross:1973id ; Politzer:1973fx . Thanks to the hierarchy between short- and long-distance scales, the dynamics at different regimes can be systematically disentangled within the framework of effective field theory. Schematically, the cross section admits a factorized structure
$\sigma(Q,\Lambda_{\text{QCD}})=\mathcal{C}(Q,\mu)\otimes\langle\mathcal{O}\rangle(\mu,\Lambda_{\text{QCD}})+\mathcal{O}\left(\frac{\Lambda^{2}_{\text{QCD}}}{Q^{2}}\right)\,$, (1)
where $\mu$ is the factorization scale. Factorization based on Collins-Soper-Sterman (CSS) formalism Collins:1988ig has been proven for Drell-Yan process and single inclusive hadron production Bodwin:1984hc ; Collins:1984kg ; Collins:1988ig ; Collins:1989gx ; Nayak:2005rt .
The principle of factorization underlies all theoretical predictions of QCD. To the UV, the renormalization of $\alpha_{s}$ is independent of the external momenta flowing into the Green functions, so that UV physics is completely factorized from collider measurements performed on asymptotic final states. To the IR, long-distance physics is parametrized while short-distance dynamics are computed from Feynman amplitudes, again thanks to the asymptotic freedom. UV renormalization and IR factorization scheme ambiguities drop out in the relations among physical quantities.
In the context of TMDs, factorization provides the gateway that bridges partonic high-energy interactions to what is observed at the detectors: the confined motion can never been probed directly, it manifests through correlations with the distribution of final-state gauge bosons, leptons, and hadrons, owing to the unification of Standard Model gauge group ${\text{SU}}(3)\times{\text{SU}}(2)\times{\text{U}}(1)_{\text{Y}}~({\text{SU}}(2)\times{\text{U}}(1)_{\text{Y}}\to\text{U}(1)_{\text{EM}}$. Indeed, TMDs are indispensable non-perturbative input for a variety of benchmark observables in the Standard Model, including Drell-Yan process Dokshitzer:1978yd ; Parisi:1979se ; Collins:1984kg ; Arnold:1990yk ; Ladinsky:1993zn ; Bozzi:2010xn ; Becher:2010tm ; Becher:2011xn ; Bertone:2019nxa , semi-inclusive deep-inelastic scattering (SIDIS) Ji:2004wu ; Ji:2004xq ; Liu:2018trl ; Fang:2024auf ; Gao:2022bzi ; Ebert:2021jhy ; Fang:2024auf , electron-positron annihilation to hadrons and jets Collins:1981uk ; Collins:1981va ; Neill:2016vbi ; Gutierrez-Reyes:2018qez ; Gutierrez-Reyes:2019vbx ; Gutierrez-Reyes:2019msa , Higgs boson production Berger:2002ut ; Bozzi:2005wk ; Gao:2005iu ; Echevarria:2015uaa ; Neill:2015roa ; Bizon:2017rah ; Chen:2018pzu ; Bizon:2018foh , top quark pair production Zhu:2012ts ; Li:2013mia ; Catani:2014qha ; Catani:2018mei , $J/\psi$ production Echevarria:2024idp ; Copeland:2023wbu ; Kishore:2018ugo ; Maxia:2024cjh ; Banu:2024ywv ; Kishore:2024bdd ; Chakrabarti:2022rjr ; Boer:2023zit ; Bor:2022fga ; Bacchetta:2018ivt ; Boer:2016bfj , as well as Energy-Energy (EEC) or Charge-Charge Correlator Moult:2018jzp ; Gao:2019ojf ; Gao:2024wcg ; Kang:2024otf ; Kang:2023big ; Liu:2024kqt ; Monni:2025zyv at lepton and hadron colliders.
Helicity TMDs in particular probe spin-momentum correlations and are central to understanding the decomposition of nucleon spin. SIDIS provides direct access to such distributions through transverse momentum and spin asymmetry measurements. In this paper, we consider lepton-jet transverse momentum imbalance in lepton-hadron colliders with polarized high-energy lepton beams scattering off longitudinally polarized nucleon targets:
$\ell(p_{\ell},\lambda_{\ell})+N(P_{N},S_{\parallel})\to\ell^{\prime}(p_{\ell^{\prime}},\lambda_{\ell^{\prime}})+h(P_{h})+X(P_{X})\,.$ (2)
We derive the TMD factorization formula for this observable within the framework of Soft-Collinear Effective Theory (SCET) Bauer:2000ew ; Bauer:2000yr ; Bauer:2001ct ; Bauer:2001yt , which enables a systematic separation of perturbative and nonperturbative contributions in the small transverse momentum regime. Using our new analytic expressions, we obtain a high-precision prediction for the $q_{\ast}$ spectrum, which provides a complementary probe to the TMD dynamics, alongside the conventional $q_{T}$-differential cross section in SIDIS.
This work is motivated by the broader program of global analyses of TMD distributions. Over the past decade, global fits of unpolarized and polarized TMDs have been performed using SIDIS, Drell-Yan, and $e^{+}e^{-}$ annihilation data within the TMD factorization framework Sun:2014dqm ; Bertone:2019nxa ; Scimemi:2019cmh ; Moos:2023yfa ; Moos:2025sal ; Almasy:2011eq ; Bacchetta:2017gcc ; Bacchetta:2019sam ; Bacchetta:2022awv ; Bacchetta:2024qre ; Bacchetta:2025ara . These fits have reached approximately N4LL accuracy in the unpolarized sector, while efforts to include spin-dependent TMDs–such as helicity and transversity distributions–are currently underway Yang:2024drd ; Bacchetta:2024yzl . Despite this progress, theoretical control over helicity-sensitive observables still lags behind that of their unpolarized counterparts. In particular, the full next-to-next-to-next-to-leading order (N3LO) matching for helicity TMDs has so far been unavailable. This limits the precision of phenomenological fits, especially in light of the high-statistics spin-polarized SIDIS measurements anticipated from the Electron-Ion Collider (EIC).
The study of helicity TMDs is inherently linked to the evolution of helicity PDFs and FFs, which is governed by polarized DGLAP Altarelli:1977zs ; Gribov:1972ri ; Dokshitzer:1977sg splitting functions. Unlike in the unpolarized case, where splitting functions are known up to NNLO Moch:2004pa ; Vogt:2004mw ; Mitov:2006ic ; Moch:2007tx ; Chen:2020uvt and partially known to N3LO Moch:2017uml ; Gehrmann:2023ksf ; Gehrmann:2023cqm ; Gehrmann:2023iah ; Falcioni:2023luc ; Falcioni:2023vqq ; Falcioni:2023tzp ; Moch:2023tdj ; Basdew-Sharma:2022vya ; Falcioni:2024xyt , their polarized counterparts have long remained incomplete beyond NLO Mertig:1995ny ; Vogelsang:1995vh ; Vogelsang:1996im ; Zijlstra:1993sh ; Behring:2019tus ; Stratmann:1996hn ; Rijken:1997rg . In recent years, significant effort has been devoted to computing the NNLO helicity splitting functions Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , driven by the need for high-precision QCD evolution in global analyses and improved theoretical control over spin asymmetries in polarized deep-inelastic scattering. These splitting functions are also essential for implementing helicity-sensitive parton evolution in Monte Carlo simulations of polarized scattering processes. Moreover, they play a key role in the ongoing effort to resolve the longstanding proton spin puzzle–the question of how the proton’s spin is distributed among its constituent quarks and gluons. Helicity PDFs provide a direct link between measurable observables and the partonic spin content of the nucleon, and their precise extraction requires accurate knowledge of both their functional forms and scale evolution. In this context, higher-order polarized splitting functions and matching coefficients are indispensable for reducing theoretical uncertainties in spin decomposition. In particular, recent progress in understanding helicity evolution at small $x$ has revealed potentially non-negligible contributions to the proton spin from the low-$x$ region Kovchegov:2015pbl ; Kovchegov:2016weo ; Kovchegov:2017jxc ; Kovchegov:2017lsr . Beyond their phenomenological relevance, higher-order splitting functions also offer a complementary theoretical motivation: they provide the database for investigating crossing symmetry and beyond: the generalization of Gribov-Lipatov reciprocity beyond amplitude-level crossing Chen:2020uvt .
In this paper, we take a significant step toward completing the perturbative framework for helicity-dependent observables in QCD. We compute the twist-2 matching coefficients for helicity TMD PDFs and FFs at N3LO accuracy. As an outcome, we derive the full set of NNLO polarized DGLAP splitting functions in analytic form. These results establish a new level of precision for helicity evolution and TMD matching, placing helicity observables on equal theoretical footing with their unpolarized counterparts. We further illustrate the numerical impact of higher-order corrections by performing a N3LL-resummed calculation of the transverse momentum spectrum in SIDIS. Together, these developments provide a solid foundation for future global analyses of spin-dependent data and for interpreting precision measurements at the EIC.
2 The TMD Factorization of $q_{\ast}$ Spectrum
We define leptonic momentum transfer as $q=p_{l}-p_{\ell^{\prime}}$, with virtuality $Q^{2}=-q^{2}$. We also introduce the standard set of kinematic invariants
$x=\frac{Q^{2}}{2P_{N}\cdot q}\,,\quad y=\frac{P_{N}\cdot q}{P_{N}\cdot p_{\ell}}\,,\quad z=\frac{P_{N}\cdot P_{h}}{P_{N}\cdot q}\,.$ (3)
In conventional SIDIS analyses, one measures the transverse-momentum spectrum $q_{T}$ of the virtual photon in hadron-hadron frame, or equivalently the $P^{T}_{h}$ of the identified hadron in photon-hadron frame where $\vec{P}^{\perp}_{h}=-z\vec{q}_{\perp}$. Instead of the conventional $q_{T}$ spectrum, we propose to measure the angular correlations in the transverse plane, defined as follows
$\frac{\mathrm{d}\sigma^{h}}{\mathrm{d}\cos\phi}=$ $\int\mathrm{d}\sigma_{\ell+N\to\ell^{\prime}+h+X}\times\delta(\cos\phi_{\ell^{\prime}h}-\cos\phi)\,$, (4)
where $\phi_{\ell^{\prime}h}$ is the lepton-hadron azimuthal angle in the lepton-target frame. The observable select events with a fixed azimuthal angle $\cos\phi_{\ell^{\prime}h}=\cos\phi$ . An equivalent but Lorentz-invariant way of parametrization is to select events with fixed momentum projection $q_{\perp}\cdot v$, where $v$ is the normal 4-vector to the event plane
$v_{\mu}\equiv\frac{2i\epsilon_{\mu\nu\rho\sigma}p^{\nu}_{\ell}P^{\rho}_{h}P^{\sigma}_{N}}{\sqrt{s_{\ell h}s_{\ell N}s_{hN}}}\,,\quad v\cdot v=-\vec{v}\cdot\vec{v}=-1\,.$ (5)
We introduce the notation of transverse momentum imbalance between the scattered lepton $\ell^{\prime}$ and the identified hadron $h$
$q_{\mathrm{y}}\equiv\vec{q}_{\perp}\cdot\vec{v}=-q_{\perp}\cdot v\,$, (6)
where $q_{\mathrm{y}}$ quantifies the deviation of the scattered lepton $\ell^{\prime}$ from the event plane, its relation to the azimuthal angle is
$\sin\left(\pi-\phi_{\ell^{\prime}h}\right)=\frac{q_{\mathrm{y}}}{q_{T}}\,.$ (7)
In this way, we are considering the $q_{\ast}$ spectrum ($q_{\ast}=|q_{\mathrm{y}}|$) in the lepton-target frame as a variant of the conventional $q_{T}$ spectrum in the photon-target frame. Note however, in addition to the laboratory frame construction, $\epsilon(q,p_{\ell},P_{h},P_{N})$ can also be interpreted as momentum shift of the lepton in Breit frame or momentum shift of the hadron in Trento frame Bacchetta:2004jz .
The $q_{\ast}$ spectrum with $q_{\ast}\ll Q$ delivers sensitivity to the TMD dynamics, the partonic picture can be imaged as follows: if only born-level contribution $N+\gamma^{\ast}\to h$ were considered, then the hadron, the Proton, and the leptons are coplanar, i.e., the volume form exactly vanishes $\epsilon(p_{\ell^{\prime}},p_{\ell},P_{h},P_{N})\equiv 0$. In general, events with multiple jets $N+\gamma^{\ast}\to h+X$ contribute, but in the back-to-back limit $\epsilon(p_{\ell^{\prime}},p_{\ell},P_{h},P_{N})\to 0$, all the final-state radiations in $X$ must be collimated, with momenta collinear either to the target $N$ or the detected hadron $h$. The collinear splitting carries a non-perturbative transverse scale of $K_{T}$ – the intrinsic scale of TMD motion. As a result, the momentum of the outgoing lepton $\ell^{\prime}$ should balance that transverse momentum flow out of the event plane, leaving small deviations of $q_{\mathrm{y}}\sim K_{T}$ in the $v$-direction. Such an observable is known as transverse momentum imbalance between the scattering-off lepton $\ell^{\prime}$ and the detected hadron $h$, whose differential cross section is given by
$\frac{\mathrm{d}\sigma_{\ell+N\to\ell^{\prime}+h+X}}{\mathrm{d}\delta\vec{p}_{J}^{\perp}\mathrm{d}xy\mathrm{d}yz\mathrm{d}z}=\frac{2\pi\alpha^{2}}{4Q^{4}}L_{\mu\nu}(P_{\ell},P_{\ell^{\prime}})W^{\mu\nu}(q,P_{N},P_{h})\,.$ (8)
where $\delta\vec{p}_{J}^{\perp}=\vec{P}_{h}^{\perp}/z-\vec{q}_{\perp}$ is the difference in transverse momentum between the jet and the photon, $q_{\mathrm{y}}=\delta\vec{p}_{J}^{\perp}\cdot v$ in the lepton-hadron frame. Moreover, the hadronic/lepton tensors are
$W^{\mu\nu}(q,P_{N},P_{h})=$ $\prod_{X}\frac{\mathrm{d}^{3}P_{X}}{(2\pi)^{3}2E_{X}}\delta^{(4)}(q+P_{N}-P_{h}-P_{X})\langle N|J^{\dagger\mu}|h,X\rangle\langle h,X|J^{\nu}|N\rangle\,.$
$L^{\mu\nu}=$ $2\delta_{\lambda_{\ell}\lambda_{\ell^{\prime}}}\left[\left(p_{\ell}^{\mu}p_{\ell^{\prime}}^{\nu}+p_{\ell}^{\nu}p_{\ell^{\prime}}^{\mu}-\frac{Q^{2}}{2}g^{\mu\nu}\right)+i\lambda_{l}\epsilon^{\mu\nu\ell\ell^{\prime}}\right]\,.$ (9)
To separate perturbative and non-perturbative dynamics associated with the scale hierarchy $q_{\ast}\ll Q$, we will derive a factorization formulae from perspectives of Soft-Collinear Effective Theory (SCET) Bauer:2000ew ; Bauer:2000yr ; Bauer:2001ct ; Bauer:2001yt . The active modes in SCET are collinear partons moving in well-separated light-like directions and soft gluons mediating long-range interactions among the jets. After BPS field redefinition Bauer:2001yt , soft and collinear sectors decouple in the Lagrangian. Physically, this reflects that soft gluons cannot resolve the internal structure of a jet, but are only sensitive to its total quantum numbers and direction. In SCET, multi-direction collinear jets interact through a complete set of hard operators $\mathcal{O}_{i}$. The corresponding Wilson coefficients $\mathcal{C}_{i}$ encode the short-distance dynamics, with the latter modded out from the low-energy theory. The hard operators are constructed such that their ultraviolet divergences reproduce the infrared divergences of multi-leg amplitudes. In renormalized perturbation theory, the SCET Lagrangian takes the form
$\mathcal{L}_{\text{SCET}}=\sum_{{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}j}=1}^{N}\mathcal{L}_{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}j}+\mathcal{L}_{\color[rgb]{0.7,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.0,0.0}\text{soft}}+\sum_{i}Z_{i}\mathcal{C}_{i}\mathcal{O}_{i}\,.$ (10)
The Lagrangian is soft zero-bin subtracted, which is equivalent to dividing the collinear fields by the vacuum expectation values of zero-bin Wilson lines Feige:2014wja ; Lee:2006nr . At amplitude level when $|X\rangle\simeq|X_{\color[rgb]{0.7,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.0,0.0}\text{soft}}\rangle\prod^{N}_{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}j}|X_{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}j}\rangle$, the amplitude is mediated through effective vertices of $\mathcal{O}_{i}$ and is factorized Feige:2014wja , e.g., for the leading jet production in SIDIS
[…]
each hard operator $\mathcal{O}^{\vec{\lambda}}_{\text{scet}}$ is a local product of helicity fields with definite little group scalings Moult:2015aoa . The hard amplitude $\mathcal{C}(Q^{2},\mu)$ disentangles from the infrared (IR) dynamics associated with the hadronic state $\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}N$} , and can be computed at partonic level
[…]
In this case, the leading-power collinear expansion is exact (== instead of $\simeq$), as each collinear trajectory contains just one particle–taking the collinear limit simply does nothing. Furthermore, since pure virtual-loop corrections in SCET are scaleless, they vanish identically, consequently
$\langle{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}q}|\mathcal{O}^{\vec{\lambda}}_{\text{scet}}|{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}q}\rangle=\langle{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}q}|\mathcal{O}^{\vec{\lambda}}_{\text{scet}}|{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}q}\rangle_{\text{tree}}\left(1+\delta_{Z_{\vec{\lambda}}}(\epsilon_{\text{IR}})\right)\,$, (13)
so that the IR poles of a helicity amplitude is reproduced by UV renormalization counter-term diagrams provided by $\delta{Z_{\vec{\lambda}}}={Z_{\vec{\lambda}}}-1$ Moult:2015aoa . This is as expected since the IR divergences in the theory above a matching scale must match with the UV divergences of the low-energy effective theory below that matching scale. The discussion above can be summarized as in the following factorization formulae Feige:2014wja ; Bauer:2000ew ; Bauer:2000yr ; Bauer:2001ct ; Bauer:2001yt
[…]
Once the hard and soft-collinear degrees of freedom are disentangled, the next step is to enumerate all allowed hard events, and within each hard label, zoom in to the soft-collinear subprocesses and count over particles in a proper way: the classes of momenta entering into the soft-collinear phase space differ only by their typical size of energy or rapidity, as a result, the phase space should be integrated over with a rapidity cut-off $\nu$, here we employ exponential cut-off scheme Li:2016axz . Additionally, a zero-bin subtraction is performed to remove double-counting between collinear and soft sectors, yielding genuine collinear beam functions Manohar:2006nz . Consequently, we have hard-collinear-soft factorization
$W^{\mu\nu}\simeq\prod\mathrm{d}X_{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}{\text{c}}\color[rgb]{0.7,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.0,0.0}{\text{s}}}n\cdot\bar{n}\delta(\bar{n}\cdot p-x\bar{n}\cdot P_{N})\delta(n\cdot p_{J}-n\cdot P_{h}/z)\delta^{(2)}(\delta\vec{p}_{J}^{\perp}+\vec{P}^{\perp}_{X_{\color[rgb]{0.0,0.0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.9}{\text{c}}\color[rgb]{0.7,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.0,0.0}{\text{s}}}})\sum_{f}|\mathcal{C}_{f}(Q^{2},\mu)|^{2}$
[…]
$=$ $2n\cdot\bar{n}\,z\sum_{f}H_{f}(Q^{2},\mu)\int\frac{\mathrm{d}b}{2\pi}e^{ib\delta p_{J}^{\mathrm{y}}}\mathcal{S}_{n\bar{n}}(b,\mu)\mathrm{Tr}[\mathscr{B}_{f/N}(S_{\parallel},x,b,E_{n},\mu)\gamma^{\mu}\mathscr{D}_{h/f}(z,b,E_{\bar{n}},\mu)\gamma^{\nu}]\,$, (15)
the factor $2$ in the front originates from spin average for the operator definition of TMD FFs. Furthermore, in the last step of eq. (2), the $\mathrm{x}$-direction integral is performed, so that the impact parameter $b_{\perp}=b\,v$ is aligned along the $\mathrm{y}$-direction. By applying SCET Fierz identity
[…]
to the TMD factorization formula in eq. (2), the spin-correlation decomposes into a spin-sum and a spin-asymmetry contribution
$xyz\frac{\mathrm{d}\sigma_{\ell+N\to\ell^{\prime}+h+X}}{2\mathrm{d}q_{\ast}\mathrm{d}x\mathrm{d}y\mathrm{d}z}\simeq$ $\,\sum_{f}\int\frac{\mathrm{d}b}{2\pi}e^{ibq_{\mathrm{y}}}\,\left[H_{f}(Q^{2},\mu)\mathcal{S}_{n\bar{n}}\left(b_{\perp}\,,\mu\,,\nu\right)\right]z\mathcal{F}_{h/f}\left(z\,,\frac{b_{\perp}}{z}\,,E_{\bar{n}},\mu,\nu\right)$
$\times$ $\left(\sigma^{\text{U}}_{0}\times x\mathcal{B}_{f/N}(x\,,b_{\perp}\,,E_{n},\mu,\nu)+\lambda_{\ell}S_{\parallel}\times\sigma^{\text{L}}_{0}\times x\Delta\mathcal{B}_{f/N}(x\,,b_{\perp}\,,E_{n},\mu,\nu)\right)\,$, (17)
where $\alpha=e_{R}^{2}/(4\pi)$ and $H_{f}(Q^{2},\mu)=|\mathcal{C}_{f}(Q,\mu)\rangle\langle\mathcal{C}_{f}(Q,\mu)|$ is the square of the hard matching coefficient Becher:2006mr ; Baikov:2009bg . The born-level unpolarized cross-section $\sigma^{\text{U}}_{0}$ and its polarized counterpart $\sigma^{\text{L}}_{0}$, are given by
$\sigma^{\text{U}}_{0}=2\pi\alpha^{2}\frac{1+(1-y)^{2}}{Q^{2}}\,,\quad\sigma^{\text{L}}_{0}=2\pi\alpha^{2}\frac{1-(1-y)^{2}}{Q^{2}}\,.$ (18)
The denominator $2\mathrm{d}q_{\ast}$ takes into account the reflection symmetry of the events with respect to the event plane. Again, we stress that our factorization formula consists of genuine collinear beam functions  Collins:1982wa ; Collins:2011zzd ; Echevarria:2012js ; Ji:2004wu
$\mathcal{B}_{q/N}$ $(x\,,b_{\perp}\,,E_{n},\mu,\nu)\equiv\lim_{\nu\to\infty}Z^{q}_{B}(b_{\perp}\,,E_{n},\mu,\nu)\frac{\mathcal{B}^{0}_{q/N}(x\,,b_{\perp}\,,E_{n},\mu,\nu)}{\mathcal{S}^{0}_{n\bar{n}}(b_{\perp}\,,\mu,\nu)}\,$, (19)
where the superscript 0 denotes unsubtracted collinear beam functions. The beam or soft function is defined with a rapidity regulator  Collins:2011zzd ; Becher:2010tm ; Echevarria:2011epo ; Echevarria:2012js ; Chiu:2012ir ; Li:2016axz ; Echevarria:2015usa ; Echevarria:2015byo ; Echevarria:2016scs ; Becher:2011dz ; Ebert:2018gsn . In this work, we adopt exponential regulator Li:2016axz for both soft Li:2016ctv and collinear sectors Luo:2019bmw ; Luo:2019hmp ; Luo:2019szz ; Luo:2020epw , as it explicitly preserves non-abelian exponentiation Gatheral:1983cz ; Frenkel:1984pz , and the zero-bin soft function within exponential regularization scheme is identical to the TMD soft function.
The TMD beam functions and fragmentation functions are the conventional ones but are parameterized one-dimensionally along the $v$-direction since $\frac{b_{\perp}}{b}=v$. Note that for the beam functions or TMD PDFs, the incoming beam provides a canonical reference frame for transverse momentum. In the case of TMD FFs, we have used parton frame TMD fragmentation functions $\mathcal{F}_{h/f}$, where the parton initialing the decay processes has zero transverse momentum. It is related to the hadron-frame TMD fragmentation functions by
$\mathcal{F}_{h/f}\left(z\,,\frac{b_{\perp}}{z}\,,E,\mu,\nu\right)=z^{2}\mathcal{D}_{h/f}\left(z\,,b_{\perp}\,,E,\mu,\nu\right)\,$, (20)
the later is defined by enforcing that the detected hadron has zero transverse momentum Collins:2011zzd ; Luo:2019hmp ; Luo:2019bmw ; Luo:2020epw . The TMD soft correlation function $\mathcal{S}_{n\bar{n}}$ is defined as a vaccum expectation value of the time-ordered Wilson lines
$\qquad\mathbf{S}_{12\dots m}(b_{\perp})=\langle 0|T[\boldsymbol{O}_{n_{1}\dots n_{m}}(0)]\overline{T}[\boldsymbol{O}_{n_{1}\dots n_{m}}^{\dagger}(b_{\perp})]|0\rangle\,$, (21)
where $\boldsymbol{O}_{n_{1}\dots n_{m}}(x)=\prod_{i}^{m}\boldsymbol{Y}_{n_{i}}(x)$, with $\boldsymbol{Y}_{n_{i}}(x)=\exp[i\int\mathrm{d}s\,n_{i}\cdot A_{s}(sn_{i}+x)\mathbf{T}_{i}]$ a semi-infinite light-like soft Wilson line, and $n_{i}=p_{i}/p_{i}^{0}$ the light-like direction of the incoming or outgoing parton with $b_{\perp}\cdot n_{i}=0$. Each Wilson line is aligned with the classic trajectories of the hard particles, from the origin where hard scattering take place to the distant future where final-state particles are detected. The Wilson line can couple to infinitely many soft gluons through eikonal interactions. Thanks to the non-abelian exponentiation theorem Gatheral:1983cz ; Frenkel:1984pz , the TMD soft correlation function is an exponential of web diagrams with a dipole color structure, provided that no more than three Wilson lines are considered Catani:1999ss ; Catani:2000pi ; Gao:2023ivm
$\mathbf{S}_{12\dots m}\equiv\exp\Bigl[-\sum_{i<j\in\{12\dots m\}}\mathbf{T}_{i}\cdot\mathbf{T}_{j}\,\mathbf{s}_{ij}\Bigr]\,,\quad m\leq 3$ (22)
where $\mathbf{s}_{ij}$ is the universal soft exponent, it is related to the conventional back-to-back TMD soft function Li:2016ctv through a boost
$\mathbf{s}_{ij}=\mathbf{s}_{\perp}\,\Bigl(L_{b},L_{\nu}+\ln\frac{n_{i}\cdot n_{j}}{2}\Bigr)\,.$ (23)
For SIDIS soft function, the dipole color structure leads to naive Casimir scaling up to N33LO. Indeed, by color conservation
$\mathcal{S}^{i}_{n\bar{n}}\left(b_{\perp}\,,\mu\,,\nu\right)=\exp\Bigl[C_{i}\,\mathbf{s}_{n\bar{n}}\left(b_{\perp}\,,\mu\,,\nu\right)+\mathcal{O}(\alpha^{4}_{s})\Bigr]\,$, (24)
$C_{i}=C_{F}$ for fundamental representation, $C_{i}=C_{A}$ for adjoint representation. We have assumed here that the TMD soft function $\mathcal{S}^{i}_{n\bar{n}}$ takes the same expression no matter $n$ and $\bar{n}$ being incoming or outgoing. In Collins:2004nx , it was argued with contour deformation that the TMD soft function takes the same universal expression for Drell-Yan, $e^{+}e^{-}$ and semi-inclusive deep elastic scattering.
3 Renormalization Group Evolution and $q_{\ast}$ Spectrum with RG Improved Perturbation Theory
We first introduce two conventional integrals in the context of Sudakov problems Becher:2006mr
$K_{\Gamma}(\mu_{0},\mu)=-\int_{\alpha_{s}(\mu_{0})}^{\alpha_{s}(\mu)}\!\mathrm{d}\alpha_{s}\,\frac{\Gamma(\alpha_{s})}{\beta(\alpha_{s})}\int_{\alpha_{s}(\mu_{0})}^{\alpha_{s}}\!\mathrm{d}\alpha_{s}^{\prime}\,\frac{1}{\beta(\alpha_{s}^{\prime})}\,,\quad A_{\gamma}(\mu_{0},\mu)=-\int_{\alpha_{s}(\mu_{0})}^{\alpha_{s}(\mu)}\!\mathrm{d}\alpha_{s}\,\frac{\gamma(\alpha_{s})}{\beta(\alpha_{s})}\,.$ (25)
From RG consistence, it follows that the $\mu$-evolution is linear in $\ln\mu$ to all orders in $\alpha_{s}$, for all the relevant building blocks B. The hard Wilson coefficients obey the following renormalization group equation
$\frac{\mathrm{d}|\mathcal{C}(\mu)\rangle}{\mathrm{d}\ln\mu}\equiv$ $\,\widehat{\Gamma}_{H}(\mu)|\mathcal{C}(\mu)\rangle\,,\quad\widehat{\Gamma}_{H}(\mu)=\,\gamma_{\mathrm{cusp}}(\alpha_{s})\mathbf{\Delta}_{H}(\mu)+\widehat{\gamma}_{H}(\alpha_{s})\,$, (26)
where the non-cusp piece $\widehat{\gamma}_{H}(\alpha_{s})$ is proportional to the identity operator
$\widehat{\gamma}_{H}(\alpha_{s})=\sum_{i=q,g}\frac{\gamma^{H}_{i}}{2}\mathbf{1}\,.$ (27)
$\mathbf{\Delta}_{H}(\mu)$ is a process-dependent color matrix that does not depend on $\alpha_{s}$
$\mathbf{\Delta}_{H}(\mu)=-\sum_{i<j}{\bf T}_{i}\cdot{\bf T}_{j}\ln\frac{\sigma_{ij}\,s_{ij}-i0}{\mu^{2}}\,$, (28)
$\sigma_{ij}=-1$ if both $i$ and $j$ are incoming or outgoing, and $1$ otherwise. By color identity
$2\sum_{i<j}{\bf{T}}_{i}\cdot{\bf{T}}_{j}=\sum_{i}{\bf{T}}_{i}^{2}$ (29)
we find $\mathbf{\Delta}_{H}(\mu)-\mathbf{\Delta}_{H}(\mu_{0})$ is linear in $\ln\mu$ with trivialized color factors
$\mathbf{\Delta}_{H}(\mu)-\mathbf{\Delta}_{H}(\mu_{0})=\mathbf{1}(n_{g}C_{A}+n_{q}C_{F})\ln\left(\frac{\mu_{0}}{\mu}\right)\,.$ (30)
Consequently, the solution for the evolution in eq. (26) is
$\widehat{U}_{H}(\mu_{0},\mu)=$ $e^{(n_{g}C_{A}+n_{q}C_{F})K_{\mathrm{cusp}}(\mu_{0},\mu)}\times\mathcal{P}\exp\Bigg[-A_{\mathrm{cusp}}(\mu_{0},\mu)\,\mathbf{\Delta}_{H}(\mu_{0})-A_{H}(\mu_{0},\mu)\Bigg]\,.$ (31)
The $\mu$-evolution of the TMD soft function is Kidonakis:1998bk ; Kidonakis:1998nf ; Aybat:2006mz ; Aybat:2006wq
$\frac{\mathrm{d}\mathbf{S}_{1\dots m}}{\mathrm{d}\ln\mu}=\mathbf{\Gamma}_{S}^{\dagger}\cdot\mathbf{S}_{1\dots m}+\mathbf{S}_{1\dots m}\cdot\mathbf{\Gamma}_{S}\,$, (32)
the anomalous dimension $\mathbf{\Gamma}_{S}$ takes the form
$\mathbf{\Gamma}_{S}=\gamma_{\mathrm{cusp}}(\alpha_{s})\mathbf{\Delta}_{S}(\mu,\nu)+\widehat{\gamma}_{S}(\alpha_{s})\,$, (33)
where $\mathbf{\Delta}_{S}(\nu;\mu)$ is a process-dependent color matrix that does not depend on $\alpha_{s}$
$\mathbf{\Delta}_{S}(\nu;\mu)=\sum_{i<j}{\bf T}_{i}\cdot{\bf T}_{j}\ln\left(\frac{\sigma_{ij}\,n_{i}\cdot n_{j}-i0}{2}\right)\left(\frac{\nu^{2}}{\mu^{2}}\right)\,$, (34)
and the soft anomalous $\widehat{\gamma}_{S}(\alpha_{s})$ takes the form
$\widehat{\gamma}_{S}(\alpha_{s})=-\sum_{i}\frac{1}{2}\gamma^{S}_{i}\mathbf{1}\,.$ (35)
The $\mu$-evolution for the TMD soft correlation functions is solved by
$\widehat{U}_{S}(\nu;\mu_{0},\mu)=$ $e^{-(n_{g}C_{A}+n_{q}C_{F})K_{\mathrm{cusp}}(\mu_{0},\mu)-A_{S}(\mu_{0},\mu)}\times\mathcal{P}\exp\Bigg[-A_{\mathrm{cusp}}(\mu_{0},\mu)\,\mathbf{\Delta}_{S}(\mu_{0})\Bigg]\,.$ (36)
The rapidity renormalization group (RRGs) Chiu:2011qc ; Chiu:2012ir evolution is
$\frac{\mathrm{d}\mathbf{S}_{1\dots m}}{\mathrm{d}\ln\nu}=\frac{1}{2}\left(\mathbf{\Gamma}_{R}^{\dagger}(\mu,\mu_{b})\cdot\mathbf{S}_{1\dots m}+\mathbf{S}_{1\dots m}\cdot\mathbf{\Gamma}_{R}(\mu,\mu_{b})\right)\,$, (37)
where $\mathbf{\Gamma}_{R}(\mu,\mu_{b})$ denotes the Collins-Soper (CS) kernel Collins:1981uk ; Collins:1987pm ; Collins:2011zzd
$\mathbf{\Gamma}_{R}(\mu,\mu_{b})=$ $\,\sum_{i}\left(-\int^{\mu^{2}}_{\mu_{b}^{2}}\frac{d\bar{\mu}^{2}}{\bar{\mu}^{2}}\gamma^{i}_{\rm cusp}[\alpha_{s}(\bar{\mu})]+\gamma^{i}_{R}[\alpha_{s}(\mu_{b})]\right)\,$, (38)
its $\mu$-evolution is controlled by the cusp anomalous dimension $\gamma^{i}_{\rm cusp}$ which originates from the UV renormalization of the cusped Wilson loops Brandt:1981kf ; Brandt:1982gz . On the other hand, the rapidity anomalous dimension $\gamma^{i}_{R}(\mu_{b})$ has both perturbative part as $b_{T}\to 0$ and genuine non-perturbative part as $b_{T}\to\infty$. Tremendous progress has been made in the nonperturbative study of the Collins–Soper kernel Kang:2024dja ; LatticePartonLPC:2022eev ; Shanahan:2020zxr ; Shanahan:2021tst ; Avkhadiev:2024mgd ; Schlemmer:2021aij ; Li:2021wvl ; LatticeParton:2020uhz , and perturbative accuracy has now reached N4LL Li:2016ctv ; Moult:2022xzt ; Duhr:2022yyp . Up to N3LL, naive Casimir scaling holds and one may write, e.g.,
$\gamma_{i}^{S}=C_{i}\gamma^{S},\quad\gamma^{i}_{R}=C_{i}\gamma_{R}\,.$ (39)
To N4LL strict Casimir scaling is violated due to the presence of the quartic Casimirs, requiring a generalized scaling scheme Boels:2017ftb ; Boels:2017fng ; Boels:2017skl ; Lee:2019zop ; Lee:2021uqq ; Lee:2022nhh ; Das:2019btv ; Das:2020adl ; Moch:2018wjh ; Henn:2019rmi ; Grozin:2017css ; vonManteuffel:2020vjv ; Agarwal:2021zft ; Moult:2022xzt ; Duhr:2022yyp . The rapidity RG for TMD soft function is solved by
$U_{R}(\nu_{S},\nu;\mu_{S})=\exp\left[\int_{\nu_{S}}^{\nu}\mathbf{\Gamma}_{R}(\mu_{S},\mu_{b})\mathrm{d}\ln\bar{\nu}\right]=\left(\frac{\nu^{2}}{\nu_{S}^{2}}\right)^{(n_{g}C_{A}+n_{q}C_{F})\left(-A_{\mathrm{cusp}}(\mu_{S},\mu_{b})+\frac{1}{2}\gamma_{R}(\mu_{b})\right)}\,.$ (40)
The combined solution of $\mu$-RG and $\nu$-RG leads to a two-dimensional kernel, which simultaneously resums large logarithms in factorization and rapidity scales. A convenient scale evolution scheme is to evolve the hard and soft sectors from their natural scales to the common factorization scale $(\mu\,,\nu)$ which we set to be the natural scales of the TMD beam functions, this leads to the following evolution kernels for perturbative resummation
$\widehat{U}_{H}\widehat{U}^{\dagger}_{H}(\mu_{H},\mu)\widehat{U}_{S}\widehat{U}^{\dagger}_{S}(\nu;\mu_{S},\mu)U_{R}(\nu_{S},\nu;\mu_{S})=\exp\bigg[-2A_{H}(\mu_{H},\mu)-2A_{S}(\mu_{S},\mu)$
$+$ $(n_{g}C_{A}+n_{q}C_{F})\left(2K_{\mathrm{cusp}}(\mu_{H},\mu_{S})+\left(\frac{1}{2}\gamma_{R}(\mu_{b})-A_{\mathrm{cusp}}(\mu_{S},\mu_{b})\right)\ln\frac{\nu^{2}}{\nu_{S}^{2}}\right)-A_{\mathrm{cusp}}(\mu_{H},\mu_{S})$
$\times$ $\left(-2\sum_{i<j}{\bf{T}}_{i}\cdot{\bf{T}}_{j}\ln\frac{s_{ij}}{\mu_{H}^{2}}\right)-A_{\mathrm{cusp}}(\mu_{S},\mu)\left(-2\sum_{i<j}{\bf{T}}_{i}\cdot{\bf{T}}_{j}\left(\ln\frac{s_{ij}}{\nu^{2}}-\ln\frac{n_{i}\cdot n_{j}}{2}\right)\right)\bigg]\,.$ (41)
Eq. (3) provides a general evolution function for processes admitting a color-dipole structure, and is expected to hold for reactions with no more than 3-jets. RG invariance of the cross section is guaranteed by linear relations among the anomalous dimensions. Indeed, with a bit color algebra we find
$2\mathbf{\Delta}_{H}(\mu)+2\mathbf{\Delta}_{S}(\nu;\mu)=2\sum_{i}\gamma^{i}_{\textrm{cusp}}\ln{\frac{2E_{i}}{\nu}}\,.$ (42)
This is consistent with the RG equations of various TMD beam/fragmentation functions $\mathcal{G}_{i}\,,i\in\{q\,,g\}$
$\frac{\mathrm{d}}{\mathrm{d}\ln\mu}\ln\mathcal{G}_{i}=-\gamma^{i}_{\rm cusp}\ln\frac{4E_{i}^{2}}{\nu^{2}}+\gamma^{B}_{i}\,$, (43)
provided that $\gamma_{i}^{H}+\gamma_{i}^{B}-\gamma_{i}^{S}=0$ holds to all orders. The beam/fragmentation functions are regulator dependent, regulator independent physical TMDs are obtained by dressing them with the cloud of soft gluons
$f^{q}_{1}\left(x\,,b_{\perp}\,,\xi^{n},\mu\right)=$ $\,\mathcal{B}_{q/N}(x\,,b_{\perp}\,,E_{n},\mu,\nu)\sqrt{\mathcal{S}_{n\bar{n}}(b_{\perp},\mu,\nu)}\,$,
$D^{q}_{1}\left(z\,,\frac{b_{\perp}}{z}\,,\xi^{\bar{n}},\mu\right)=$ $\,\mathcal{F}_{h/q}\left(z\,,\frac{b_{\perp}}{z}\,,E_{\bar{n}},\mu,\nu\right)\sqrt{\mathcal{S}_{n\bar{n}}(b_{\perp},\mu,\nu)}\,.$ (44)
They obey Collins-Soper equations Collins:1981uk ; Collins:1987pm ; Collins:2011zzd
$\frac{\mathrm{d}}{\mathrm{d}\ln\sqrt{\xi^{i}}}\ln\mathcal{\widetilde{G}}_{i}(x_{i},b_{\perp},\xi^{i},\mu)\equiv$ $\,\Gamma_{R}^{i}(\mu,\mu_{b})=-2A^{i}_{\rm cusp}(\mu,\mu_{b})+\gamma^{i}_{R}(\mu_{b})\,$,
$\frac{\mathrm{d}}{\mathrm{d}\ln\mu}\ln\mathcal{\widetilde{G}}_{i}(x_{i},b_{\perp},\xi^{i},\mu)\equiv$ $\,\Gamma_{\mu}^{i}(\mu,\xi^{i})=\,-\gamma^{i}_{\rm cusp}(\alpha_{s}(\mu))\ln\frac{\xi^{i}}{\mu^{2}}-\gamma^{i}_{H}\,$, (45)
where $\mu_{b}=b_{0}/b,b_{0}=2e^{-\gamma_{E}}$, $\xi^{i}$ are the Collins-Soper scales Collins:1981uk ; Collins:1987pm ; Collins:2011zzd . These equations are consistent with SCET RRGs Chiu:2011qc ; Chiu:2012ir since the rapidity and hard scale in eq. (43) are paired. Solution to the two-dimensional Collins-Soper equations are
$\frac{\mathcal{\widetilde{G}}_{i}(x_{i},b_{\perp},\xi^{i},\mu)}{\mathcal{\widetilde{G}}_{i}(x^{i},b_{\perp},\xi_{0}^{i},\mu_{0})}=$ $\exp\bigg[\int_{\mu_{0}}^{\mu}\Gamma_{\bar{\mu}}^{i}(\bar{\mu},\xi^{i})\mathrm{d}\ln\bar{\mu}+\int_{\xi^{i}_{0}}^{\xi^{i}}{\Gamma}_{R}(\mu_{0},\mu_{b})\mathrm{d}\ln\sqrt{\bar{\xi^{i}}}\bigg]$
$=$ $e^{-2K^{i}_{\rm cusp}(\mu_{0},\mu)+A^{i}_{H}(\mu_{0},\mu)}\times\left(\frac{\xi^{i}}{\mu^{2}_{0}}\right)^{A^{i}_{\rm cusp}(\mu_{0},\mu)}\times\left(\frac{\xi^{i}}{\xi^{i}_{0}}\right)^{-A^{i}_{\rm cusp}(\mu_{0},\mu_{b})+\frac{1}{2}\gamma_{R}^{i}(\mu_{b})}\,.$ (46)
Putting everything together, we obtain a TMD factorization theorem for $q_{\ast}$ spectrum, expressed directly in terms of physical TMDs
$xyz\frac{\mathrm{d}\sigma_{\ell+N\to\ell^{\prime}+h+X}}{2\mathrm{d}q_{\ast}\mathrm{d}x\mathrm{d}y\mathrm{d}z}\simeq$ $\,\sum_{q}\int\frac{\mathrm{d}b}{2\pi}e^{ibq_{\mathrm{y}}}\,H_{q}(Q^{2},\mu)zD_{1}^{q}\left(z,\frac{b_{\perp}}{z},\xi^{\bar{n}}_{0},\mu_{0}\right)$
$\times$ $\left(\sigma^{\text{U}}_{0}\times xf_{1}^{q}(x,b_{\perp},\xi^{n}_{0},\mu_{0})+\lambda_{\ell}S_{\parallel}\times\sigma^{\text{L}}_{0}\times xg_{1}^{q}(x,b_{\perp},\xi^{n}_{0},\mu_{0})\right)$
$\times$ $\,\prod_{i}e^{-2K^{i}_{\rm cusp}(\mu_{0},\mu)+A^{i}_{H}(\mu_{0},\mu)}\times\left(\frac{\xi^{i}}{\mu^{2}_{0}}\right)^{A^{i}_{\rm cusp}(\mu_{0},\mu)}\times\left(\frac{\xi^{i}}{\xi^{i}_{0}}\right)^{-A^{i}_{\rm cusp}(\mu_{0},\mu_{b})+\frac{1}{2}\gamma_{R}^{i}(\mu_{b})}\,$, (47)
where the Collins-Soper scales are
$\xi^{n}=Q^{2}\frac{x}{x+y-xy}\,,\quad\xi^{\bar{n}}=Q^{2}\frac{x+y-xy}{x}\,.$ (48)
The factorization framework of eq. (3) was pioneered by Collins, Soper, and Sterman (CSS) Collins:1981uk ; Collins:1981uw ; Collins:1984kg and expanded in Collins:2011zzd ; Ji:2004wu ; deFlorian:2001zd ; Catani:2000vq ; Catani:2010pd ; Aybat:2011zv ; Ji:2005nu , and later reformulated in the framework of Soft Collinear Effective Theory Becher:2010tm ; Becher:2011xn ; Becher:2012yn ; Echevarria:2011epo ; Echevarria:2012js ; Echevarria:2014rua ; Ebert:2019tvc ; Collins:2017oxh .
Figure 1: Resummed $q_{\ast}$ distributions at EIC energies.
Figure 2: Resummed $q_{\ast}$ distributions at COMPASS energies.
Figure 3: Impact of non-perturbative model function of BLNY18 Sun:2014dqm for pion $q_{\ast}$-distribution. Here, all the results are implemented with $b_{\ast}$ prescription, for otherwise the integration diverges.
The N3LL predictions for eq. (3) is shown in fig. (1,2) for both unpolarized and polarized distributions, where we incorporate the non-perturbative model given in Ref Sun:2014dqm
$S_{\text{NP}}^{\text{SIDIS}}=\left({g_{1}\over 2}+{g_{h}\over z_{h}^{2}}\right)b^{2}+g_{2}\ln\left(\frac{b}{b_{\ast}}\right)\ln\left(\frac{Q}{Q_{0}}\right)\,$,
and accordingly, the $b^{\ast}$ prescription Collins:2014jpa to regulate Landau poles
$b\to b^{\ast}=\frac{b}{\sqrt{1+b^{2}/b^{2}_{\text{max}}}}\,,\quad b_{\text{max}}=1.5\text{GeV}^{-1}\,.$ (50)
The impact of the non-perturbative model is shown in fig. (3).
4 Twist-2 Matching of helicity TMDs
The factorization formulae in eq. (3) does not require a strict scale hierarchy between $q_{\ast}$ and $\Lambda_{\text{QCD}}$. However when $Q\gg q_{\ast}\gg\Lambda_{\text{QCD}}$, we will have two equivalent constructions of the effective field theory that has dynamic degrees of freedom carrying momenta of order $\Lambda_{\text{QCD}}$. To obtain the EFTs associated with the scale hierarchy $Q\gg q_{\ast}\gg\Lambda_{\text{QCD}}$ to the leading power, one may either integrate out momenta of off-shellness $\simeq q_{\ast}\gg\Lambda_{\text{QCD}}$ from the TMDs, or instead integrate out the modes above $\Lambda_{\text{QCD}}$ once for all. This corresponds to twist-2 matching onto the PDFs/FFs, with DGLAPs resumming large logarithms associated with $\Lambda_{\text{QCD}}$. Owing to the consistency between TMD factorization and the standard collinear factorization, TMDs admit matching onto conventional PDFs/FFs, we derive their N3LO matching coefficients in this section, which provide the boundary conditions for the numerical results presented in Figs. (1,2).
4.1 Operator definitions for the TMD helicity PDFs and FFs
The TMD helicity parton distribution functions (TMD helicity PDFs) can be defined in terms of SCET Bauer:2000ew ; Bauer:2000yr ; Bauer:2001yt ; Bauer:2002nz ; Beneke:2002ph collinear fields
$\mathcal{B}_{q/N}^{\rm bare}(x,b_{\perp},\lambda_{N})=$ $\int\frac{db^{-}}{2\pi}\,e^{-ixb^{-}P^{+}}\langle N(P),\lambda_{N}|\bar{\chi}_{n}(0,b^{-},b_{\perp})\frac{\not{\bar{n}}}{2}\gamma_{5}\chi_{n}(0)|N(P),\lambda_{N}\rangle\,$,
${\cal B}_{g/N}^{{\rm bare}}(x,b_{\perp},\lambda_{N})=$ $-xP_{+}\int\frac{db^{-}}{2\pi}e^{-ixb^{-}P^{+}}\langle N(P),\lambda_{N}|\widetilde{\cal A}_{n\perp}^{a,\mu}(0,b^{-},b_{\perp}){\cal A}_{n\perp}^{a,\mu}(0)|N(P),\lambda_{N}\rangle\,$, (51)
where $N(P)$ is a hadron state with momentum $P^{\mu}=(\bar{n}\cdot P)n^{\mu}/2=P^{+}n^{\mu}/2$, and $n^{\mu}=(1,0,0,1)$ and $\bar{n}^{\mu}=(1,0,0,-1)$. $\chi_{n}=W_{n}^{\dagger}\xi_{n}$ is the gauge invariant collinear quark field Bauer:2001ct in SCET, constructed from collinear quark field $\xi_{n}$ and path-ordered collinear Wilson line $W_{n}(x)={\cal P}\exp\left(ig\int_{-\infty}^{0}ds\,\bar{n}\cdot A_{n}(x+\bar{n}s)\right)$. $A_{n\perp}^{a,\mu}$ is the gauge invariant collinear gluon field with color index $a$ and Lorentz index $\mu$, and the dual collinear gluon field $\widetilde{A}_{n\perp}^{a,\mu}$ is defined in terms of the transverse Levi-Civita tensor
$\widetilde{A}_{n\perp}^{a,\mu}\equiv i\epsilon_{\perp}^{\mu\nu}A_{n\perp,\nu}^{a}\,,\quad\text{and}\quad\epsilon_{\perp}^{\mu\nu}=\frac{P_{\sigma}\bar{n}_{\rho}}{P\cdot\bar{n}}\epsilon^{\mu\nu\sigma\rho}\,.$ (52)
The helicity distribution function depends on the target helicity $\lambda_{N}=\pm$, however, by rotation invariance and parity symmetry, the dependence is trivially linear in the target helicity, and we define $\Delta\mathcal{B}_{i/N}$ through
$\lambda_{N}\Delta\mathcal{B}_{i/N}^{\rm bare}(x,b_{\perp})\equiv$ $\,\mathcal{B}_{i/N}^{\rm bare}(x,b_{\perp},\lambda_{N})\,.$ (53)
For sufficiently small $b_{\perp}$, the TMD PDFs in Eq. (4.1) admit operator product expansion onto the collinear helicity PDFs,
$\Delta\mathcal{B}_{i/N}^{\rm bare}(x,b_{\perp})=$ $\sum_{i}\int_{x}^{1}\frac{d\xi}{\xi}\,\Delta\mathcal{I}_{ij}^{\rm bare}(\xi,b_{\perp})\,\Delta\phi_{j/N}^{\rm bare}(x/\xi)+\mathcal{O}(\Lambda_{\text{QCD}}b_{\perp})\,$, (54)
[…]
Xn
dPSXn
|XnXn|
[…]
into the operator definition, the bare TMD PDFs can be computed from splitting amplitudes integrated over collinear phase space
[…]
where $K^{\mu}$ is the total momentum of $|X_{n}\rangle$, and $d\text{PS}_{X_{n}}$ is the collinear phase space measure, and $\Delta{\rm\mathbb{P}}^{\sigma\rho}_{i\leftarrow j}$ is the space-like spin correlator in the target helicity space
$\Delta{\rm\mathbb{P}}^{\sigma\tau}_{q\leftarrow j}\equiv{\rm\mathbb{Sp}}_{X_{n}q_{l}^{*}\leftarrow j_{\sigma}}^{*}\Gamma^{ls}{\rm\mathbb{Sp}}_{X_{n}q_{s}^{*}\leftarrow j_{\rho}}\equiv\langle j_{\sigma}|\bar{\chi}_{n}|X_{n}\rangle\frac{\not{\bar{n}}}{2}\gamma_{5}\langle X_{n}|\chi_{n}|j_{\rho}\rangle$
$\Delta{\rm\mathbb{P}}^{\sigma\tau}_{g\leftarrow j}\equiv{\rm\mathbb{Sp}}_{X_{n}g_{\mu}^{*}\leftarrow j_{\sigma}}^{*}\Gamma^{\mu\nu}{\rm\mathbb{Sp}}_{X_{n}g_{\nu}^{*}\leftarrow j_{\rho}}\equiv-xP_{+}\langle j_{\sigma}|{\cal A}^{n\perp}_{a,\mu}|X_{n}\rangle i\epsilon^{\mu\nu}_{\perp}\langle X_{n}|{\cal A}^{n\perp}_{a,\nu}|j_{\rho}\rangle\,.$ (64)
The spin density matrices for the partons in Eq. (4.1) are
$\Gamma^{ls}=\left[\frac{\not{\bar{n}}}{2}\gamma_{5}\right]^{ls}\,,\quad\text{and}\quad\Gamma^{\mu\nu}=i\epsilon^{\mu\nu}_{\perp}\,.$ (65)
According to eq. (53), the target spin projectors $\bar{\Gamma}$ should be given by
[…]
where the factor $\mathcal{N}_{\text{sch.}}$ depends on the definition of $\gamma_{5}$ in dimensional regularization. Furthermore, the integral in Eq. (63) is properly defined only with a rapidity cutoff implemented through the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$. The TMD FFs are defined as crossings of TMD beam functions, and according to parton-hardron frame relation given in eq. (20Collins:2011zzd ; Luo:2019hmp ; Luo:2019bmw , we have, in dimensional regularization
[…]
where $P^{\mu}=(\bar{n}\cdot P)n^{\mu}/2=P^{+}n^{\mu}/2$ is the momentum of the final state detected hadron. Similar to Eq.(63), we can compute the partonic TMD helicity FFs by
[…]
where $\Delta{\rm\mathbb{P}}^{T,\sigma\rho}_{i\leftarrow j}$ is the square of the time-like splitting amplitude, which can be obtained from the space-like ones in Eq. (4.1) by analytical continuation. The spin projectors $\Gamma$ in Eq. (75) and the dual spin projectors $\bar{\Gamma}$ in Eq. (93) are formally identical to those of the space-like ones in Eq. (65), but $P$ is now the momentum of the detected hardron.
4.2 Collinear mass factorization and renormalization group equations
The bare TMD helicity PDFs or FFs have divergencies both of UV and IR origin, the UV renormalization, the zero-bin subtraction and mass-factorization are technically identical to those of the unpolarized case Luo:2020epw , and are summarized in the following collinear mass factorization formula
$\frac{1}{Z_{i}^{B}}\frac{\Delta{\cal B}_{ij}^{{\rm bare}}(x,b_{\perp})}{\mathcal{S}_{0\rm b}}=$ $\sum_{k}\Delta\mathcal{I}_{ik}(x,b_{\perp},\mu,\nu)\otimes\Delta\phi_{kj}(x,\mu)\,$,
$\frac{1}{Z_{i}^{B}}\frac{\Delta{\cal F}_{ij}^{{\rm bare}}(z,b_{\perp}/z)}{\mathcal{S}_{0\rm b}}=$ $\sum_{k}\Delta d_{ik}(z,\mu)\otimes\Delta\mathcal{C}_{kj}(z,b_{\perp}/z,\mu,\nu)\,$, (94)
where $\mathcal{S}_{0\rm b}(\alpha_{s})$ is the bare zero-bin soft function which is the same as the TMD soft function Li:2016ctv . $Z_{i}^{B}$ (see in Sec. E) are the multiplicative operator renormalization constants for TMD PDFs and FFs. $\Delta\mathcal{I}_{ij}$ (and $\Delta\mathcal{C}_{ij}$) are the finite coefficient functions. $\Delta\phi_{ki}$ (and $\Delta d_{ik}$) are the partonic helicity PDFs (and FFs), they evolve with the helicity splitting functions Mertig:1995ny ; Vogelsang:1995vh ; Vogelsang:1996im ; Zijlstra:1993sh ; Moch:2014sna ; Moch:2015usa ; Behring:2019tus ; Stratmann:1996hn ; Rijken:1997rg . Up to $\alpha_{s}^{3}$, we have
$\Delta\phi_{ij}(x,\alpha_{s})=$ $\delta_{ij}\delta(1-x)-\frac{\alpha_{s}}{4\pi}\frac{{\Delta P}^{(0)}_{ij}(x)}{\epsilon}$
$+$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\bigg[\frac{1}{2\epsilon^{2}}\biggl(\sum_{k}{\Delta P}^{(0)}_{ik}\otimes{\Delta P}^{(0)}_{kj}(x)+\beta_{0}{\Delta P}^{(0)}_{ij}(x)\biggl)-\frac{1}{2\epsilon}{\Delta P}^{(1)}_{ij}(x)\bigg]$
$+$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg[\frac{-1}{6\epsilon^{3}}\biggl(\sum_{m\,,k}{\Delta P}^{{(0)}}_{im}\otimes{\Delta P}^{{(0)}}_{mk}\otimes{\Delta P}^{{(0)}}_{kj}(x)+3\beta_{0}\sum_{k}{\Delta P}^{{(0)}}_{ik}\otimes{\Delta P}^{{(0)}}_{kj}(x)$
$+$ $2\beta_{0}^{2}{\Delta P}^{{(0)}}_{ij}(x)\biggl)+\frac{1}{6\epsilon^{2}}\biggl(\sum_{k}{\Delta P}^{{(0)}}_{ik}\otimes{\Delta P}^{{(1)}}_{kj}(x)+2\sum_{k}{\Delta P}^{{(1)}}_{ik}\otimes{\Delta P}^{{(0)}}_{kj}(x)$
$+$ $2\beta_{0}{\Delta P}^{{(1)}}_{ij}(x)+2\beta_{1}{\Delta P}^{{(0)}}_{ij}(x)\biggl)-\frac{1}{3\epsilon}{\Delta P}^{(2)}_{ij}(x)\bigg]+\mathcal{O}(\alpha_{s}^{3})\,$, (95)
and for the FFs
$\Delta d_{ij}(z,\alpha_{s})=$ $\delta_{ij}\delta(1-z)-\frac{\alpha_{s}}{4\pi}\frac{{\Delta P}^{T{(0)}}_{ij}(z)}{\epsilon}$
$+$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\bigg[\frac{1}{2\epsilon^{2}}\biggl(\sum_{k}{\Delta P}^{T{(0)}}_{ik}\otimes{\Delta P}^{T{(0)}}_{kj}(z)+\beta_{0}{\Delta P}^{T{(0)}}_{ij}(z)\biggl)-\frac{1}{2\epsilon}{\Delta P}^{T{(1)}}_{ij}(z)\bigg]$
$+$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg[\frac{-1}{6\epsilon^{3}}\biggl(\sum_{m\,,k}{\Delta P}^{T{(0)}}_{im}\otimes{\Delta P}^{T{(0)}}_{mk}\otimes{\Delta P}^{T{(0)}}_{kj}(z)+3\beta_{0}\sum_{k}{\Delta P}^{T{(0)}}_{ik}\otimes{\Delta P}^{T{(0)}}_{kj}(z)$
$+$ $2\beta_{0}^{2}{\Delta P}^{T{(0)}}_{ij}(z)\biggl)+\frac{1}{6\epsilon^{2}}\biggl(2\sum_{k}{\Delta P}^{T{(0)}}_{ik}\otimes{\Delta P}^{T{(1)}}_{kj}(z)+\sum_{k}{\Delta P}^{T{(1)}}_{ik}\otimes{\Delta P}^{T{(0)}}_{kj}(z)$
$+$ $2\beta_{0}{\Delta P}^{T{(1)}}_{ij}(z)+2\beta_{1}{\Delta P}^{T{(0)}}_{ij}(z)\biggl)-\frac{1}{3\epsilon}{\Delta P}^{T{(2)}}_{ij}(z)\bigg]+\mathcal{O}(\alpha_{s}^{3})\,$, (96)
where ${\Delta P}_{ij}^{(n)}$ is the $\text{N}^{n}$LO space-like helicity splitting function, which is presently known to NNLO Mertig:1995ny ; Vogelsang:1995vh ; Vogelsang:1996im ; Zijlstra:1993sh ; Moch:2014sna ; Moch:2015usa ; Behring:2019tus . ${\Delta P}_{ij}^{T(n)}$ is the $\text{N}^{n}$LO time-like helicity splitting function, which is computed at NLO in the work of Refs Stratmann:1996hn ; Rijken:1997rg .
The factorized finite coefficient functions obey the following $\mu$-RG equations
$\frac{d}{d\ln\mu}\Delta\mathcal{I}_{ji}(x,b_{\perp},\mu,\nu)=2\bigg[\Gamma^{\text{cusp}}_{j}(\alpha_{s}(\mu))\ln\frac{\nu}{xP_{+}}+$ $\gamma^{B}_{j}(\alpha_{s}(\mu))\bigg]\Delta\mathcal{I}_{ji}(x,b_{\perp},\mu,\nu)$
- $2\sum_{k}\Delta\mathcal{I}_{jk}(x,b_{\perp},\mu,\nu)\otimes{\Delta P}_{ki}(x,\alpha_{s}(\mu))\,$, (97)
$\frac{d}{d\ln\mu}\Delta\mathcal{C}_{ij}(z,b_{\perp}/z,\mu,\nu)=2\bigg[\Gamma^{\text{cusp}}_{j}(\alpha_{s}(\mu))\ln\frac{z\nu}{P_{+}}+$ $\gamma^{B}_{j}(\alpha_{s}(\mu))\bigg]\Delta\mathcal{C}_{ij}(z,b_{\perp}/z,\mu,\nu)$
- $2\sum_{k}{\Delta P}^{T}_{ik}(z,\alpha_{s}(\mu))\otimes\Delta\mathcal{C}_{kj}(z,b_{\perp}/z,\mu,\nu)\,$, (98)
and the rapidity evolution equations Chiu:2011qc ; Chiu:2012ir
$\frac{d}{d\ln\nu}\Delta\mathcal{I}_{ji}(x,b_{\perp},\mu,\nu)=$ $-2\left[\int_{\mu}^{b_{0}/b_{T}}\frac{d\bar{\mu}}{\bar{\mu}}\Gamma^{\text{cusp}}_{j}(\alpha_{s}(\bar{\mu}))+\gamma^{R}_{j}(\alpha_{s}(b_{0}/b_{T}))\right]\Delta\mathcal{I}_{ji}(x,b_{\perp},\mu,\nu)\,$,
$\frac{d}{d\ln\nu}\Delta\mathcal{C}_{ij}(z,b_{\perp}/z,\mu,\nu)=$ $-2\left[\int_{\mu}^{b_{0}/b_{T}}\frac{d\bar{\mu}}{\bar{\mu}}\Gamma^{\text{cusp}}_{j}(\alpha_{s}(\bar{\mu}))+\gamma^{R}_{j}(\alpha_{s}(b_{0}/b_{T}))\right]\Delta\mathcal{C}_{ij}(z,b_{\perp}/z,\mu,\nu)\,.$ (99)
We present the coefficient function oder-by-order in terms of strong coupling $\alpha_{s}(\mu)/(4\pi)$, up to $\mathcal{O}(\alpha_{s}^{3})$, the solution to above evolution equations reads
$\Delta\mathcal{I}^{(0)}_{ji}(x,b_{\perp},$ $\,\mu,\nu)=\delta_{ji}\delta(1-x)\,$,
$\Delta\mathcal{I}^{(1)}_{ji}(x,b_{\perp},$ $\,\mu,\nu)=\left(-\frac{\Gamma^{\text{cusp}}_{0}}{2}L_{\perp}L_{Q}+\gamma_{0}^{B}L_{\perp}+\gamma_{0}^{R}L_{Q}\right)\delta_{ji}\delta(1-x)-{\Delta P}_{ji}^{(0)}(x)L_{\perp}+\Delta I_{ji}^{(1)}(x)\,$,
$\Delta\mathcal{I}^{(2)}_{ji}(x,b_{\perp},$ $\,\mu,\nu)=\bigg[\frac{1}{8}\left(-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}\right)\left(-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}+2\beta_{0}\right)L_{\perp}^{2}$
$+$ $\left((-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}+2\beta_{0})\frac{\gamma_{0}^{R}}{2}L_{Q}-\frac{\Gamma^{\text{cusp}}_{1}}{2}L_{Q}+\gamma^{B}_{1}\right)L_{\perp}$
$+$ $\frac{(\gamma_{0}^{R})^{2}}{2}L_{Q}^{2}+\gamma_{1}^{R}L_{Q}\bigg]\,\delta_{ji}\delta(1-x)+\bigg(\frac{1}{2}\sum_{l}{\Delta P}^{(0)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)$
$+$ $\frac{{\Delta P}^{(0)}_{ji}(x)}{2}(\Gamma^{\text{cusp}}_{0}L_{Q}-2\gamma_{0}^{B}-\beta_{0})\bigg)L_{\perp}^{2}+\bigg[-{\Delta P}^{(1)}_{ji}(x)-{\Delta P}^{(0)}_{ji}(x)\gamma_{0}^{R}L_{Q}$
- $\sum_{l}\Delta I^{(1)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)+\left(-\frac{\Gamma^{\text{cusp}}_{0}}{2}L_{Q}+\gamma_{0}^{B}+\beta_{0}\right)\Delta I^{(1)}_{ji}(x)\bigg]L_{\perp}$
$+$ $\gamma_{0}^{R}L_{Q}\Delta I^{(1)}_{ji}(x)+\Delta I^{(2)}_{ji}(x)\,$,
$\Delta\mathcal{I}^{(3)}_{ji}(x,b_{\perp},$ $\,\mu,\nu)=L_{\perp}^{3}\bigg[\left(\frac{1}{2}\beta_{0}+\frac{1}{4}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\sum_{l}{\Delta P}^{(0)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)$
- $\frac{1}{6}\sum_{l\,k}{\Delta P}^{(0)}_{jl}\otimes{\Delta P}^{(0)}_{lk}\otimes{\Delta P}^{(0)}_{ki}(x)+\delta_{ji}\delta(1-x)\left(\frac{1}{6}\beta_{0}^{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right$.
$+$ $\left.\frac{1}{8}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}+\frac{1}{48}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{3}\right)$
$+$ ${\Delta P}^{(0)}_{ji}\left(-\frac{1}{2}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})-\frac{1}{3}\beta_{0}^{2}-\frac{1}{8}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}\right)\bigg]$
$+$ $L_{\perp}^{2}\bigg[\left(-\frac{3}{2}\beta_{0}-\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\sum_{l}\Delta I^{(1)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)$
$+$ $\frac{1}{2}\sum_{l\,k}\Delta I^{(1)}_{jl}\otimes{\Delta P}^{(0)}_{lk}\otimes{\Delta P}^{(0)}_{ki}(x)+\frac{1}{2}\sum_{l}{\Delta P}^{(0)}_{jl}\otimes{\Delta P}^{(1)}_{li}(x)$
$+$ ${\Delta P}^{(0)}_{ji}(x)\left(-\frac{1}{2}\beta_{1}-\frac{1}{2}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)+\delta_{ji}\delta(1-x)\left(\frac{1}{4}\beta_{1}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right$.
$+$ $\left.\frac{1}{2}\beta_{0}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})+\frac{1}{4}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)$
$+$ $\Delta I^{(1)}_{ji}(x)\left(\frac{3}{4}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})+\beta_{0}^{2}+\frac{1}{8}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}\right)$
$+$ ${\Delta P}^{(1)}_{ji}(x)\left(-\beta_{0}-\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)+\frac{1}{2}\sum_{l}{\Delta P}^{(1)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)\bigg]$
$+$ $L_{\perp}\bigg[-\sum_{l}\Delta I^{(1)}_{jl}\otimes{\Delta P}^{(1)}_{li}(x)-\sum_{l}\Delta I^{(2)}_{jl}\otimes{\Delta P}^{(0)}_{li}(x)-{\Delta P}^{(0)}_{ji}(x)\gamma^{R}_{1}L_{Q}-{\Delta P}^{(2)}_{ji}(x)$
$+$ $\delta_{ji}\delta(1-x)\biggl(2\beta_{0}\gamma_{1}^{R}L_{Q}+\frac{1}{2}\gamma_{1}^{R}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})L_{Q}+\frac{1}{2}(2\gamma_{2}^{B}-\Gamma^{\text{cusp}}_{2}L_{Q})\biggl)$
$+$ $\Delta I^{(1)}_{ji}(x)\left(\beta_{1}+\frac{1}{2}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)+\Delta I^{(2)}_{ji}(x)\left(2\beta_{0}+\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\bigg]$
$+$ $\delta_{ji}\delta(1-x)\gamma^{R}_{2}L_{Q}+\Delta I^{(1)}_{ji}(x)\gamma^{R}_{1}L_{Q}+\Delta I^{(3)}_{ji}(x)\,$, (100)
where $\Delta I^{(n)}_{ji}(z)$ are the scale-independent coefficient functions. In $\Delta\mathcal{I}_{ji}^{{(3)}}$ we have used $\gamma_{0}^{R}=0$ to simplify the expression. The scale logarithms are defined by
$L_{\perp}=\ln\frac{b_{T}^{2}\mu^{2}}{b_{0}^{2}},\quad L_{Q}=2\ln\frac{x\,P_{+}}{\nu},\quad L_{\nu}=\ln\frac{\nu^{2}}{\mu^{2}}\,,\quad b_{0}=2e^{-\gamma_{E}}\,.$ (101)
Similarly, the RG solution to the fragmentation coefficient functions are
$\Delta\mathcal{C}^{(0)}_{ji}(z,b_{\perp}/z,$ $\,\mu,\nu)=\delta_{ji}\delta(1-z)\,$,
$\Delta\mathcal{C}^{(1)}_{ji}(z,b_{\perp}/z,$ $\,\mu,\nu)=\left(-\frac{\Gamma^{\text{cusp}}_{0}}{2}L_{\perp}L_{Q}+\gamma_{0}^{B}L_{\perp}+\gamma_{0}^{R}L_{Q}\right)\delta_{ji}\delta(1-z)-{\Delta P}_{ji}^{T{(0)}}(z)L_{\perp}+\Delta C_{ji}^{(1)}(z)\,$,
$\Delta\mathcal{C}^{(2)}_{ji}(z,b_{\perp}/z,$ $\,\mu,\nu)=\bigg[\frac{1}{8}\left(-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}\right)\left(-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}+2\beta_{0}\right)L_{\perp}^{2}$
$+$ $\left((-\Gamma^{\text{cusp}}_{0}L_{Q}+2\gamma^{B}_{0}+2\beta_{0})\frac{\gamma_{0}^{R}}{2}L_{Q}-\frac{\Gamma^{\text{cusp}}_{1}}{2}L_{Q}+\gamma^{B}_{1}\right)L_{\perp}$
$+$ $\frac{(\gamma_{0}^{R})^{2}}{2}L_{Q}^{2}+\gamma_{1}^{R}L_{Q}\bigg]\,\delta_{ji}\delta(1-z)+\bigg(\frac{1}{2}\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes{\Delta P}^{T{(0)}}_{li}(z)$
$+$ $\frac{{\Delta P}^{T{(0)}}_{ji}(z)}{2}(\Gamma^{\text{cusp}}_{0}L_{Q}-2\gamma_{0}^{B}-\beta_{0})\bigg)L_{\perp}^{2}+\bigg[-{\Delta P}^{T{(1)}}_{ji}(z)-{\Delta P}^{T{(0)}}_{ji}(z)\gamma_{0}^{R}L_{Q}$
- $\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes\Delta C^{(1)}_{li}(z)+\left(-\frac{\Gamma^{\text{cusp}}_{0}}{2}L_{Q}+\gamma_{0}^{B}+\beta_{0}\right)\Delta C^{(1)}_{ji}(z)\bigg]L_{\perp}$
$+$ $\gamma_{0}^{R}L_{Q}\Delta C^{(1)}_{ji}(z)+\Delta C^{(2)}_{ji}(z)\,$,
$\Delta\mathcal{C}^{(3)}_{ji}(z,b_{\perp}/z,$ $\,\mu,\nu)=L_{\perp}^{3}\bigg[\left(\frac{1}{2}\beta_{0}+\frac{1}{4}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes{\Delta P}^{T{(0)}}_{li}(z)$
- $\frac{1}{6}\sum_{l\,k}{\Delta P}^{T{(0)}}_{jl}\otimes{\Delta P}^{T{(0)}}_{lk}\otimes{\Delta P}^{T{(0)}}_{ki}(z)+\delta_{ji}\delta(1-z)\left(\frac{1}{6}\beta_{0}^{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right$.
$+$ $\left.\frac{1}{8}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}+\frac{1}{48}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{3}\right)$
$+$ ${\Delta P}^{T{(0)}}_{ji}\left(-\frac{1}{2}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})-\frac{1}{3}\beta_{0}^{2}-\frac{1}{8}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}\right)\bigg]$
$+$ $L_{\perp}^{2}\bigg[\left(-\frac{3}{2}\beta_{0}-\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes\Delta C^{(1)}_{li}(z)$
$+$ $\frac{1}{2}\sum_{l\,k}{\Delta P}^{T{(0)}}_{jl}\otimes{\Delta P}^{T{(0)}}_{lk}\otimes\Delta C^{(1)}_{ki}(z)+\frac{1}{2}\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes{\Delta P}^{T{(1)}}_{li}(z)$
$+$ ${\Delta P}^{T{(0)}}_{ji}(z)\left(-\frac{1}{2}\beta_{1}-\frac{1}{2}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)+\delta_{ji}\delta(1-z)\left(\frac{1}{4}\beta_{1}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right$.
$+$ $\left.\frac{1}{2}\beta_{0}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})+\frac{1}{4}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)$
$+$ $\Delta C^{(1)}_{ji}(z)\left(\frac{3}{4}\beta_{0}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})+\beta_{0}^{2}+\frac{1}{8}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})^{2}\right)$
$+$ ${\Delta P}^{T{(1)}}_{ji}(z)\left(-\beta_{0}-\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)+\frac{1}{2}\sum_{l}{\Delta P}^{T{(1)}}_{jl}\otimes{\Delta P}^{T{(0)}}_{li}(z)\bigg]$
$+$ $L_{\perp}\bigg[-\sum_{l}{\Delta P}^{T{(1)}}_{jl}\otimes\Delta C^{(1)}_{li}(z)-\sum_{l}{\Delta P}^{T{(0)}}_{jl}\otimes\Delta C^{(2)}_{li}(z)-{\Delta P}^{T{(0)}}_{ji}(z)\gamma^{R}_{1}L_{Q}$
- ${\Delta P}^{T{(2)}}_{ji}(z)+\delta_{ji}\delta(1-z)\biggl(2\beta_{0}\gamma_{1}^{R}L_{Q}+\frac{1}{2}\gamma_{1}^{R}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})L_{Q}+\frac{1}{2}(2\gamma_{2}^{B}-\Gamma^{\text{cusp}}_{2}L_{Q})\biggl)$
$+$ $\Delta C^{(1)}_{ji}(z)\left(\beta_{1}+\frac{1}{2}(2\gamma^{B}_{1}-\Gamma^{\text{cusp}}_{1}L_{Q})\right)+\Delta C^{(2)}_{ji}(z)\left(2\beta_{0}+\frac{1}{2}(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q})\right)\bigg]$
$+$ $\delta_{ji}\delta(1-z)\gamma^{R}_{2}L_{Q}+\Delta C^{(1)}_{ji}(z)\gamma^{R}_{1}L_{Q}+\Delta C^{(3)}_{ji}(z)\,.$ (102)
The anomalous dimensions appeared above are identical to those in the space-like ones and we have suppressed their dependence on the exact flavor. The logarithms appeared in the fragmentation coefficient functions are defined as
$L_{\perp}=\ln\frac{b_{T}^{2}\mu^{2}}{b_{0}^{2}},\quad L_{Q}=2\ln\frac{P_{+}}{z\,\nu},\quad L_{\nu}=\ln\frac{\nu^{2}}{\mu^{2}}\,,\quad b_{0}=2e^{-\gamma_{E}}\,$, (103)
which differ from thsoe in Eq. (101) only in $L_{Q}$. Both space-like and time-like coefficient functions depend on the rapidity regulator being used. The physical and rapidity-regulator-independent coefficient functions can be obtained by multiplying the genuine collinear coefficient functions with the squared root of the TMD soft functions ${\cal S}(b_{\perp},\mu,\nu)$ Luo:2019hmp ; Luo:2019bmw
$\Delta\mathscr{I}_{ij}(x,b_{\perp},\mu)=$ $\Delta{\cal I}_{ij}(x,b_{\perp},\mu,\nu)\sqrt{{\cal S}_{i}(b_{\perp},\mu,\nu)}\,$,
$\Delta\mathscr{C}_{ij}(z,b_{\perp}/z,\mu)=$ $\Delta{\cal C}_{ij}(z,b_{\perp}/z,\mu,\nu)\sqrt{{\cal S}_{j}(b_{\perp},\mu,\nu)}\,.$ (104)
4.3 $\gamma_{5}$ in dimensional regularization and scheme transformations
The computation of polarized TMDs requires careful treatment of the genuinely four-dimensional objects $\gamma_{5}$ and the Levi-Civita tensor $\epsilon^{\mu\nu\rho\sigma}$, whose definitions must be consistently extended to $D=4-2\epsilon$ dimensions. However, the anti-commutativity of $\gamma_{5}$ and the cyclicity of the Dirac trace cannot be simultaneously preserved in dimensional regularization. One approach is to retain the four-dimensional Dirac algebra of $\gamma_{5}$ by evaluating the trace from a prescribed ‘reading point’ Korner:1991sx ; Kreimer:1993bh , thereby avoiding ambiguities associated with non-cyclic traces. This method is employed in Refs. Matiounine:1998re ; Rijken:1997rg to obtain the polarized DIS coefficient functions and DGLAP splitting functions in $\overline{\text{MS}}$ scheme. The “reading point” prescription is re-examined in the course of computing helicity TMDs, where we reproduce the two-loop scheme transformation factor $z_{\text{ps}}^{(2)}$, introduced below. Another approach is to abandon the anti-commutativity of $\gamma_{5}$, and instead read in $\gamma_{5}$ from the effective vertex. This leads to the use of the HVBM scheme tHooft:1972tcz ; Breitenlohner:1977hr , or the closely related Larin prescription Larin:1991tj ; Zijlstra:1992kj ; Larin:1993tq , for the consistent treatment of $\gamma_{5}$ in dimensional regularization. In HVBM scheme, one proceeds as Ravindran:2003gi
  1. 1.
    Evaluate integrals first
  2. a)
[…]
  3. b)
    Evaluate tensor-like Feynman integrals and phase space integrals in D-dimensions
  4. 2.
    Begin explicit definition of $\gamma_{5}$
  5. c)
    Replace the $\gamma_{5}$-matrix by
    $\gamma_{\mu}\gamma_{5}=\frac{i}{6}\epsilon_{\mu\rho\sigma\tau}\gamma^{\rho}\gamma^{\sigma}\gamma^{\tau}\quad\text{or}\quad\gamma_{5}=\frac{i}{24}\epsilon_{\mu\rho\sigma\tau}\gamma^{\mu}\gamma^{\rho}\gamma^{\sigma}\gamma^{\tau}$
  6. d)
    Compute trace of Dirac matrix in D dimensions
  7. e)
    Contract the Levi-Civita tensors in four dimensions
In $\texttt{Larin}^{+}$ scheme Gutierrez-Reyes:2017glx , one replaces the $\gamma_{5}$-matrix by
$\gamma^{+}\gamma_{5}\to\frac{i\epsilon_{\perp}^{\alpha\beta}}{2}\gamma_{\alpha}\gamma_{\beta}\,$, (105)
and supplement it by the D-dimensional relation
$\epsilon_{\perp}^{\alpha_{1}\beta_{1}}\epsilon_{\perp}^{\alpha_{2}\beta_{2}}=-g_{\perp}^{\alpha_{1}\alpha_{2}}g_{\perp}^{\beta_{1}\beta_{2}}+g_{\perp}^{\alpha_{1}\beta_{2}}g_{\perp}^{\beta_{1}\alpha_{2}}\,.$ (106)
Accordingly, the normalization factors presented in Eq. (66) are
$\mathcal{N}_{\text{sch.}}=\left\{\begin{array}[]{cc}1&\texttt{HVBM}\,,\\ (1-\epsilon)^{-1}(1-2\epsilon)^{-1}&\texttt{Larin}^{+}\,.\end{array}\right$. (108)
We have verified that the HVBM and $\texttt{Larin}^{+}$ prescriptions yield identical results for helicity TMDs at $\text{N}^{3}$LO. Both HVBM and $\texttt{Larin}^{+}$ schemes break the anticommutativity of the $\gamma_{5}$ matrix, leading to a renormalization of the non-singlet baryon current despite its conservation (number of quarks minus antiquarks of each flavor) Larin:1991tj , and causing a violation of the Adler–Bardeen theorem for the non-renormalization of the axial anomaly beyond one-loop Adler:1969er . To restore anticommutativity and the associated Ward identities, it is necessary to introduce additional evanescent counterterms to correct the axial current prior to collinear mass factorization. Alternatively, after mass factorization, the finite coefficient functions and the corresponding collinear PDFs (or FFs) are subjected to a factorization-scheme transformation to obtain the correct expression in $\overline{\text{MS}}$
$\Delta{\cal B}^{\text{sub.}}\equiv\frac{1}{Z^{B}}\frac{\Delta{\cal B}^{{\rm bare}}}{\mathcal{S}_{0\rm b}}=$ $\Delta\mathcal{I}_{L}\otimes\Delta\phi_{L}=\left(\Delta\mathcal{I}_{L}\otimes Z_{i}^{-1}\right)\otimes\left(Z_{i}\otimes\Delta\phi_{L}\right)=\Delta\mathcal{I}\otimes\Delta\phi\,$,
$\Delta{\cal D}^{{\text{sub.}}}\equiv\frac{1}{Z^{B}}\frac{\Delta{\cal D}^{{\rm bare}}}{\mathcal{S}_{0\rm b}}=$ $\Delta d_{L}\otimes\Delta\mathcal{C}_{L}=\left(\Delta d_{L}\otimes Z_{i}^{T}\right)\otimes\left((Z^{T}_{i})^{-1}\otimes\Delta\mathcal{C}_{L}\right)=\Delta d\otimes\Delta\mathcal{C}\,$, (109)
where above we introduce the UV and soft zero-bin subtracted functions $\Delta{\cal B}^{\text{sub.}}$ and $\Delta{\cal D}^{{\text{sub.}}}$, and suppress all the function arguments and matrix indices. The transformation factors are defined for each of the flavor group indices (see the flavor group decomposition in apendix A). For flavor group index $i=\pm,\text{v},\text{s}$, they allow perturbative expansions of the form in Mellin-$N$ space
$Z^{i}=1+\sum_{n}\left(\frac{\alpha_{s}}{4\pi}\right)^{n}z_{i}^{(n)}\,.$ (110)
We denote the difference of the splitting functions as $\delta[\Delta P]=\Delta P_{\overline{\text{MS}}}-\Delta P_{L}$, from DGLAP evolution equations we derive up to $\text{N}^{3}$LO Moch:2014sna :
$\delta[\Delta P]=$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\left\{\left[Z_{i}^{(1)},P_{L}^{(0)}\right]-\beta_{0}Z_{i}^{(1)}\right\}+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg\{\left[Z_{i}^{(1)},P_{L}^{(1)}\right]+\left[Z_{i}^{(2)},P_{L}^{(0)}\right]$
$+$ $\left[Z_{i}^{(1)},P_{L}^{(0)}\right]Z_{i}^{(1)}-\beta_{1}Z_{i}^{(1)}+\beta_{0}\left(\left(Z_{i}^{(1)}\right)^{2}-2Z_{i}^{(2)}\right)\bigg\}+\mathcal{O}(\alpha_{s}^{4})\,$,
$\delta[\Delta P^{T}]=$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\left\{\left[P_{L}^{T,(0)},Z_{i}^{T,(1)}\right]-\beta_{0}Z_{i}^{T,(1)}\right\}+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg\{\left[P_{L}^{T,(1)},Z_{i}^{T,(1)}\right]+\left[P_{L}^{T,(0)},Z_{i}^{T,(2)}\right]$
$+$ $Z_{i}^{T,(1)}\left[P_{L}^{T,(0)},Z_{i}^{T,(1)}\right]-\beta_{1}Z_{i}^{T,(1)}+\beta_{0}\left(\left(Z_{i}^{T,(1)}\right)^{2}-2Z_{i}^{T,(2)}\right)\bigg\}+\mathcal{O}(\alpha_{s}^{4})\,$, (111)
[…]
$\delta[\Delta P_{\text{ns}}]=$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\left\{-\beta_{0}z_{\text{ns}}^{(1)}\right\}+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg\{\beta_{0}\left(z_{\text{ns}}^{(1)}\right)^{2}-\beta_{1}z_{\text{ns}}^{(1)}-2\beta_{0}z_{\text{ns}}^{(2)}\bigg\}+\mathcal{O}(\alpha_{s}^{4})\,$,
$\delta[\Delta P_{\text{ns}}^{T}]=$ $\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\left\{-\beta_{0}z_{\text{ns}}^{T(1)}\right\}+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\bigg\{\beta_{0}\left(z_{\text{ns}}^{T(1)}\right)^{2}-\beta_{1}z_{\text{ns}}^{T(1)}-2\beta_{0}z_{\text{ns}}^{T(2)}\bigg\}+\mathcal{O}(\alpha_{s}^{4})\,.$ (112)
For the singlet sector with $i=\text{s}$, the transformation factors are $2\times 2$ matrices in Mellin-$N$ space, but only $q\to q$ entry is non-trivial. Indeed, we have
$(z_{\text{s}}^{(n)})_{ik}=\delta_{iq}\delta_{kq}z_{qq}^{(n)}=\delta_{iq}\delta_{kq}\left\{z^{(n)}_{\text{ns},+}+z^{(n)}_{\text{ps}}\right\}\,$, (113)
and
$Z^{\text{s}}_{ik}=\delta_{ik}+\delta_{iq}\delta_{kq}\left(\left(\frac{\alpha_{s}}{4\pi}\right)z^{(1)}_{\text{ns},+}+\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\left\{z^{(2)}_{\text{ns},+}+z^{(2)}_{\text{ps}}\right\}+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\left\{z^{(3)}_{\text{ns},+}+z^{(3)}_{\text{ps}}\right\}+\mathcal{O}(\alpha_{s}^{4})\right)\,.$ (114)
Substituting Eq. (114) into Eq. (4.3) , we find for the singlet space-like splitting functions
$\delta[\Delta P_{qq}^{\text{s}\,(1)}]=$ $-\beta_{0}z_{qq}^{(1)}\,,\quad\delta[\Delta P_{qg}^{\text{s}\,(1)}]=z_{qq}^{(1)}\otimes\Delta P^{\text{s}\,(0)}_{qg}$
$\delta[\Delta P_{gq}^{\text{s}\,(1)}]=$ $-\Delta P^{\text{s}\,(0)}_{gq}\otimes z_{qq}^{(1)}\,,\quad\delta[\Delta P_{gg}^{\text{s}\,(1)}]=0\,$, (115)
and
$\delta[\Delta P_{qq}^{\text{s}\,(2)}]=$ $\beta_{0}\left[\left(z_{qq}^{(1)}\right)^{2}-2z_{qq}^{(2)}\right]-\beta_{1}z_{qq}^{(1)}\,$,
$\delta[\Delta P_{qg}^{\text{s}\,(2)}]=$ $z_{qq}^{(2)}\otimes\Delta P^{\text{s}\,(0)}_{qg}+z_{qq}^{(1)}\otimes\Delta P^{\text{s}\,(1)}_{qg,L}\,$,
$\delta[\Delta P_{gq}^{\text{s}\,(2)}]=$ $-\left(\Delta P^{\text{s}\,(1)}_{gq,L}-\Delta P^{\text{s}\,(0)}_{gq}\otimes z_{qq}^{(1)}\right)\otimes z_{qq}^{(1)}-\Delta P^{\text{s}\,(0)}_{gq}\otimes z_{qq}^{(2)}\,$,
$\delta[\Delta P_{gg}^{\text{s}\,(2)}]=$ $0\,.$ (116)
For the singlet time-like splitting functions, we find
$\delta[\Delta P_{qq}^{T,\text{s}\,(1)}]=$ $-\beta_{0}z_{qq}^{T(1)}\,,\quad\delta[\Delta P_{gq}^{T,\text{s}\,(1)}]=z_{qq}^{T(1)}\otimes\Delta P^{T,\text{s}\,(0)}_{gq}$
$\delta[\Delta P_{qg}^{T,\text{s}\,(1)}]=$ $-\Delta P^{T,\text{s}\,(0)}_{qg}\otimes z_{qq}^{T(1)}\,,\quad\delta[\Delta P_{gg}^{T,\text{s}\,(1)}]=0\,$, (117)
and
$\delta[\Delta P_{qq}^{T,\text{s}\,(2)}]=$ $\beta_{0}\left[\left(z_{qq}^{T(1)}\right)^{2}-2z_{qq}^{T(2)}\right]-\beta_{1}z_{qq}^{T(1)}\,$,
$\delta[\Delta P_{gq}^{T,\text{s}\,(2)}]=$ $z_{qq}^{T(2)}\otimes\Delta P^{T,\text{s}\,(0)}_{gq}+z_{qq}^{T(1)}\otimes\Delta P^{T,\text{s}\,(1)}_{gq,L}\,$,
$\delta[\Delta P_{qg}^{T,\text{s}\,(2)}]=$ $-\left(\Delta P^{T,\text{s}\,(1)}_{qg,L}-\Delta P^{T,\text{s}\,(0)}_{qg}\otimes z_{qq}^{T(1)}\right)\otimes z_{qq}^{T(1)}-\Delta P^{T,\text{s}\,(0)}_{qg}\otimes z_{qq}^{T(2)}\,$,
$\delta[\Delta P_{gg}^{T,\text{s}\,(2)}]=$ $0\,.$ (118)
[…]
$Z^{\text{ns},\pm}\equiv\frac{\Delta\mathcal{I}^{(\mp)}_{\text{naive}}}{\Delta\mathcal{I}^{(\pm)}_{\texttt{HVBM}}}=1+\sum_{k}\left(\frac{\alpha_{s}}{4\pi}\right)^{k}z^{(k)}_{\text{ns},\pm}\,$. (119)
With this method we determine $Z^{\text{ns}}$ up to three loops, reproducing the known one- and two-loop results Moch:2014sna . The first moment of $Z^{\text{ns},+}$ yields
$Z^{\text{ns},+}|_{N=1}=$ $1-\left(\frac{\alpha_{s}}{4\pi}\right)\*4\,\*C_{F}+\left(\frac{\alpha_{s}}{4\pi}\right)^{2}\*\bigg[22\,\*C_{F}-\frac{107}{9}\>\*C_{F}\*C_{A}+\frac{2}{9}\>\*C_{F}\*N_{f}\bigg]+\left(\frac{\alpha_{s}}{4\pi}\right)^{3}\*\bigg[$
$C_{F}\*\bigg(-\frac{370}{3}+96\,\*\zeta_{3}\bigg)+C_{F}\*C_{A}\*\bigg(\>\frac{5834}{27}-160\,\*\zeta_{3}\bigg)+C_{F}\*C_{A}\*\bigg(-\frac{2147}{27}+56\,\*\zeta_{3}\bigg)$
$+$ $C_{F}\*N_{f}\*\bigg(-\frac{62}{27}-\frac{32}{3}\>\*\zeta_{3}\bigg)+C_{A}\*C_{F}\*N_{f}\*\bigg(\>\frac{356}{81}+\frac{32}{3}\>\*\zeta_{3}\bigg)+\frac{52}{81}\>\*C_{F}\*N_{f}\,\bigg]\,$, (120)
which is exactly the same finite renormalization constant for the non-singlet axial vector obtained originally by Larin Larin:1991tj , and hence serves as an independent check for our calculation. Furthermore, we find at NLO, the Gribov-Lipatov relation breaks down as was originally observed in vanNeerven:1998ma ; Blumlein:2000wh
$z^{T(2)}_{\text{ns},\pm}(\xi)=-\xi z^{(2)}_{\text{ns},\pm}(1/\xi)+\beta_{0}z^{(1)}_{\text{ns},\pm}(\xi)\ln\xi\,.$ (121)
For the pure-singlet case, two $\gamma_{5}$ matrices appear, one in each fermion trace, so that the loops are topologically unconnected. Following the procedure of ‘reading in’ the $\gamma_{5}$ matrix from the position suggested in Ref. Vogelsang:1996im —namely, by reading in the $\gamma_{5}$ inside the triangle diagrams (appearing in $g\to q$ channel at NLO and $q^{\prime}\to q$ at NNLO) from the gluon-quark vertex rather than from the projector vertex, and subtracting the collinear poles accordingly, we reproduce the NNLO pure-singlet transformation factor $z_{\text{ps}}^{(2)}$ Matiounine:1998re ; Rijken:1997rg as the difference in IR-subtracted coefficient functions between Kreimer’s prescription and the HVBM prescription. The N3LO pure-singlet transformation factor $z_{\text{ps}}^{(3)}$ remains unkown.
4.4 The NNLO helicity-dependent splitting functions
4.4.1 Space-like results in $\overline{\text{MS}}$}
The space-like helicity dependent splitting functions were obtained in Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , by comparison we find all agreement except for the term with cubic color structure $d_{abc}^{2}$ for ‘sea’ quark difference $N_{f}(\Delta P_{qq^{\prime}}-\Delta P_{q\bar{q}^{\prime}})$:  [^1] Our results agree at transcendental weight-4, but discrepancies appear for the weight-3 terms, indicating differences of $\gamma_{5}$ prescriptions between the one used in Ref Moch:2015usa and our tensor reduction prescription following Ravindran:2003gi .
$\Delta P_{d_{abc}^{2}}^{S,(2)}-\Delta P_{d_{abc}^{2}}^{\text{Moch:2015}}=\frac{16N_{f}}{3}\frac{d_{abc}^{2}}{N_{c}}\bigg[-18\zeta_{3}+3H_{0}+84H_{2}+36H_{3}+12H_{-2,0}+(1-z)\times\left(174H_{1}-12H_{1,0,0}-84\zeta_{2}\right)+(1+z)\times\left(36H_{-1,0}-48H_{-1,2}-24H_{-1,0,0}+48H_{-1}\zeta_{2}-36H_{0}\zeta_{2}\right)+z\times\left(-174H_{0}+90H_{2}+12H_{3}-24H_{-2,0}+54H_{0,0}+24H_{0,0,0}-54\zeta_{2}-72\zeta_{3}\right)\bigg]\,.$ (122)
It’s interesting to investigate the first moment of the helicity splitting functions, up to $\text{N}^{3}$LO we found
$\Delta P^{(n)}_{gg}(N=1)=$ $\beta_{n}\,$,
$\Delta P_{\text{ns},+}^{(n)}(N=1)=$ $\Delta P_{qg}^{(n)}(N=1)=0\,$,
$\Delta P_{\text{ps}}^{(n)}(N=1)=$ $-2N_{f}\Delta P_{gq}^{(n-1)}(N=1)\,.$ (123)
These relations follow from the Adler–Bardeen theorem–ensuring the non-renormalization of the non-singlet axial anomaly in pure QCD Adler:1969er –and from the one-loop exactness of the singlet axial anomaly in QCD Larin:1993tq :
$\int_{0}^{1}d\xi\,\xi^{n-1}[\Delta f_{j}(\xi)-(-1)^{n}\Delta f_{\bar{\jmath}}]=\frac{i^{n-1}}{2(P^{+})^{n}}\langle P|\bar{\psi}_{j}\gamma^{+}\gamma_{5}(\partial_{+})^{n-1}\psi_{j}|P\rangle_{\text{c}}\,$,
$\langle\psi(x)\not{\partial}\gamma_{5}\psi(x)\mathcal{O}(x_{1}\,,\dots\,,x_{n})\rangle=\frac{\alpha_{s}}{8\pi}\langle G^{a}_{\mu\nu}\widetilde{G}^{a\mu\nu}(x)\mathcal{O}(x_{1}\,,\dots\,,x_{n})\rangle\,.$ (124)
4.4.2 Time-like results in $\overline{\text{MS}}$}
The NLO time-like helicity-dependent splitting functions were obtained in Refs. Stratmann:1996hn ; Rijken:1997rg . Our results extend the precision to NNLO for the first time, reproducing the known NLO expressions. We now present the analytic NNLO results for both non-singlet and singlet sectors. The $\pm$’ non-singlet results coincide with the corresponding unpolarized ones $\mp$
$\Delta P_{\text{ns},\pm}^{T,(n)}=P_{\text{ns},\mp}^{T,(n)}\,.$ (125)
The results for the singlet sectors are
$\Delta P_{qq}^{T,\text{s}(0)}=4C_{F}\left[\frac{1}{1-z}\right]_{+}+3C_{F}\delta(1-z)+(-2z-2)C_{F}\,$, (126)
$\Delta P_{qq}^{T,\text{s}(1)}=\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\frac{268}{9}-8\zeta_{2}\right)C_{A}C_{F}-\frac{40C_{F}N_{f}}{9}\bigg]+C_{A}C_{F}\bigg[\frac{z^{2}}{z+1}\left(-8H_{-1,0}-\frac{446}{9}\right)+\left(-\frac{8z^{2}}{(z-1)(z+1)}\right)H_{0,0}+\frac{1}{z+1}\left(\frac{178}{9}-8H_{-1,0}\right)+\left(-\frac{8}{(z-1)(z+1)}\right)H_{0,0}+\frac{2z^{2}}{3(z-1)}H_{0}+\left(-\frac{46}{3(z-1)}\right)H_{0}+\frac{z\left(8\zeta_{2}-\frac{268}{9}\right)}{z+1}\bigg]+C_{F}N_{f}\bigg[z\left(8H_{0,0}+\frac{440}{9}\right)+8H_{0,0}+\left(-\frac{80z^{2}}{3(z-1)}\right)H_{0}+\frac{40}{3(z-1)}H_{0}+\frac{16z}{z-1}H_{0}-\frac{400}{9}\bigg]+C_{F}^{2}\bigg[\frac{12z^{3}}{(z-1)(z+1)}H_{0,0}+\frac{z^{2}}{z+1}\left(16H_{-1,0}+8\zeta_{2}+36\right)+\frac{z^{2}}{z-1}\left(8H_{0,1}+8H_{1,0}-20H_{0}\right)+\frac{28z^{2}}{(z-1)(z+1)}H_{0,0}+\frac{1}{z+1}\left(16H_{-1,0}+8\zeta_{2}-36\right)+\frac{1}{z-1}\left(8H_{0,1}+8H_{1,0}+16H_{0}\right)+\frac{20}{(z-1)(z+1)}H_{0,0}+\frac{4z}{(z-1)(z+1)}H_{0,0}+\left(-\frac{8z}{z-1}\right)H_{0}\bigg]+\delta(1-z)\bigg[\left(\frac{44}{3}\zeta_{2}-12\zeta_{3}+\frac{17}{6}\right)C_{A}C_{F}+\left(\left(-\frac{8}{3}\right)\zeta_{2}-\frac{1}{3}\right)C_{F}N_{f}+\left(-12\zeta_{2}+24\zeta_{3}+\frac{3}{2}\right)C_{F}^{2}\bigg]\,$, (127)
$\Delta P_{qq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\frac{1}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}-56H_{0,1,0}+96H_{1,0,0}-208H_{0,0,0,1}-64H_{0,0,1,1}-48H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z}{z-1}\left(-192\zeta_{2}H_{1}-800H_{1}+64H_{0,1,0}\right)+\frac{z}{z-1}\left(-144\zeta_{2}H_{0}-72H_{0}-12\zeta_{2}-192\zeta_{3}-324\zeta_{4}-352\zeta_{2}H_{0,-1}+144\zeta_{2}H_{0,0}+176H_{0,0}-96H_{0,0,0}-64H_{0,-1,-1,0}+128H_{0,-1,0,0}-96H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(-16\zeta_{2}H_{0}+128\zeta_{3}H_{0}+10H_{0}+204\zeta_{2}+48\zeta_{3}+252\zeta_{4}+160\zeta_{2}H_{0,-1}+208\zeta_{2}H_{0,0}+172H_{0,0}+528H_{0,0,0}-320H_{0,-1,-1,0}+192H_{0,-1,0,0}+288H_{0,0,-1,0}-592H_{0,0,0,0}-192H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}+88H_{0,1,0}+96H_{1,0,0}-240H_{0,0,0,1}-64H_{0,0,1,1}-144H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z^{3}}{z-1}\left(432\zeta_{2}H_{0}+114H_{0}+12\zeta_{2}+336\zeta_{3}-28\zeta_{4}-352\zeta_{2}H_{0,-1}+48\zeta_{2}H_{0,0}-228H_{0,0}+240H_{0,0,0}-64H_{0,-1,-1,0}-160H_{0,0,-1,0}-112H_{0,0,0,0}-128H_{0,0,1,0}+\frac{1}{z+1}+302\right)+\frac{1}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-12H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-64H_{0,-1,0}+64H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(16\zeta_{2}H_{0}-64\zeta_{3}H_{0}-64H_{0}-204\zeta_{2}-192\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}-224H_{0,0}-96H_{0,0,0}-320H_{0,-1,-1,0}+64H_{0,-1,0,0}+96H_{0,0,-1,0}-176H_{0,0,0,0}-128H_{0,0,1,0}+302\right)+\frac{z}{z+1}\left(-576H_{-1}\zeta_{2}-192H_{-1,0}-312H_{0,1}-64H_{-1,0,0}+768H_{-1,0,1}-320H_{0,-1,0}-96H_{0,0,1}\right)+\frac{z^{2}}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-300H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-160H_{0,-1,0}-256H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+144H_{1,0}+192H_{-1,-1,0}+z\left(192H_{-1,-1,0}-144H_{1,0}\right)\bigg]+C_{F}^{2}N_{f}\bigg[\frac{668}{3}H_{1}+\frac{1}{z-1}\left(\left(-\frac{2104}{9}\right)H_{0,1}+\left(-\frac{464}{3}\right)H_{0,1,0}+\left(-\frac{580}{9}\right)H_{1,0}-160\zeta_{2}H_{0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{5048}{9}\right)H_{0,0}+\left(-\frac{496}{3}\right)H_{0,0,1}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{4772}{9}\zeta_{2}+\frac{7868}{9}H_{0}+\left(-\frac{112\zeta_{2}}{3}\right)H_{0}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-152H_{0,0,0}+\frac{1}{z+1}-\frac{1727}{3}\right)+\frac{z}{z-1}\left(\frac{220}{3}H_{0,1}+\frac{280}{3}H_{1,0}+144H_{0,1,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{580}{9}\right)H_{1,0}+\left(-\frac{32}{3}\right)H_{0,1,0}+\frac{1124}{9}H_{0,1}+160\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{80}{3}\right)\zeta_{3}+\frac{20}{9}H_{0}+\frac{56}{3}H_{0,0,0}+\frac{760}{9}H_{0,0}+\frac{1672}{9}\zeta_{2}-64H_{0}\zeta_{2}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+64H_{0,0,1}+\frac{1}{z+1}-539\right)$ (128)
$\frac{z^{3}}{z-1}\left(\left(-\frac{7664}{9}\right)H_{0}+\left(-\frac{5092}{9}\right)\zeta_{2}+\frac{112}{3}\zeta_{3}+\frac{368}{3}H_{0,0,1}+\frac{1624}{3}H_{0,0}+\frac{176\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+120H_{0,0,0}+\frac{1}{z+1}+\frac{1727}{3}\right)+\frac{1}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{1576}{9}\right)H_{0,0}+\left(-\frac{1352}{9}\right)\zeta_{2}+\left(-\frac{280}{3}\right)H_{0,0,0}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{184}{9}H_{0}+64H_{0}\zeta_{2}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-64H_{0,0,1}+539\right)+\left(-\frac{128z}{3(z+1)}\right)H_{-1,0}+\frac{z^{2}}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+400H_{1}\zeta_{2}-52H_{1,1}+48H_{0,1,1}-440H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}+z\left(\left(-\frac{668}{3}\right)H_{1}-400H_{1}\zeta_{2}+52H_{1,1}-32H_{0,1,1}+440H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)-32H_{0,1,1,1}\bigg]+C_{A}^{2}C_{F}\bigg[\frac{1}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+36H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1184}{3}\right)H_{1}-64H_{1}\zeta_{2}-64H_{1,0,0}+32H_{0,-1,-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1744}{9}\right)\zeta_{2}+\left(-\frac{722}{9}\right)H_{0,0}+\left(-\frac{1258}{27}\right)H_{0}+\frac{160}{3}\zeta_{3}+\frac{344}{3}H_{0,-1,0}-100H_{0}\zeta_{2}+7\zeta_{4}+60\zeta_{2}H_{0,0}+100H_{0,0,1}+72H_{0,-1,0,0}+96H_{0,-1,0,1}-8H_{0,0,-1,0}-60H_{0,0,0,1}+\frac{1}{z+1}+\frac{5333}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{416}{3}\right)H_{0,0,0}+\left(-\frac{1378}{27}\right)H_{0}+\left(-\frac{4}{3}\right)H_{0,0,1}+\frac{328}{3}H_{0,-1,0}+\frac{1504}{9}\zeta_{2}+\frac{332\zeta_{2}}{3}H_{0}-144H_{0}\zeta_{3}+112\zeta_{3}-121\zeta_{4}-68\zeta_{2}H_{0,0}-130H_{0,0}-56H_{0,-1,0,0}-160H_{0,-1,0,1}-40H_{0,0,-1,0}+40H_{0,0,0,0}+28H_{0,0,0,1}+\frac{1}{z+1}+\frac{923}{27}\right)+\frac{z^{2}}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+28H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{376}{3}\right)H_{0,-1,0}+\left(-\frac{2266}{27}\right)H_{0}+\frac{238}{9}H_{0,0}+\frac{224}{3}\zeta_{2}+188H_{0}\zeta_{2}+48\zeta_{3}-31\zeta_{4}+36\zeta_{2}H_{0,0}-188H_{0,0,1}+88H_{0,-1,0,0}+96H_{0,-1,0,1}+8H_{0,0,-1,0}-36H_{0,0,0,1}+\frac{1}{z+1}-\frac{5333}{27}\right)+\frac{1}{z+1}\left(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-120\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{560}{3}\right)\zeta_{3}+\left(-\frac{3442}{27}\right)H_{0}+\left(-\frac{80}{3}\right)H_{0,0,0}+\left(-\frac{8}{3}\right)H_{0,-1,0}+\frac{92}{3}H_{0,0,1}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}-112H_{0}\zeta_{3}-63\zeta_{4}-28\zeta_{2}H_{0,0}-162H_{0,0}-40H_{0,-1,0,0}-160H_{0,-1,0,1}-24H_{0,0,-1,0}+24H_{0,0,0,0}+4H_{0,0,0,1}-\frac{923}{27}\right)+\frac{z}{z+1}\left(\left(-\frac{448}{3}\right)H_{0,1}+\left(-\frac{160}{3}\right)H_{-1,0}-256H_{-1}\zeta_{2}-16\zeta_{2}H_{0,1}+96H_{-1,0,0}+320H_{-1,0,1}\right)$ (129)
$\frac{z^{2}}{z+1}\left(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-136\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\right)+64zH_{-1,-1,0}+64H_{-1,-1,0}\bigg]+C_{F}N_{f}^{2}\bigg[\left(-\frac{416}{9}\right)H_{1}+\left(-\frac{152}{9}\right)H_{0,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,1}+\frac{152}{9}\zeta_{2}+\frac{64}{3}\zeta_{3}+\frac{1}{z-1}\left(\frac{8}{3}H_{0,0}+\frac{112}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}\right)+\frac{z}{z-1}\left(\left(-\frac{64}{3}\right)H_{0,0}+\frac{320}{9}H_{0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{128}{3}\right)H_{0}+\frac{152}{9}H_{0,0}+\left(-\frac{16\zeta_{2}}{3}\right)H_{0}\right)+40H_{1,0}+16H_{0,1,0}+z\left(\left(-\frac{296}{9}\right)H_{0,1}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{64}{3}\zeta_{3}+\frac{296}{9}\zeta_{2}+\frac{416}{9}H_{1}-40H_{1,0}+16H_{0,1,0}+\frac{1952}{27}\right)-\frac{1936}{27}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[C_{F}\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{2}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)C_{F}N_{f}C_{A}-\frac{16}{27}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)\bigg]+\delta(1-z)\bigg[\left(-32\zeta_{3}\zeta_{2}+18\zeta_{2}+68\zeta_{3}+144\zeta_{4}-240\zeta_{5}+\frac{29}{2}\right)C_{F}^{3}+\left(\left(-\frac{136}{3}\right)\zeta_{3}+\frac{20}{3}\zeta_{2}+\frac{116}{3}\zeta_{4}-23\right)N_{f}C_{F}^{2}+\left(\left(-\frac{16}{9}\right)\zeta_{3}+\frac{80}{27}\zeta_{2}-\frac{17}{9}\right)N_{f}^{2}C_{F}+C_{A}^{2}\left(\left(-\frac{1552}{9}\right)\zeta_{3}+\frac{4496}{27}\zeta_{2}-5\zeta_{4}+40\zeta_{5}-\frac{1657}{36}\right)C_{F}+C_{A}\left(\left(\left(-\frac{494}{3}\right)\zeta_{4}+\left(-\frac{410}{3}\right)\zeta_{2}+\frac{844}{3}\zeta_{3}+16\zeta_{2}\zeta_{3}+120\zeta_{5}+\frac{151}{4}\right)C_{F}^{2}+N_{f}\left(\left(-\frac{1336}{27}\right)\zeta_{2}+\frac{200}{9}\zeta_{3}+2\zeta_{4}+20\right)C_{F}\right)\bigg]+C_{A}C_{F}^{2}\bigg[\frac{1}{z-1}\left(\left(-\frac{2008}{3}\right)H_{1}+\frac{272}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-80H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{832}{3}\right)H_{1,0}+\frac{4016}{3}H_{1}+224H_{1}\zeta_{2}+128H_{1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1072}{3}\right)H_{0,-1,0}+\left(-\frac{2648}{9}\right)H_{0,1}+\left(-\frac{536}{3}\right)H_{0,0,1}+\left(-\frac{832}{9}\right)H_{0,0}+\frac{56}{3}\zeta_{3}+\frac{244}{9}H_{0}+\frac{2296}{9}\zeta_{2}+\frac{904\zeta_{2}}{3}H_{0}+268\zeta_{4}+416\zeta_{2}H_{0,-1}-192\zeta_{2}H_{0,0}+64\zeta_{2}H_{0,1}+140H_{0,0,0}+128H_{0,-1,-1,0}-208H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+192H_{0,0,0,1}+\frac{1}{z+1}+\frac{1516}{3}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3880}{9}\right)\zeta_{2}+\left(-\frac{832}{3}\right)\zeta_{3}+\left(-\frac{416}{3}\right)H_{0,-1,0}+\left(-\frac{350}{9}\right)H_{0}+\frac{1280}{9}H_{0,0}+\frac{652}{3}H_{0,0,0}+\frac{2296}{9}H_{0,1}-184H_{0}\zeta_{2}+224H_{0}\zeta_{3}+28\zeta_{4}-352\zeta_{2}H_{0,-1}+32\zeta_{2}H_{0,0}-128\zeta_{2}H_{0,1}+40H_{0,0,1}+256H_{0,-1,-1,0}+16H_{0,-1,0,0}+480H_{0,-1,0,1}-64H_{0,0,-1,0}+160H_{0,0,0,0}-32H_{0,0,0,1}+32H_{0,0,1,0}+\frac{1}{z+1}+\frac{1516}{3}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{2008}{3}\right)H_{1}+\frac{80}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-16H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{880}{3}\right)\zeta_{3}+\left(-\frac{1798}{9}\right)H_{0}+\left(-\frac{152}{9}\right)\zeta_{2}+\frac{776}{3}H_{0,0}+\frac{992}{3}H_{0,-1,0}+\frac{4792}{9}H_{0,1}+\frac{1672}{3}H_{0,0,1}+\left(-\frac{1688\zeta_{2}}{3}\right)H_{0}-12\zeta_{4}+448\zeta_{2}H_{0,-1}-96\zeta_{2}H_{0,0}+96\zeta_{2}H_{0,1}+36H_{0,0,0}+192H_{0,-1,-1,0}-176H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+96H_{0,0,0,1}+\frac{1}{z+1}-\frac{1516}{3}\right)$ (130)
$\frac{1}{z+1}\left(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{80}{3}\right)H_{0,-1,0}+\left(-\frac{152}{9}\right)H_{0,1}+\frac{92}{9}H_{0}+\frac{436}{3}H_{0,0,0}+\frac{1736}{9}\zeta_{2}+\frac{1496}{3}\zeta_{3}+\frac{4504}{9}H_{0,0}+88H_{0}\zeta_{2}+256H_{0}\zeta_{3}+268\zeta_{4}-320\zeta_{2}H_{0,-1}-160\zeta_{2}H_{0,1}+8H_{0,0,1}+320H_{0,-1,-1,0}+48H_{0,-1,0,0}+480H_{0,-1,0,1}+32H_{0,0,0,0}+64H_{0,0,0,1}+32H_{0,0,1,0}-\frac{1516}{3}\right)+\frac{z}{z+1}\left(\frac{608}{3}H_{-1,0}+800H_{-1}\zeta_{2}-160H_{-1,0,0}-1024H_{-1,0,1}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\right)-224zH_{-1,-1,0}-224H_{-1,-1,0}\bigg]+N_{f}C_{A}C_{F}\bigg[\left(-\frac{1960}{9}\right)H_{1}+\left(-\frac{844}{9}\right)H_{0,1}+\left(-\frac{88}{3}\right)H_{0,1,1}+\frac{260}{3}H_{1,1}+\left(-\frac{356}{z-1}\right)H_{1,0,0}+\frac{z}{z-1}\left(\left(-\frac{544}{3}\right)H_{0}+\left(-\frac{164}{3}\right)H_{0,0,1}+\frac{472}{3}H_{0,0,0}+\frac{4700}{9}H_{0,0}+\left(-\frac{220\zeta_{2}}{3}\right)H_{0}-248\zeta_{2}+160H_{0}\zeta_{3}-168\zeta_{3}+426\zeta_{4}+8\zeta_{2}H_{0,0}+\frac{1}{z+1}-\frac{2950}{9}\right)+\frac{696z}{z-1}H_{1,0,0}+\left(-\frac{356z^{2}}{z-1}\right)H_{1,0,0}+\frac{z^{2}}{z-1}\left(\left(-\frac{440}{3}\right)H_{0,0,0}+\left(-\frac{1664}{27}\right)H_{0}+\frac{448}{9}H_{0,0}+\frac{332}{3}\zeta_{2}+\frac{656}{3}\zeta_{3}+\frac{548\zeta_{2}}{3}H_{0}-32H_{0}\zeta_{3}-406\zeta_{4}+104\zeta_{2}H_{0,0}-52H_{0,0,1}+\frac{1}{z+1}-\frac{8014}{27}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{1508}{3}\right)H_{0,0}+\left(-\frac{472}{3}\right)H_{0,0,0}+\frac{212}{3}H_{0,0,1}+\frac{488}{3}\zeta_{3}+\frac{6232}{27}H_{0}+\frac{2392}{9}\zeta_{2}+\frac{172\zeta_{2}}{3}H_{0}-160H_{0}\zeta_{3}-426\zeta_{4}-8\zeta_{2}H_{0,0}+\frac{1}{z+1}+\frac{2950}{9}\right)+\frac{1}{z+1}\left(\left(-\frac{128}{3}\right)H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{280}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{1156}{9}\right)\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{140}{3}H_{0,0,1}+\frac{1000}{9}H_{0}-188H_{0}\zeta_{2}+32H_{0}\zeta_{3}-192\zeta_{3}+406\zeta_{4}-104\zeta_{2}H_{0,0}+168H_{0,0,0}+\frac{8014}{27}\right)+\frac{z}{z+1}\left(\frac{112}{3}H_{-1,0}-16H_{-1}\zeta_{2}+368H_{-1,0,0}-64H_{-1,0,1}+240H_{0,-1,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{128}{3}\right)H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{472}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\right)-360H_{1}\zeta_{2}-16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}+268H_{1,0}-80H_{-1,-1,0}-64H_{0,1,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-32H_{0,-1,-1,0}+80H_{0,-1,0,0}+112H_{0,0,-1,0}-304H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+z\left(\left(-\frac{2536}{9}\right)H_{0,1}+\left(-\frac{260}{3}\right)H_{1,1}+\frac{152}{3}H_{0,1,1}+\frac{1960}{9}H_{1}+360H_{1}\zeta_{2}+16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}-268H_{1,0}-80H_{-1,-1,0}+216H_{0,1,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}+32H_{0,-1,-1,0}-80H_{0,-1,0,0}-48H_{0,0,-1,0}+64H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}\right)\bigg]\,.$ (131)
$\Delta P_{gq}^{T,\text{s}(0)}=(4-2z)C_{F}\,$, (132)
$\Delta P_{gq}^{T,\text{s}(1)}=C_{A}C_{F}\left(z\left(8H_{-1,0}-8H_{0,0}-8H_{0,1}+24H_{1,0}+8H_{1,1}+44H_{0}+16H_{1}+8\zeta_{2}-16\right)+16H_{-1,0}-48H_{0,0}+16H_{0,1}-48H_{1,0}-16H_{1,1}-16H_{0}-16H_{1}+20\right)+C_{F}^{2}\left(z\left(4H_{0,0}+16H_{0,1}-16H_{1,0}-8H_{1,1}-18H_{0}-16H_{1}-28\right)-8H_{0,0}-32H_{0,1}+32H_{1,0}+16H_{1,1}+32H_{0}+16H_{1}+38\right)\,$, (133)
$\Delta P_{gq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\left(-\frac{48}{z}\right)H_{-1,0}+447H_{0}+626H_{1}+32H_{-1}\zeta_{2}+24H_{0}\zeta_{2}-112H_{1}\zeta_{2}-424\zeta_{2}-400H_{0}\zeta_{3}+64H_{1}\zeta_{3}-224\zeta_{3}+84\zeta_{4}+48H_{-1,0}+192\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+174H_{0,0}+400\zeta_{2}H_{0,1}+40H_{0,1}+128\zeta_{2}H_{1,0}+304H_{1,0}+96\zeta_{2}H_{1,1}+240H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-168H_{0,0,0}-344H_{0,0,1}+40H_{0,1,0}-40H_{0,1,1}+80H_{1,0,0}-16H_{1,0,1}+80H_{1,1,0}+96H_{1,1,1}-128H_{-1,-1,0,0}-64H_{-1,0,-1,0}+32H_{-1,0,0,0}+384H_{0,-1,-1,0}-128H_{0,-1,0,0}-192H_{0,0,-1,0}+192H_{0,0,0,0}+128H_{0,0,0,1}+176H_{0,0,1,0}+240H_{0,0,1,1}+80H_{0,1,0,0}-144H_{0,1,0,1}-80H_{0,1,1,0}-208H_{0,1,1,1}-64H_{1,0,-1,0}-192H_{1,0,0,0}-256H_{1,0,0,1}+32H_{1,0,1,0}-128H_{1,0,1,1}-192H_{1,1,0,0}-96H_{1,1,0,1}+32H_{1,1,1,0}+z\left(-60\zeta_{2}H_{0}-40\zeta_{3}H_{0}-221H_{0}-592H_{1}+32H_{-1}\zeta_{2}+76H_{1}\zeta_{2}+44\zeta_{2}-32H_{1}\zeta_{3}+76\zeta_{3}-194\zeta_{4}+64H_{-1,0}+64\zeta_{2}H_{0,-1}-32\zeta_{2}H_{0,0}-73H_{0,0}-232\zeta_{2}H_{0,1}+244H_{0,1}-64\zeta_{2}H_{1,0}-248H_{1,0}-48\zeta_{2}H_{1,1}-184H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-144H_{0,-1,0}+108H_{0,0,0}+140H_{0,0,1}+56H_{0,1,0}+112H_{0,1,1}+28H_{1,0,0}+52H_{1,0,1}-44H_{1,1,0}-84H_{1,1,1}-64H_{-1,-1,0,0}-32H_{-1,0,-1,0}+16H_{-1,0,0,0}+128H_{0,-1,-1,0}-32H_{0,-1,0,0}-32H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,0,1}-88H_{0,0,1,0}-120H_{0,0,1,1}-8H_{0,1,0,0}+72H_{0,1,0,1}+40H_{0,1,1,0}+104H_{0,1,1,1}+32H_{1,0,-1,0}+96H_{1,0,0,0}+128H_{1,0,0,1}-16H_{1,0,1,0}+64H_{1,0,1,1}+96H_{1,1,0,0}+48H_{1,1,0,1}-16H_{1,1,1,0}-32H_{1,1,1,1}-\frac{3133}{4}\right)+64H_{1,1,1,1}+\frac{1569}{2}\bigg]+N_{f}C_{F}^{2}\bigg[\left(-\frac{80}{3}\right)H_{1,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{112}{9}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,1,1}+\left(-\frac{16}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{1,0,1}+\left(-\frac{16}{3}\right)H_{1,1,0}+\frac{76}{27}H_{1}+\frac{16}{3}H_{1,1,1}+\frac{64}{3}H_{0,0,1}+\frac{224}{9}H_{0,1}+\frac{380}{3}H_{0,0}+\frac{4334}{27}H_{0}-16H_{1}\zeta_{2}-64H_{0,0,0,0}+z\left(\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{2038}{27}\right)H_{0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{64}{9}\right)H_{0,1}+\left(-\frac{8}{3}\right)H_{1,1,1}+\frac{8}{3}H_{1,0,1}+\frac{8}{3}H_{1,1,0}+\frac{32}{9}H_{1,1}+\frac{112}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,0,0}+8H_{1}\zeta_{2}+154H_{0,0}+32H_{0,0,0,0}-\frac{1567}{18}\right)+\frac{767}{9}\bigg]+C_{A}^{2}C_{F}\bigg[\left(-\frac{7520}{9}\right)\zeta_{2}+\left(-\frac{7376}{9}\right)H_{1,0}+\left(-\frac{1384}{3}\right)H_{-1,0,0}+\left(-\frac{904}{3}\right)H_{0,-1,0}+\left(-\frac{2248}{9}\right)H_{1,1}+\left(-\frac{2216}{9}\right)H_{-1,0}+\left(-\frac{268}{3}\right)H_{1,0,0}+\frac{8}{3}H_{0,1,1}+\frac{88}{3}H_{1,1,1}+\frac{112}{3}H_{1,1,0}+\frac{932}{3}H_{0,0,1}+\frac{5584}{9}H_{0,0}+\frac{2216}{3}H_{0,0,0}+\frac{7624}{9}H_{0,1}+\frac{28256}{27}H_{0}+\frac{32362}{27}H_{1}+\left(-\frac{24}{z}\right)H_{-1,0}+\left(-\frac{1468\zeta_{2}}{3}\right)H_{0}+104H_{-1}\zeta_{2}+264H_{1}\zeta_{2}-16H_{-1}\zeta_{3}+480H_{0}\zeta_{3}+656H_{1}\zeta_{3}-712\zeta_{3}+1570\zeta_{4}-32\zeta_{2}H_{-1,-1}$ (134)
$160\zeta_{2}H_{-1,0}-16\zeta_{2}H_{0,-1}-312\zeta_{2}H_{0,0}+496\zeta_{2}H_{0,1}-128\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}+288H_{0,1,0}-96H_{1,0,1}-192H_{-1,-1,-1,0}+192H_{-1,-1,0,0}-64H_{-1,-1,0,1}+352H_{-1,0,-1,0}-336H_{-1,0,0,0}+64H_{-1,0,0,1}-192H_{-1,0,1,0}-64H_{-1,0,1,1}+96H_{0,-1,-1,0}-272H_{0,-1,0,0}+64H_{0,-1,0,1}-496H_{0,0,-1,0}+1264H_{0,0,0,0}-184H_{0,0,0,1}+960H_{0,0,1,0}+320H_{0,0,1,1}-152H_{0,1,0,0}-352H_{0,1,0,1}-224H_{0,1,1,0}-160H_{0,1,1,1}-96H_{1,0,-1,0}+560H_{1,0,0,0}+160H_{1,0,0,1}+640H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}+128H_{1,1,0,1}+192H_{1,1,1,0}+z\left(\left(-\frac{31412}{27}\right)H_{1}+\left(-\frac{20644}{27}\right)H_{0}+\left(-\frac{1472}{3}\right)H_{-1,0,0}+\left(-\frac{992}{3}\right)H_{0,-1,0}+\left(-\frac{2752}{9}\right)H_{-1,0}+\left(-\frac{580}{3}\right)H_{0,0,1}+\left(-\frac{484}{3}\right)H_{0,1,1}+\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{200}{3}\right)H_{1,1,0}+\left(-\frac{172}{3}\right)H_{1,0,0}+\left(-\frac{44}{3}\right)H_{1,1,1}+\frac{1712}{9}H_{1,1}+\frac{5380}{9}H_{0,1}+\frac{5944}{9}H_{1,0}+\frac{10172}{9}H_{0,0}+104H_{-1}\zeta_{2}-156H_{0}\zeta_{2}-352H_{1}\zeta_{2}-796\zeta_{2}-8H_{-1}\zeta_{3}+176H_{0}\zeta_{3}-328H_{1}\zeta_{3}-464\zeta_{3}+143\zeta_{4}-16\zeta_{2}H_{-1,-1}+80\zeta_{2}H_{-1,0}-72\zeta_{2}H_{0,-1}+76\zeta_{2}H_{0,0}+24\zeta_{2}H_{0,1}+64\zeta_{2}H_{1,0}-48\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}-848H_{0,1,0}+96H_{1,0,1}-96H_{-1,-1,-1,0}+96H_{-1,-1,0,0}-32H_{-1,-1,0,1}+176H_{-1,0,-1,0}-168H_{-1,0,0,0}+32H_{-1,0,0,1}-96H_{-1,0,1,0}-32H_{-1,0,1,1}-80H_{0,-1,-1,0}+120H_{0,-1,0,0}+32H_{0,-1,0,1}+56H_{0,0,-1,0}-64H_{0,0,0,0}-76H_{0,0,0,1}+192H_{0,0,1,0}+64H_{0,0,1,1}+116H_{0,1,0,0}+32H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+48H_{1,0,-1,0}-280H_{1,0,0,0}-80H_{1,0,0,1}-320H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}-64H_{1,1,0,1}-96H_{1,1,1,0}-32H_{1,1,1,1}-\frac{13346}{9}\right)+64H_{1,1,1,1}+\frac{13840}{9}\bigg]+C_{A}C_{F}^{2}\bigg[\left(-\frac{52072}{27}\right)H_{1}+\left(-\frac{10868}{9}\right)H_{0,1}+\left(-\frac{1466}{3}\right)H_{0,0}+\left(-\frac{1136}{3}\right)H_{0,0,0}+\left(-\frac{3719}{27}\right)H_{0}+\left(-\frac{376}{3}\right)H_{1,1,1}+\left(-\frac{368}{3}\right)H_{1,1,0}+\left(-\frac{184}{3}\right)H_{0,0,1}+\left(-\frac{344}{9}\right)H_{1,1}+\left(-\frac{88}{3}\right)H_{1,0}+\frac{56}{3}H_{0,1,1}+\frac{272}{3}H_{1,0,0}+\frac{352}{3}H_{1,0,1}+\frac{1}{z}\left(-72H_{-1}\zeta_{2}-24H_{1}\zeta_{2}+72H_{-1,0}-48H_{-1,-1,0}+48H_{-1,0,0}+48H_{-1,0,1}\right)-256H_{-1}\zeta_{2}+296H_{0}\zeta_{2}-240H_{1}\zeta_{2}+716\zeta_{2}-176H_{-1}\zeta_{3}-752H_{0}\zeta_{3}-816H_{1}\zeta_{3}+728\zeta_{3}-2144\zeta_{4}+288\zeta_{2}H_{-1,-1}-192\zeta_{2}H_{-1,0}+136H_{-1,0}-32\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-912\zeta_{2}H_{0,1}+32\zeta_{2}H_{1,0}-192\zeta_{2}H_{1,1}-256H_{-1,-1,0}+144H_{-1,0,0}+128H_{-1,0,1}+96H_{0,-1,0}-472H_{0,1,0}+192H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-192H_{-1,-1,0,1}-160H_{-1,0,-1,0}-16H_{-1,0,0,0}-64H_{-1,0,0,1}+128H_{-1,0,1,0}+64H_{-1,0,1,1}-192H_{0,-1,-1,0}-64H_{0,-1,0,0}-64H_{0,-1,0,1}+96H_{0,0,-1,0}+128H_{0,0,0,0}+752H_{0,0,0,1}-688H_{0,0,1,0}-432H_{0,0,1,1}+416H_{0,1,0,0}+624H_{0,1,0,1}+432H_{0,1,1,0}+368H_{0,1,1,1}+96H_{1,0,-1,0}-48H_{1,0,0,0}+224H_{1,0,0,1}-544H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}-32H_{1,1,0,1}-224H_{1,1,1,0}-128H_{1,1,1,1}+z\left(\left(-\frac{6224}{9}\right)H_{0,1}+\left(-\frac{688}{3}\right)H_{0,0,1}+\left(-\frac{452}{3}\right)H_{1,0,1}+\left(-\frac{448}{3}\right)H_{1,0,0}+\frac{56}{3}H_{1,0}+\frac{68}{3}H_{0,1,1}+\frac{376}{9}H_{1,1}+\frac{296}{3}H_{1,1,1}+\frac{340}{3}H_{1,1,0}+\frac{364}{3}H_{0,0,0}+\frac{50168}{27}H_{1}+\frac{57247}{27}H_{0}-184H_{-1}\zeta_{2}+72H_{0}\zeta_{2}+380H_{1}\zeta_{2}+1176\zeta_{2}-88H_{-1}\zeta_{3}-136H_{0}\zeta_{3}+408H_{1}\zeta_{3}+368\zeta_{3}-152\zeta_{4}+144\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}+64H_{-1,0}-64\zeta_{2}H_{0,-1}-8\zeta_{2}H_{0,0}-889H_{0,0}+136\zeta_{2}H_{0,1}-16\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}-208H_{-1,-1,0}+24H_{-1,0,0}+80H_{-1,0,1}+264H_{0,-1,0}+528H_{0,1,0}+96H_{-1,-1,-1,0}+32H_{-1,-1,0,0}-96H_{-1,-1,0,1}-80H_{-1,0,-1,0}-8H_{-1,0,0,0}-32H_{-1,0,0,1}+64H_{-1,0,1,0}+32H_{-1,0,1,1}-64H_{0,-1,-1,0}-16H_{0,-1,0,0}+32H_{0,-1,0,1}-16H_{0,0,-1,0}+32H_{0,0,0,0}+168H_{0,0,0,1}-104H_{0,0,1,0}-8H_{0,0,1,1}-160H_{0,1,0,0}-168H_{0,1,0,1}-168H_{0,1,1,0}-136H_{0,1,1,1}-48H_{1,0,-1,0}+24H_{1,0,0,0}-112H_{1,0,0,1}+272H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}+16H_{1,1,0,1}+112H_{1,1,1,0}+64H_{1,1,1,1}-\frac{2737}{36}\right)+\frac{1577}{18}\bigg]$ (135)
$N_{f}C_{A}C_{F}\bigg[\left(-\frac{728}{27}\right)H_{0}+\left(-\frac{112}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{88}{9}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,1,1}+\frac{8}{9}\zeta_{2}+\frac{32}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{32}{3}H_{1,1,0}+\frac{112}{9}H_{1,1}+\frac{368}{9}H_{1,0}+\frac{160}{3}H_{1,0,0}+\frac{584}{9}H_{0,0}+\frac{256}{3}H_{0,0,0}+\left(-\frac{32\zeta_{2}}{3}\right)H_{0}+32H_{1}\zeta_{2}-80\zeta_{3}+z\left(\left(-\frac{356}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{208}{9}\right)H_{1,0}+\left(-\frac{518}{27}\right)H_{0}+\left(-\frac{184}{27}\right)H_{1}+\left(-\frac{16}{3}\right)H_{-1,0,0}+\left(-\frac{16}{3}\right)H_{0,-1,0}+\left(-\frac{16}{3}\right)H_{1,1,0}+\left(-\frac{32}{9}\right)H_{-1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\left(-\frac{8}{3}\right)H_{0,1,1}+\frac{8}{3}H_{1,1,1}+\frac{32}{9}H_{0,1}+\frac{16}{3}H_{0,0,0}+\frac{16}{3}H_{0,0,1}-16H_{1}\zeta_{2}-16\zeta_{2}+32\zeta_{3}+\frac{416}{9}\right)-\frac{526}{9}\bigg]\,$, (136)
$\Delta P_{qg}^{T,\text{s}(0)}=(4z-2)N_{f}\,$, (137)
$\Delta P_{qg}^{T,\text{s}(1)}=C_{A}N_{f}\bigg[z\left(-16H_{-1,0}+16H_{0,0}-32H_{0,1}+32H_{1,0}+16H_{1,1}+\left(-\frac{40}{3}\right)H_{1}+\left(-\frac{8}{3}\right)H_{0}+\frac{740}{9}\right)-8H_{-1,0}+24H_{0,0}+16H_{0,1}-16H_{1,0}-8H_{1,1}+\left(-\frac{32}{3}\right)H_{0}+\left(-\frac{4}{3}\right)H_{1}-8\zeta_{2}-\frac{436}{9}\bigg]+C_{F}N_{f}\bigg[z\left(-8H_{0,0}+16H_{0,1}-48H_{1,0}-16H_{1,1}+8H_{1}-16\zeta_{2}-50\right)+4H_{0,0}-8H_{0,1}+24H_{1,0}+8H_{1,1}+22H_{0}+4H_{1}+8\zeta_{2}+32\bigg]+N_{f}^{2}\bigg[\left(\left(-\frac{16}{3}\right)H_{0}+\frac{16}{3}H_{1}-\frac{32}{9}\right)z+\left(-\frac{8}{3}\right)H_{1}+\frac{8}{3}H_{0}-\frac{8}{9}\bigg]\,$, (138)
$\Delta P_{qg}^{T,\text{s}(2)}=N_{f}^{3}\bigg[\left(-\frac{8}{3}\right)\zeta_{2}+\left(-\frac{8}{3}\right)H_{0,0}+\left(-\frac{8}{3}\right)H_{1,1}+\left(-\frac{8}{9}\right)H_{1}+\frac{8}{9}H_{0}+\frac{8}{3}H_{0,1}+\frac{8}{3}H_{1,0}+z\left(\left(-\frac{16}{3}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1}+\frac{32}{9}H_{0}+\frac{16}{3}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{16}{3}H_{1,1}-\frac{16}{3}\right)+\frac{40}{9}\bigg]+C_{F}N_{f}^{2}\bigg[\left(16\zeta_{2}+16H_{-1,0}-16H_{0,0}\right)z^{2}+\left(\left(-\frac{128}{3}\right)H_{0,0,1}+\left(-\frac{112}{3}\right)H_{1,1,0}+\left(-\frac{280}{9}\right)H_{0,1}+\left(-\frac{260}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,1}+\left(-\frac{80}{3}\right)H_{1,1,1}+\left(-\frac{680}{27}\right)H_{1}+\frac{280}{9}\zeta_{2}+\frac{328}{9}H_{1,1}+\frac{128}{3}H_{0,1,1}+\frac{160}{3}H_{0,1,0}+\frac{232}{3}H_{1,0}+\frac{320}{3}H_{0,0,0}+\frac{5090}{27}H_{0}+\frac{608}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}-16H_{1}\zeta_{2}-64H_{-1,0}+64H_{0,-1,0}+64H_{1,0,0}-64H_{0,0,0,0}-\frac{1157}{3}\right)z$ (139)
$\left(-\frac{304}{3}\right)\zeta_{3}+\left(-\frac{848}{9}\right)\zeta_{2}+\left(-\frac{80}{3}\right)H_{0,1,0}+\left(-\frac{64}{3}\right)H_{0,1,1}+\left(-\frac{56}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\frac{184}{27}H_{1}+\frac{40}{3}H_{1,0,1}+\frac{40}{3}H_{1,1,1}+\frac{56}{3}H_{1,1,0}+\frac{64}{3}H_{0,0,1}+\frac{272}{9}H_{0,1}+\frac{898}{9}H_{0,0}+\frac{320}{3}H_{0,0,0}+\frac{3830}{27}H_{0}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}+8H_{1}\zeta_{2}-80H_{-1,0}-32H_{0,-1,0}-32H_{1,0,0}+32H_{0,0,0,0}+\frac{6587}{18}\bigg]+N_{f}C_{A}^{2}\bigg[\left(72H_{-1}\zeta_{2}-96H_{0}\zeta_{2}-24H_{1}\zeta_{2}+24\zeta_{2}-120\zeta_{3}+24H_{-1,0}-24H_{0,0}+48H_{-1,-1,0}-48H_{-1,0,0}-48H_{-1,0,1}-48H_{0,-1,0}+48H_{0,0,0}+48H_{0,0,1}\right)z^{2}+\left(\left(-\frac{8836}{9}\right)H_{0,1}+\left(-\frac{976}{3}\right)H_{0,0,0}+\left(-\frac{848}{3}\right)H_{-1,-1,0}+\left(-\frac{2348}{9}\right)H_{0,0}+\left(-\frac{592}{3}\right)H_{1,0,1}+\left(-\frac{592}{3}\right)H_{1,1,0}+\left(-\frac{440}{3}\right)H_{1,1,1}+\left(-\frac{3796}{27}\right)H_{0}+\left(-\frac{404}{3}\right)H_{1,0,0}+\left(-\frac{320}{3}\right)H_{-1,0,1}+\left(-\frac{308}{3}\right)\zeta_{2}+\left(-\frac{304}{3}\right)H_{-1,0}+\frac{284}{3}H_{0,0,1}+\frac{536}{3}H_{-1,0,0}+\frac{5464}{27}H_{1}+\frac{1828}{9}H_{1,1}+\frac{2068}{9}H_{1,0}+\frac{824}{3}H_{0,-1,0}+\left(-\frac{104\zeta_{2}}{3}\right)H_{-1}+\frac{472\zeta_{2}}{3}H_{1}+60H_{0}\zeta_{2}+272H_{-1}\zeta_{3}+32H_{0}\zeta_{3}+496H_{1}\zeta_{3}-184\zeta_{3}+170\zeta_{4}-416\zeta_{2}H_{-1,-1}+224\zeta_{2}H_{-1,0}+144\zeta_{2}H_{0,-1}-56\zeta_{2}H_{0,0}+336\zeta_{2}H_{0,1}+96\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}-24H_{0,1,0}+192H_{0,1,1}-192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}+320H_{-1,-1,0,1}+224H_{-1,0,-1,0}+16H_{-1,0,0,0}-128H_{-1,0,1,0}-64H_{-1,0,1,1}+32H_{0,-1,-1,0}-208H_{0,-1,0,0}-128H_{0,-1,0,1}-112H_{0,0,-1,0}+128H_{0,0,0,0}-200H_{0,0,0,1}+128H_{0,0,1,0}+352H_{0,0,1,1}+184H_{0,1,0,0}-160H_{0,1,1,1}-32H_{1,0,-1,0}-240H_{1,0,0,0}-256H_{1,0,0,1}-64H_{1,0,1,1}-320H_{1,1,0,0}-64H_{1,1,0,1}-64H_{1,1,1,0}+64H_{1,1,1,1}+\frac{5408}{3}\right)z+\left(-\frac{19762}{27}\right)H_{0}+\left(-\frac{9620}{27}\right)H_{1}+\left(-\frac{928}{3}\right)H_{0,0,0}+\left(-\frac{640}{3}\right)H_{-1,-1,0}+\left(-\frac{1502}{9}\right)H_{0,0}+\left(-\frac{352}{3}\right)H_{-1,0,1}+\left(-\frac{566}{9}\right)H_{1,1}+\left(-\frac{434}{9}\right)H_{1,0}+\frac{220}{3}H_{1,1,1}+\frac{232}{3}H_{-1,0}+\frac{304}{3}\zeta_{3}+\frac{680}{3}H_{1,0,1}+\frac{680}{3}H_{1,1,0}+\frac{776}{3}H_{0,-1,0}+\frac{2390}{9}H_{0,1}+\frac{856}{3}H_{-1,0,0}+\frac{1240}{3}H_{1,0,0}+\frac{1342}{3}\zeta_{2}+\left(-\frac{224\zeta_{2}}{3}\right)H_{1}+\frac{32\zeta_{2}}{3}H_{-1}+\frac{856\zeta_{2}}{3}H_{0}+136H_{-1}\zeta_{3}+480H_{0}\zeta_{3}-248H_{1}\zeta_{3}+63\zeta_{4}-208\zeta_{2}H_{-1,-1}+112\zeta_{2}H_{-1,0}-88\zeta_{2}H_{0,-1}+316\zeta_{2}H_{0,0}-216\zeta_{2}H_{0,1}-48\zeta_{2}H_{1,0}+16\zeta_{2}H_{1,1}-40H_{0,0,1}+4H_{0,1,0}-56H_{0,1,1}-96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}+160H_{-1,-1,0,1}+112H_{-1,0,-1,0}+8H_{-1,0,0,0}-64H_{-1,0,1,0}-32H_{-1,0,1,1}-176H_{0,-1,-1,0}+216H_{0,-1,0,0}+248H_{0,0,-1,0}-632H_{0,0,0,0}-388H_{0,0,0,1}-48H_{0,0,1,0}-96H_{0,0,1,1}+204H_{0,1,0,0}+144H_{0,1,0,1}+144H_{0,1,1,0}+128H_{0,1,1,1}+16H_{1,0,-1,0}+120H_{1,0,0,0}+128H_{1,0,0,1}+32H_{1,0,1,1}+160H_{1,1,0,0}+32H_{1,1,0,1}+32H_{1,1,1,0}-32H_{1,1,1,1}-\frac{4670}{3}\bigg]+N_{f}C_{F}^{2}\bigg[\left(48\zeta_{2}+48H_{-1,0}-48H_{0,0}\right)z^{2}+\left(-80\zeta_{2}H_{0}+80\zeta_{3}H_{0}+289H_{0}+600H_{1}-64H_{-1}\zeta_{2}-176H_{1}\zeta_{2}+612\zeta_{2}+992H_{1}\zeta_{3}-368\zeta_{3}+280\zeta_{4}-192\zeta_{2}H_{-1,0}+304H_{-1,0}+128\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-350H_{0,0}+208\zeta_{2}H_{0,1}-340H_{0,1}-224\zeta_{2}H_{1,0}+592H_{1,0}-32\zeta_{2}H_{1,1}+32H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+56H_{0,0,0}+16H_{0,0,1}-16H_{0,1,0}+32H_{0,1,1}-64H_{1,0,0}-48H_{1,0,1}-144H_{1,1,0}-48H_{1,1,1}-384H_{-1,-1,0,0}-192H_{-1,0,-1,0}+352H_{-1,0,0,0}+256H_{0,-1,-1,0}+64H_{0,-1,0,0}-192H_{0,0,-1,0}+32H_{0,0,0,0}+16H_{0,0,0,1}+192H_{0,0,1,0}+128H_{0,0,1,1}-352H_{0,1,0,0}-208H_{0,1,0,1}-240H_{0,1,1,0}-112H_{0,1,1,1}-192H_{1,0,-1,0}+672H_{1,0,0,0}+160H_{1,0,0,1}+608H_{1,0,1,0}+128H_{1,0,1,1}+32H_{1,1,0,0}+160H_{1,1,0,1}+96H_{1,1,1,0}+64H_{1,1,1,1}+\frac{3747}{2}\right)z$ (140)
$1505H_{0}-582H_{1}-64H_{-1}\zeta_{2}+164H_{0}\zeta_{2}+140H_{1}\zeta_{2}+394\zeta_{2}-152H_{0}\zeta_{3}-496H_{1}\zeta_{3}-548\zeta_{3}-476\zeta_{4}-96\zeta_{2}H_{-1,0}+288H_{-1,0}+32\zeta_{2}H_{0,-1}+72\zeta_{2}H_{0,0}-947H_{0,0}-72\zeta_{2}H_{0,1}-242H_{0,1}+112\zeta_{2}H_{1,0}-390H_{1,0}+16\zeta_{2}H_{1,1}-34H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+144H_{0,-1,0}-476H_{0,0,0}-68H_{0,0,1}-304H_{0,1,0}-80H_{0,1,1}-32H_{1,0,0}-60H_{1,0,1}+12H_{1,1,0}-12H_{1,1,1}-192H_{-1,-1,0,0}-96H_{-1,0,-1,0}+176H_{-1,0,0,0}+64H_{0,-1,-1,0}+64H_{0,-1,0,0}+32H_{0,0,-1,0}-128H_{0,0,0,0}-8H_{0,0,0,1}-96H_{0,0,1,0}-64H_{0,0,1,1}+144H_{0,1,0,0}+104H_{0,1,0,1}+120H_{0,1,1,0}+56H_{0,1,1,1}+96H_{1,0,-1,0}-336H_{1,0,0,0}-80H_{1,0,0,1}-304H_{1,0,1,0}-64H_{1,0,1,1}-16H_{1,1,0,0}-80H_{1,1,0,1}-48H_{1,1,1,0}-32H_{1,1,1,1}-\frac{7379}{4}\bigg]+C_{A}N_{f}^{2}\bigg[\left(-16\zeta_{2}-16H_{-1,0}+16H_{0,0}\right)z^{2}+\left(\left(-\frac{1000}{9}\right)H_{0,0}+\left(-\frac{304}{3}\right)H_{1,0,0}+\left(-\frac{2576}{27}\right)H_{0}+\left(-\frac{784}{9}\right)H_{1,0}+\left(-\frac{640}{9}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,-1,0}+\left(-\frac{128}{3}\right)H_{0,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{32}{3}\right)H_{1,0,1}+\left(-\frac{32}{3}\right)H_{1,1,0}+\frac{16}{3}\zeta_{2}+\frac{64}{3}H_{-1,0,0}+\frac{80}{3}H_{1,1,1}+\frac{128}{3}H_{-1,-1,0}+\frac{160}{3}H_{-1,0}+\frac{592}{9}H_{0,1}+\frac{208}{3}H_{0,0,1}+\frac{2456}{27}H_{1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{1}+\frac{128\zeta_{2}}{3}H_{-1}-16H_{0}\zeta_{2}-80\zeta_{3}+16H_{0,1,0}-64H_{0,1,1}+70\right)z+\left(-\frac{320}{3}\right)H_{0,0,0}+\left(-\frac{748}{9}\right)H_{0,0}+\left(-\frac{902}{27}\right)H_{0}+\left(-\frac{284}{9}\right)H_{0,1}+\left(-\frac{784}{27}\right)H_{1}+\left(-\frac{40}{3}\right)H_{1,1,1}+\left(-\frac{32}{3}\right)H_{-1,0,1}+\frac{16}{3}H_{1,0,1}+\frac{16}{3}H_{1,1,0}+\frac{32}{3}H_{-1,0,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,-1,0}+\frac{296}{9}H_{1,1}+\frac{128}{3}H_{-1,0}+\frac{152}{3}H_{1,0,0}+\frac{176}{3}\zeta_{3}+\frac{196}{3}\zeta_{2}+\frac{728}{9}H_{1,0}+\frac{8\zeta_{2}}{3}H_{0}+\frac{32\zeta_{2}}{3}H_{1}+\frac{64\zeta_{2}}{3}H_{-1}-8H_{0,0,1}+40H_{0,1,0}+48H_{0,1,1}-\frac{988}{9}\bigg]+C_{A}C_{F}N_{f}\bigg[\left(-72\zeta_{2}-72H_{-1,0}+72H_{0,0}\right)z^{2}+\left(\left(-\frac{3484}{3}\right)H_{1,0}+\left(-\frac{25828}{27}\right)H_{1}+\left(-\frac{25589}{27}\right)H_{0}+\left(-\frac{7276}{9}\right)\zeta_{2}+\left(-\frac{2284}{9}\right)H_{1,1}+\left(-\frac{536}{3}\right)H_{0,1,1}+\frac{128}{3}H_{0,0,1}+\frac{272}{3}H_{0,1,0}+\frac{280}{3}H_{0,0,0}+\frac{584}{3}H_{1,1,1}+\frac{784}{3}\zeta_{3}+\frac{848}{3}H_{1,0,1}+\frac{1168}{3}H_{1,1,0}+\frac{5774}{9}H_{0,0}+\frac{10156}{9}H_{0,1}+\frac{80\zeta_{2}}{3}H_{0}+176H_{-1}\zeta_{2}-16H_{1}\zeta_{2}-80H_{-1}\zeta_{3}-112H_{0}\zeta_{3}-1584H_{1}\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{-1,-1}+48H_{-1,0}-128\zeta_{2}H_{0,-1}+64\zeta_{2}H_{0,0}-400\zeta_{2}H_{0,1}+160\zeta_{2}H_{1,0}+64\zeta_{2}H_{1,1}+224H_{-1,-1,0}-288H_{-1,0,0}-64H_{-1,0,1}-176H_{0,-1,0}+344H_{1,0,0}+192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}-64H_{-1,-1,0,1}-160H_{-1,0,-1,0}-48H_{-1,0,0,0}+192H_{-1,0,1,0}+64H_{-1,0,1,1}-256H_{0,-1,-1,0}+288H_{0,0,-1,0}-64H_{0,0,0,0}+128H_{0,0,0,1}-320H_{0,0,1,0}-352H_{0,0,1,1}+272H_{0,1,0,0}+336H_{0,1,0,1}+368H_{0,1,1,0}+272H_{0,1,1,1}+160H_{1,0,-1,0}-112H_{1,0,0,0}+224H_{1,0,0,1}-480H_{1,0,1,0}-64H_{1,0,1,1}+416H_{1,1,0,0}-96H_{1,1,0,1}-32H_{1,1,1,0}-128H_{1,1,1,1}-\frac{9751}{6}\right)z+\left(-\frac{3016}{9}\right)\zeta_{2}+\left(-\frac{788}{3}\right)H_{1,1,0}+\left(-\frac{556}{3}\right)H_{1,0,1}+\left(-\frac{220}{3}\right)H_{0,0,1}+\left(-\frac{184}{3}\right)H_{1,1,1}+\frac{640}{9}H_{0,1}+\frac{244}{3}H_{0,0,0}+\frac{1064}{9}H_{1,1}+\frac{400}{3}H_{0,1,1}+\frac{716}{3}H_{0,1,0}+\frac{4607}{9}H_{0,0}+\frac{1904}{3}H_{1,0}+\frac{2872}{3}\zeta_{3}+\frac{28226}{27}H_{1}+\frac{31993}{27}H_{0}+\left(-\frac{364\zeta_{2}}{3}\right)H_{0}+80H_{-1}\zeta_{2}+12H_{1}\zeta_{2}-40H_{-1}\zeta_{3}+8H_{0}\zeta_{3}+792H_{1}\zeta_{3}+658\zeta_{4}+80\zeta_{2}H_{-1,-1}+8H_{-1,0}-16\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+296\zeta_{2}H_{0,1}-80\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}+128H_{-1,-1,0}-240H_{-1,0,0}-16H_{-1,0,1}+16H_{0,-1,0}-592H_{1,0,0}+96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-32H_{-1,-1,0,1}-80H_{-1,0,-1,0}-24H_{-1,0,0,0}+96H_{-1,0,1,0}+32H_{-1,0,1,1}-32H_{0,-1,-1,0}-48H_{0,-1,0,0}+16H_{0,0,-1,0}-32H_{0,0,0,0}+48H_{0,0,0,1}-80H_{0,0,1,0}+96H_{0,0,1,1}-520H_{0,1,0,0}-312H_{0,1,0,1}-328H_{0,1,1,0}-184H_{0,1,1,1}-80H_{1,0,-1,0}+56H_{1,0,0,0}-112H_{1,0,0,1}+240H_{1,0,1,0}+32H_{1,0,1,1}-208H_{1,1,0,0}+48H_{1,1,0,1}+16H_{1,1,1,0}+64H_{1,1,1,1}+\frac{52313}{36}\bigg]\,$, (141)
$\Delta P_{gg}^{T,\text{s}(0)}=4C_{A}\left[\frac{1}{1-z}\right]_{+}+\delta(1-z)\left(\frac{11C_{A}}{3}-\frac{2N_{f}}{3}\right)+\left(-\frac{8z^{2}}{z-1}+\frac{12z}{z-1}-\frac{4}{z-1}\right)C_{A}\,$, (142)
$\Delta P_{gg}^{T,\text{s}(1)}=\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\frac{268}{9}-8\zeta_{2}\right)C_{A}^{2}-\frac{40C_{A}N_{f}}{9}\bigg]+C_{A}^{2}\bigg[\frac{z^{2}}{z-1}\left(-64H_{0,0}+\frac{1}{z+1}-\frac{74}{9}\right)+\frac{z^{2}}{z-1}\left(32H_{0,1}+32H_{1,0}-20H_{0}\right)+\frac{32z^{2}}{z+1}H_{-1,0}+\frac{1}{z-1}\left(32H_{0,1}+32H_{1,0}+\frac{92}{3}H_{0}-8\zeta_{2}+\frac{268}{9}\right)+\frac{z}{z-1}\left(16H_{0,0}+\frac{1}{z+1}-16\zeta_{2}-\frac{74}{9}\right)+\frac{z}{z-1}\left(-48H_{0,1}-48H_{1,0}-40H_{0}\right)+\frac{32}{z+1}H_{-1,0}+\frac{1}{(z-1)(z+1)}\left(96H_{0,0}-\frac{194}{9}\right)+\frac{48z}{z+1}H_{-1,0}+\frac{z^{3}\left(32\zeta_{2}-\frac{194}{9}\right)}{(z-1)(z+1)}\bigg]+C_{F}N_{f}\bigg[z\left(8H_{0,0}+4H_{0}-28\right)+8H_{0,0}+28H_{0}+28\bigg]+\delta(1-z)\bigg[-\frac{8C_{A}N_{f}}{3}+\left(12\zeta_{3}+\frac{32}{3}\right)C_{A}^{2}-2C_{F}N_{f}\bigg]+C_{A}N_{f}\bigg[\frac{z^{2}}{z-1}\left(8H_{0}+\frac{152}{9}\right)+\frac{1}{z-1}\left(\frac{40}{3}H_{0}+\frac{152}{9}\right)+\frac{z}{z-1}\left(-16H_{0}-\frac{88}{3}\right)-\frac{40}{9(z-1)}\bigg]\,$, (143)
$\Delta P_{gg}^{T,\text{s}(2)}=C_{A}^{3}\bigg[\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)z^{2}+\left(336H_{-1,-1,0}+368H_{-1,0,1}\right)z+\frac{1}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{1072}{9}\right)\zeta_{2}+\left(-\frac{8}{9}\right)H_{1}+\frac{88}{3}\zeta_{3}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}+88\zeta_{4}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-208H_{0,0,1}-144H_{0,1,0}-256H_{0,0,1,1}-608H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}+\frac{490}{3}\right)+\frac{1}{z}\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{48478}{27}\right)H_{0}+\left(-\frac{2968}{3}\right)\zeta_{3}+\left(-\frac{320}{3}\right)H_{0,1}+\left(-\frac{284}{3}\right)H_{0,0}+\left(-\frac{592}{9}\right)\zeta_{2}+\frac{320}{3}H_{0,-1,0}+\frac{280\zeta_{2}}{3}H_{0}+472\zeta_{4}-96\zeta_{2}H_{0,0}+160\zeta_{2}H_{0,1}-112H_{0,0,0}+64H_{0,-1,-1,0}-224H_{0,-1,0,0}-128H_{0,-1,0,1}-64H_{0,0,-1,0}+128H_{0,0,0,0}-160H_{0,0,0,1}-64H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)+\frac{z}{z-1}\left(\left(-\frac{1232}{3}\right)H_{1,0}+\frac{16}{9}H_{1}-336H_{1}\zeta_{2}-288H_{1}\zeta_{3}+416\zeta_{2}H_{0,-1}+96\zeta_{2}H_{1,0}+584H_{0,0,1}+320H_{0,1,0}+568H_{1,0,0}+384H_{0,0,1,1}+672H_{0,1,0,0}+384H_{0,1,0,1}+384H_{0,1,1,0}-192H_{1,0,-1,0}+960H_{1,0,0,0}+384H_{1,0,0,1}+384H_{1,0,1,0}+384H_{1,1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3232}{3}\right)H_{0,-1,0}+\left(-\frac{9136}{9}\right)\zeta_{2}+\left(-\frac{592}{9}\right)H_{0,1}+\frac{16964}{9}H_{0,0}+\frac{6736}{3}H_{0,0,0}+\frac{8990}{3}H_{0}+\left(-\frac{2024\zeta_{2}}{3}\right)H_{0}-800H_{0}\zeta_{3}-920\zeta_{3}-576\zeta_{4}-608\zeta_{2}H_{0,0}-32\zeta_{2}H_{0,1}+320H_{0,-1,-1,0}-544H_{0,-1,0,0}-704H_{0,0,-1,0}+1984H_{0,0,0,0}+800H_{0,0,0,1}+512H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)$ (144)
$\frac{z^{2}}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{8}{9}\right)H_{1}+\frac{176}{3}H_{0,1,0}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-112H_{0,0,1}-256H_{0,0,1,1}-288H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,1}+\frac{88\zeta_{2}}{3}H_{1}\right)+\frac{z^{3}}{z-1}\left(\frac{1060}{9}H_{0,0}+\frac{3104}{9}H_{0,1}+\frac{2792}{3}\zeta_{3}+\frac{13402}{9}H_{0}+\frac{160\zeta_{2}}{3}H_{0}+304\zeta_{2}-680\zeta_{4}+96\zeta_{2}H_{0,0}-256\zeta_{2}H_{0,1}-48H_{0,-1,0}+464H_{0,0,0}+128H_{0,-1,0,0}+256H_{0,-1,0,1}-64H_{0,0,-1,0}-256H_{0,0,0,0}-96H_{0,0,0,1}-128H_{0,0,1,0}+\frac{1}{z+1}-\frac{47572}{9}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,0}+\frac{176}{3}H_{0,-1,0}+\frac{440}{3}\zeta_{3}+\frac{352\zeta_{2}}{3}H_{0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{89182}{27}\right)H_{0}+\left(-\frac{14612}{9}\right)H_{0,0}+\left(-\frac{4448}{3}\right)H_{0,0,0}+\frac{2144}{3}\zeta_{3}+\frac{2528}{3}H_{0,-1,0}+\frac{9136}{9}\zeta_{2}+528H_{0}\zeta_{2}+1088H_{0}\zeta_{3}+808\zeta_{4}+864\zeta_{2}H_{0,0}+304H_{0,1}-512H_{0,-1,-1,0}+768H_{0,-1,0,0}+960H_{0,0,-1,0}-2496H_{0,0,0,0}-1184H_{0,0,0,1}-768H_{0,0,1,0}-\frac{47572}{9}\right)+\frac{1}{z(z+1)}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{z}{z+1}\left(\left(-\frac{4736}{3}\right)H_{-1,0}-400H_{-1}\zeta_{2}-576H_{-1}\zeta_{3}+768\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}-1312H_{-1,0,0}+384H_{-1,-1,0,0}-768H_{-1,-1,0,1}+384H_{-1,0,-1,0}-960H_{-1,0,0,0}-192H_{-1,0,1,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{z^{3}}{z+1}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{88\zeta_{2}}{3(z-1)z}H_{1}+336H_{-1,-1,0}+368H_{-1,0,1}\bigg]+N_{f}C_{A}^{2}\bigg[\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\frac{16\zeta_{2}}{3}H_{1}-16H_{-1}\zeta_{2}\right)z^{2}+32H_{-1,0,1}+96H_{0,-1,-1,0}+\left(\left(-\frac{2174}{9}\right)H_{1}+56H_{-1}\zeta_{2}+88H_{1}\zeta_{2}-48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+32H_{-1,0,1}-96H_{0,-1,-1,0}+48H_{0,-1,0,0}+80H_{0,0,-1,0}+24H_{0,0,0,1}+24H_{0,1,0,0}\right)z+\frac{2174}{9}H_{1}-48H_{0,-1,0,0}-16H_{0,0,-1,0}+\frac{1}{z-1}\left(\left(-\frac{1844}{9}\right)H_{0,1}+\left(-\frac{356}{3}\right)H_{1,0,0}+\left(-\frac{928}{9}\right)H_{1,0}+\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}+48\zeta_{2}H_{0,1}-124H_{0,0,1}-96H_{0,1,0}-\frac{836}{27}\right)+16H_{0,0,0,0}+24H_{0,0,0,1}+24H_{0,1,0,0}+\frac{1}{z}\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\left(-\frac{16\zeta_{2}}{3}\right)H_{1}-16H_{-1}\zeta_{2}\right)+\frac{z}{z-1}\left(\left(-\frac{1396}{9}\right)H_{0,0}+\left(-\frac{368}{3}\right)\zeta_{3}+\left(-\frac{1138}{27}\right)H_{0}+\frac{376}{3}H_{0,-1,0}+\frac{508}{3}\zeta_{2}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-138\zeta_{4}+24\zeta_{2}H_{0,0}-32H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+56H_{-1}\zeta_{2}-88H_{1}\zeta_{2}+48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}$ (145)
$\frac{z}{z-1}\left(\frac{512}{3}H_{0,1}+\frac{512}{3}H_{1,0}+200H_{0,0,1}+128H_{0,1,0}+200H_{1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{356}{3}\right)H_{1,0,0}+\left(-\frac{928}{9}\right)H_{1,0}+\left(-\frac{308}{3}\right)H_{0,0,1}+\left(-\frac{224}{3}\right)H_{0,1,0}+\left(-\frac{4}{3}\right)H_{0,1}-48\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{832}{3}\right)\zeta_{3}+\left(-\frac{2140}{9}\right)\zeta_{2}+\frac{1880}{3}H_{0,0}+\frac{6308}{9}H_{0}-116H_{0}\zeta_{2}-96H_{0}\zeta_{3}+78\zeta_{4}-40\zeta_{2}H_{0,0}-104H_{0,-1,0}+392H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+\left(-\frac{32z^{3}}{3(z-1)}\right)H_{0,0,1}+\frac{z^{3}}{z-1}\left(\left(-\frac{1844}{9}\right)\zeta_{2}+\frac{1246}{9}H_{0}+\frac{496}{3}\zeta_{3}+\frac{1780}{9}H_{0,0}+\frac{20\zeta_{2}}{3}H_{0}+138\zeta_{4}-24\zeta_{2}H_{0,0}-136H_{0,-1,0}-32H_{0,0,0}+\frac{1}{z+1}-\frac{22904}{27}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{32}{3}H_{0,-1,0}+\frac{80}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}+16\zeta_{2}-16H_{0,0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{16324}{27}\right)H_{0}+\left(-\frac{5432}{9}\right)H_{0,0}+\left(-\frac{1528}{3}\right)H_{0,0,0}+\frac{376}{3}H_{0,-1,0}+\frac{1996}{9}\zeta_{2}+\frac{880}{3}\zeta_{3}+100H_{0}\zeta_{2}+96H_{0}\zeta_{3}-78\zeta_{4}+40\zeta_{2}H_{0,0}-\frac{22904}{27}\right)+\frac{1}{z(z+1)}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)+\frac{z}{z+1}\left(\left(-\frac{1232}{3}\right)H_{-1,0}-272H_{-1,0,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{z^{3}}{z+1}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)\bigg]+C_{A}N_{f}^{2}\bigg[\left(-\frac{82}{9}\right)H_{1}+\left(-\frac{32}{9}\right)H_{0,1}+z\left(\left(-\frac{32}{9}\right)H_{0,1}+\frac{82}{9}H_{1}\right)+\frac{1}{z-1}\left(\left(-\frac{176}{9}\right)H_{0,0}+\left(-\frac{496}{27}\right)H_{0}+\left(-\frac{32}{9}\right)\zeta_{2}+\frac{230}{27}\right)+\frac{z}{z-1}\left(\frac{64}{3}H_{0,0}+30H_{0}-\frac{148}{9}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{158}{9}\right)H_{0}+\left(-\frac{80}{9}\right)H_{0,0}+\frac{32}{9}\zeta_{2}+\frac{230}{27}\right)-\frac{16}{27(z-1)}\bigg]+N_{f}C_{F}C_{A}\bigg[\left(\left(-\frac{64}{3}\right)H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{64}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,0,0}+\frac{64}{3}H_{0,0,1}+\left(-\frac{32\zeta_{2}}{3}\right)H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}+48H_{0,0}\right)z^{2}+\left(\left(-\frac{1756}{9}\right)H_{1}+\left(-\frac{488}{3}\right)H_{0,1,0}+\left(-\frac{296}{3}\right)H_{0,0,0}+\left(-\frac{152}{3}\right)H_{0,1,1}+\frac{260}{3}H_{1,1}+\frac{2932}{9}H_{0,1}+\frac{1004}{3}H_{1,0}+\frac{1352}{3}H_{-1,0}+\frac{2168}{3}H_{0,0}-176H_{-1}\zeta_{2}-640H_{1}\zeta_{2}+128\zeta_{2}H_{0,-1}-16\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+272H_{0,-1,0}+8H_{0,0,1}+712H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+256H_{0,-1,-1,0}-192H_{0,-1,0,0}-256H_{0,0,-1,0}+128H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)z+\left(-\frac{3632}{3}\right)H_{0,0}+\left(-\frac{3056}{3}\right)H_{0,0,0}+\left(-\frac{1004}{3}\right)H_{1,0}+\left(-\frac{260}{3}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,1,0}+\frac{40}{9}H_{0,1}+\frac{88}{3}H_{0,1,1}+\frac{1756}{9}H_{1}+\frac{1352}{3}H_{-1,0}+\frac{1}{z-1}\left(\left(-\frac{1936}{3}\right)\zeta_{3}+\left(-\frac{2672}{9}\right)\zeta_{2}+\frac{51500}{27}H_{0}+\left(-\frac{416\zeta_{2}}{3}\right)H_{0}-512H_{0}\zeta_{3}-468\zeta_{4}+\frac{88478}{27}\right)+\frac{1}{z}\left(\left(-\frac{64}{3}\right)H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\frac{64}{3}H_{-1,-1,0}+\frac{32\zeta_{2}}{3}H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{3148}{9}\right)H_{0}+\frac{1420}{3}\zeta_{2}+168H_{0}\zeta_{2}+352H_{0}\zeta_{3}+1240\zeta_{3}+160\zeta_{4}-\frac{177946}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{41624}{27}\right)H_{0}+\left(-\frac{1624}{3}\right)\zeta_{3}+\left(-\frac{1156}{9}\right)\zeta_{2}+\frac{40\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+308\zeta_{4}+\frac{89468}{27}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{160}{3}\right)\zeta_{3}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}\right)\bigg]$ (146)
$176H_{-1}\zeta_{2}+640H_{1}\zeta_{2}-128\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+368H_{0,-1,0}-32H_{0,0,1}-712H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-256H_{0,-1,-1,0}+192H_{0,-1,0,0}+128H_{0,0,-1,0}-448H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}+\delta(1-z)\bigg[\left(\frac{8}{3}\zeta_{2}+\frac{55}{3}\zeta_{4}+\frac{536}{3}\zeta_{3}-16\zeta_{2}\zeta_{3}-80\zeta_{5}+\frac{79}{2}\right)C_{A}^{3}+\left(\left(-\frac{80}{3}\right)\zeta_{3}+\left(-\frac{10}{3}\right)\zeta_{4}+\left(-\frac{8}{3}\right)\zeta_{2}-\frac{233}{18}\right)N_{f}C_{A}^{2}+\left(\frac{29N_{f}^{2}}{18}-\frac{241C_{F}N_{f}}{18}\right)C_{A}+\frac{11}{9}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{3}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)N_{f}C_{A}^{2}+\left(C_{F}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)-\frac{16}{27}N_{f}^{2}\right)C_{A}\bigg]+C_{F}N_{f}^{2}\bigg[\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{40}{3}\right)H_{1,0}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{328}{27}\right)H_{0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{136}{9}\zeta_{2}+\frac{224}{9}H_{1}+\frac{344}{9}H_{0,1}+\frac{16\zeta_{2}}{3}H_{0}-32H_{0,0}+16H_{0,0,1}+z\left(\left(-\frac{488}{9}\right)\zeta_{2}+\left(-\frac{224}{9}\right)H_{1}+\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{8}{9}H_{0,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,1}+\frac{776}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}+16H_{0,0,1}-\frac{992}{27}\right)+\frac{992}{27}\bigg]+C_{F}^{2}N_{f}\bigg[\left(32\zeta_{2}+32H_{-1,0}-32H_{0,0}\right)z^{2}+\left(-48\zeta_{2}H_{0}-160\zeta_{3}H_{0}+1032H_{0}+580H_{1}+128H_{-1}\zeta_{2}+528H_{1}\zeta_{2}+580\zeta_{2}-48\zeta_{3}-380\zeta_{4}+80H_{-1,0}-64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}-664H_{0,0}-224\zeta_{2}H_{0,1}-404H_{0,1}-68H_{1,0}-52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-256H_{0,-1,0}+80H_{0,0,1}+112H_{0,1,0}+32H_{0,1,1}-568H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-128H_{0,-1,-1,0}+64H_{0,-1,0,0}+128H_{0,0,-1,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}-664\right)z+\frac{32}{z}H_{-1,0}+252H_{0}-580H_{1}+128H_{-1}\zeta_{2}+64H_{0}\zeta_{2}-528H_{1}\zeta_{2}+152\zeta_{2}-288H_{0}\zeta_{3}-272\zeta_{3}-460\zeta_{4}+80H_{-1,0}+64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}+16H_{0,0}-224\zeta_{2}H_{0,1}-248H_{0,1}+68H_{1,0}+52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-64H_{0,0,0}-160H_{0,0,1}-192H_{0,1,0}-48H_{0,1,1}+568H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+128H_{0,-1,-1,0}-64H_{0,-1,0,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+664\bigg]\,$, (147)
4.5 N3LO coefficient functions and small-x expansion
The analytic expressions for the coefficient functions will be provided in the ancillary files along with the arXiv submission. In this section we study the asymptotic behaviors and present their numerical fits [^2] To determine the 3-loop pure-singlet $\Delta\mathcal{I}^{(3)}_{qq^{\prime}}$ in $\overline{\text{MS}}$}, we also need $z_{\text{ps}}^{(3)}$, which is currently not known..
The coefficient functions develop end-point divergences both in the threshold and high energy limit. In the $z\to 1$ limit, we have exactly the same results as the unpolarized case Luo:2019szz ; Luo:2020epw
$\lim_{z\to 1}\Delta\mathcal{I}^{(2)}_{ij}(z)=\lim_{z\to 1}\Delta\mathcal{C}^{(2)}_{ji}(z)=\frac{2\gamma^{R}_{1,i}}{(1-z)_{+}}\delta_{ij}\,,\quad\lim_{z\to 1}\Delta\mathcal{I}^{(3)}_{ij}(z)=\lim_{z\to 1}\Delta\mathcal{C}^{(3)}_{ji}(z)=\frac{2\gamma^{R}_{2,i}}{(1-z)_{+}}\delta_{ij}\,$, (148)
where $\gamma^{R}_{1(2)}$ are the two(three)-loop rapidity anomalous dimensions Li:2016ctv ; Vladimirov:2016dll . The relation between threshold limit and rapidity anomalous dimension has been anticipated in Echevarria:2016scs ; Lustermans:2016nvk ; Billis:2019vxg . We also found that threshold limit exhibits Casimir scaling up to three loops ($n=1,2,3$),
$\lim_{z\to 1}\frac{\Delta\mathcal{I}^{(n)}_{gg}(z)}{\Delta\mathcal{I}^{(n)}_{qq}(z)}=\lim_{z\to 1}\frac{\Delta\mathcal{C}^{(n)}_{gg}(z)}{\Delta\mathcal{C}^{(n)}_{qq}(z)}=\frac{C_{A}}{C_{F}}\,.$ (149)
It is also instructive to examine the small-x limit and compare with the unpolarized case. While the unpolarized coefficient functions scale as $1/x$, the helicity-dependent ones scale as $x^{0}$. A systematic resummation of the associated logarithms is therefore of interest. Significant progress has been made toward NLL small-x resummation Catani:1994sq ; Ciafaloni:1998iv ; Salam:1998tj ; Ciafaloni:1999yw ; Ciafaloni:2003rd ; Ciafaloni:2003kd ; Ciafaloni:2003ek ; Altarelli:1999vw ; Altarelli:2001ji ; Altarelli:2003hk ; Altarelli:2005ni ; Marzani:2007gk ; Neill:2023jcd ; Kovchegov:2015pbl ; Kovchegov:2016weo ; Kovchegov:2017jxc ; Kovchegov:2017lsr . For unpolarized TMD FFs, such logarithms were resummed to all orders in $\alpha_{s}$ up to NNLL in our previous work Luo:2020epw , based on the algorithm of Ref. Vogt:2011jv . Below we collect the small-x data for the physical coefficient functions, we use superscripts ‘s’ to indicate the corresponding results are in the singlet sector
$\Delta\mathscr{I}^{\text{s}(1)}_{qq}(x)=$ $2C_{F}\,$, (150)
$\Delta\mathscr{I}^{\text{s}(2)}_{qq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{2}{3}\right)\ln^{3}x+\left(-\frac{11}{6}\right)\ln^{2}x+\left(-\frac{31}{9}\right)\ln x-12\zeta_{2}+4\zeta_{3}+\frac{397}{27}\right)$
$+$ $C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+\frac{32}{3}\ln^{2}x+\frac{344}{9}\ln x-16\zeta_{2}+\frac{2452}{27}\right)$
$+$ $C_{F}^{2}\left(\ln^{3}x+3\ln^{2}x+4\ln x+18\zeta_{2}+20\zeta_{3}-52\right)\,$, (151)
$\Delta\mathscr{I}^{\text{s}(1)}_{qg}(x)=$ $8N_{f}T_{F}\,$, (152)
$\Delta\mathscr{I}^{\text{s}(2)}_{qg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+22\ln^{2}x+72\ln x-32\zeta_{2}+8\zeta_{3}+140\right)$
$+$ $C_{F}N_{f}T_{F}\left(\frac{2}{3}\ln^{3}x+13\ln^{2}x+56\ln x-24\zeta_{2}-56\zeta_{3}+142\right)\,$, (153)
$\Delta\mathscr{I}^{\text{s}(3)}_{qg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(\frac{32308}{81}-12\zeta_{2}\right)\ln^{3}x+\left(\left(-\frac{424}{3}\right)\zeta_{2}+\left(-\frac{328}{3}\right)\zeta_{3}+\frac{19276}{9}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{38}{3}\right)\zeta_{4}+\frac{376}{3}\zeta_{2}-1024\zeta_{3}+\frac{1376872}{243}\right)\ln x+2\ln^{5}x+\frac{410}{9}\ln^{4}x+\frac{11456}{9}\zeta_{4}$
$+$ $\left(\left(-\frac{968}{3}\right)\zeta_{3}+\frac{170936}{81}\right)\zeta_{2}-560\zeta_{3}+1608\zeta_{5}-\frac{16984}{81}\bigg]$
$+$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\left(-\frac{124}{9}\right)\zeta_{2}+\frac{14821}{81}\right)\ln^{3}x+\left(\left(-\frac{458}{3}\right)\zeta_{2}-88\zeta_{3}+\frac{59860}{81}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{1288}{3}\right)\zeta_{2}+\frac{260}{3}\zeta_{4}+\frac{1496}{3}\zeta_{3}+\frac{108581}{243}\right)\ln x+\frac{32}{45}\ln^{5}x+\frac{523}{27}\ln^{4}x$
$+$ $\left(-\frac{3896}{3}\right)\zeta_{5}+\frac{4438}{9}\zeta_{4}+\frac{39656}{27}\zeta_{3}+\zeta_{2}\left(-24\zeta_{3}-\frac{80504}{81}\right)+\frac{62153}{324}\bigg]$
$+$ $C_{A}N_{f}^{2}T_{F}^{2}\bigg[\left(\frac{64}{3}\zeta_{2}-\frac{336176}{243}\right)\ln x+\left(-\frac{16}{9}\right)\ln^{4}x+\left(-\frac{3320}{81}\right)\ln^{3}x$
$+$ $\left(-\frac{3116}{9}\right)\ln^{2}x+\left(-\frac{304}{9}\right)\zeta_{4}+\frac{4640}{81}\zeta_{2}-\frac{163840}{81}\bigg]$
$+$ $C_{F}^{2}N_{f}T_{F}\bigg[\left(\left(-\frac{110}{9}\right)\zeta_{2}-\frac{187}{3}\right)\ln^{3}x+\left(\left(-\frac{467}{3}\right)\zeta_{2}+\left(-\frac{344}{3}\right)\zeta_{3}-\frac{160}{3}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{2360}{3}\right)\zeta_{3}+\left(-\frac{1360}{3}\right)\zeta_{2}+\left(-\frac{472}{3}\right)\zeta_{4}+\frac{2125}{3}\right)\ln x+\left(-\frac{8}{45}\right)\ln^{5}x$
- $6\ln^{4}x+\left(-\frac{4904}{3}\right)\zeta_{3}+\left(\frac{38}{3}\right)\zeta_{4}+\frac{3632}{3}\zeta_{5}+\left(\frac{488}{3}\zeta_{3}-750\right)\zeta_{2}+\frac{11735}{12}\bigg]$
$+$ $C_{F}N_{f}^{2}T_{F}^{2}\bigg[\left(32\zeta_{2}-\frac{127328}{81}\right)\ln^{2}x+\left(128\zeta_{2}+64\zeta_{3}-\frac{1108564}{243}\right)\ln x+\left(-\frac{16}{15}\right)\ln^{5}x$
$+$ $\left(-\frac{764}{27}\right)\ln^{4}x+\left(-\frac{23300}{81}\right)\ln^{3}x+\left(-\frac{848}{9}\right)\zeta_{4}+\frac{2416}{81}\zeta_{2}+\frac{5984}{27}\zeta_{3}-\frac{423361}{81}\bigg]\,$, (154)
$\Delta\mathscr{I}^{\text{s}(1)}_{gq}(x)=$ $-4C_{F}\,$, (155)
$\Delta\mathscr{I}^{\text{s}(2)}_{gq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{4}{3}\right)\ln^{3}x-16\ln^{2}x-50\ln x+28\zeta_{2}+48\zeta_{3}-\frac{4454}{27}\right)$
$+$ $\frac{544}{27}C_{F}N_{f}T_{F}+C_{F}^{2}\left(\left(-\frac{2}{3}\right)\ln^{3}x-10\ln^{2}x-19\ln x+23\right)\,$, (156)
$\Delta\mathscr{I}^{\text{s}(3)}_{gq}(x)=$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\frac{64}{9}\zeta_{2}+\frac{19232}{81}\right)\ln^{2}x+\left(\frac{3904}{27}\zeta_{2}+\frac{1472}{9}\zeta_{3}+\frac{110612}{243}\right)\ln x$
$+$ $\frac{32}{27}\ln^{4}x+\frac{728}{27}\ln^{3}x+\left(-\frac{304}{9}\right)\zeta_{4}+\left(-\frac{3448}{81}\right)\zeta_{2}+\frac{9296}{27}\zeta_{3}+\frac{219356}{243}\bigg]$
$+$ $C_{A}C_{F}^{2}\bigg[\left(\frac{86}{9}\zeta_{2}-\frac{11407}{81}\right)\ln^{3}x+\left(\left(-\frac{56}{3}\right)\zeta_{3}+\frac{842}{9}\zeta_{2}-\frac{67001}{162}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{4208}{9}\right)\zeta_{3}+\frac{556}{3}\zeta_{4}+\frac{7025}{27}\zeta_{2}-\frac{34130}{243}\right)\ln x+\left(-\frac{8}{9}\right)\ln^{5}x+\left(-\frac{523}{27}\right)\ln^{4}x$
$+$ $\left(-\frac{40228}{27}\right)\zeta_{3}+\left(-\frac{3212}{9}\right)\zeta_{4}+\left(\left(-\frac{256}{3}\right)\zeta_{3}-\frac{15107}{81}\right)\zeta_{2}+856\zeta_{5}+\frac{241279}{108}\bigg]$
$+$ $C_{A}^{2}C_{F}\bigg[\left(\frac{184}{9}\zeta_{2}-\frac{7852}{27}\right)\ln^{3}x+\left(\frac{1486}{9}\zeta_{2}+\frac{776}{3}\zeta_{3}-\frac{128950}{81}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{5342}{27}\right)\zeta_{2}+\frac{8120}{9}\zeta_{3}-78\zeta_{4}-\frac{811912}{243}\right)\ln x-2\ln^{5}x+\left(-\frac{973}{27}\right)\ln^{4}x$
$+$ $\left(-\frac{6214}{9}\right)\zeta_{4}+\frac{22508}{27}\zeta_{3}+\left(\frac{376}{3}\zeta_{3}-\frac{29860}{81}\right)\zeta_{2}-1736\zeta_{5}-\frac{209558}{243}\bigg]$
$+$ $C_{F}^{2}N_{f}T_{F}\bigg[\left(\frac{32}{9}\zeta_{2}+\frac{68282}{81}\right)\ln^{2}x+\left(\left(-\frac{896}{9}\right)\zeta_{3}+\left(-\frac{2176}{27}\right)\zeta_{2}+\frac{568504}{243}\right)\ln x$
$+$ $\frac{16}{15}\ln^{5}x+\frac{608}{27}\ln^{4}x+\frac{15428}{81}\ln^{3}x+\left(-\frac{14080}{27}\right)\zeta_{3}+\frac{496}{9}\zeta_{4}+\frac{10960}{81}\zeta_{2}+\frac{66509}{27}\bigg]$
$+$ $C_{F}N_{f}^{2}T_{F}^{2}\bigg[\left(-\frac{128}{3}\right)\zeta_{3}-\frac{5696}{243}\bigg]+C_{F}^{3}\bigg[\left(8\zeta_{2}+75\right)\ln^{3}x+\left(\frac{200}{3}\zeta_{2}+72\zeta_{3}+\frac{1129}{6}\right)\ln^{2}x$
$+$ $\left(\frac{220}{3}\zeta_{2}+\frac{1232}{3}\zeta_{3}-24\zeta_{4}+\frac{328}{3}\right)\ln x+\frac{16}{45}\ln^{5}x+\frac{28}{3}\ln^{4}x+\frac{1184}{3}\zeta_{4}+\frac{2516}{3}\zeta_{3}$
$+$ $\zeta_{2}\left(144\zeta_{3}+284\right)-640\zeta_{5}-\frac{12365}{12}\bigg]\,$, (157)
$\Delta\mathscr{I}^{\text{s}(1)}_{gg}(x)=$ $-8C_{A}\,$, (158)
$\Delta\mathscr{I}^{\text{s}(2)}_{gg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{2}x+\frac{136}{9}\ln x+\frac{1882}{27}\right)+C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+18\ln^{2}x+84\ln x+120\right)$
$+$ $C_{A}^{2}\left(\left(-\frac{8}{3}\right)\ln^{3}x+\left(-\frac{101}{3}\right)\ln^{2}x+\left(-\frac{1109}{9}\right)\ln x+56\zeta_{2}+96\zeta_{3}-\frac{9020}{27}\right)\,$, (159)
$\Delta\mathscr{I}^{\text{s}(3)}_{gg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(8\zeta_{2}+\frac{3416}{81}\right)\ln^{3}x+\left(\frac{488}{9}\zeta_{2}+112\zeta_{3}+\frac{10708}{27}\right)\ln^{2}x$
$+$ $\left(\frac{2992}{27}\zeta_{2}+\frac{452}{3}\zeta_{4}+\frac{6160}{9}\zeta_{3}+\frac{93668}{81}\right)\ln x+\left(-\frac{4}{15}\right)\ln^{5}x+\left(-\frac{34}{27}\right)\ln^{4}x$
$+$ $\frac{586}{9}\zeta_{4}+\left(\frac{304}{3}\zeta_{3}-\frac{5158}{81}\right)\zeta_{2}+\frac{3520}{9}\zeta_{3}-336\zeta_{5}+\frac{1899556}{729}\bigg]$
$+$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\left(-\frac{236}{9}\right)\zeta_{2}+\frac{52516}{81}\right)\ln^{3}x+\left(\left(-\frac{2242}{9}\right)\zeta_{2}-288\zeta_{3}+\frac{287876}{81}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{14528}{9}\right)\zeta_{3}+\left(-\frac{20716}{27}\right)\zeta_{2}+\left(-\frac{1664}{3}\right)\zeta_{4}+\frac{2760224}{243}\right)\ln x+\frac{16}{5}\ln^{5}x+\frac{1886}{27}\ln^{4}x$
$+$ $\left(-\frac{48208}{27}\right)\zeta_{3}+\left(-\frac{6224}{9}\right)\zeta_{4}+\frac{2192}{3}\zeta_{5}+\left(\left(-\frac{736}{3}\right)\zeta_{3}-\frac{28640}{81}\right)\zeta_{2}+\frac{3080480}{243}\bigg]$
$+$ $C_{A}N_{f}^{2}T_{F}^{2}\bigg[\left(\left(-\frac{128}{27}\right)\zeta_{2}-\frac{1024}{81}\right)\ln x+\frac{32}{81}\ln^{3}x+\frac{224}{81}\ln^{2}x+\left(-\frac{2816}{81}\right)\zeta_{2}$
$+$ $\left(-\frac{512}{27}\right)\zeta_{3}-\frac{75328}{729}\bigg]+C_{A}^{3}\bigg[\left(\frac{152}{3}\zeta_{2}-\frac{16990}{27}\right)\ln^{3}x+\left(\frac{4055}{9}\zeta_{2}+\frac{1376}{3}\zeta_{3}-\frac{278705}{81}\right)\ln^{2}x$
$+$ $\left(\frac{1559}{3}\zeta_{2}+\frac{14896}{9}\zeta_{3}+24\zeta_{4}-\frac{74936}{9}\right)\ln x+\left(-\frac{56}{15}\right)\ln^{5}x+\left(-\frac{1948}{27}\right)\ln^{4}x$
$+$ $\left(-\frac{12338}{9}\right)\zeta_{4}+\frac{41864}{27}\zeta_{3}+\left(\frac{896}{3}\zeta_{3}-\frac{16004}{81}\right)\zeta_{2}-2848\zeta_{5}-\frac{2412238}{729}\bigg]$
$+$ $C_{F}N_{f}^{2}T_{F}^{2}\bigg[\left(\left(-\frac{64}{9}\right)\zeta_{2}-\frac{2224}{81}\right)\ln^{2}x+\left(\left(-\frac{2368}{27}\right)\zeta_{2}+\left(-\frac{512}{9}\right)\zeta_{3}-\frac{30400}{243}\right)\ln x$
$+$ $\frac{8}{27}\ln^{4}x+\frac{112}{81}\ln^{3}x+\left(-\frac{16064}{81}\right)\zeta_{2}+\left(-\frac{2560}{27}\right)\zeta_{3}+\left(-\frac{416}{9}\right)\zeta_{4}-\frac{28672}{243}\bigg]$
$+$ $C_{F}^{2}N_{f}T_{F}\bigg[\left(\left(-\frac{64}{9}\right)\zeta_{2}+\frac{1420}{9}\right)\ln^{3}x+\left(\left(-\frac{224}{3}\right)\zeta_{2}+\frac{32}{3}\zeta_{3}+\frac{2320}{3}\right)\ln^{2}x$
$+$ $\left(\left(-\frac{736}{3}\right)\zeta_{2}+\left(-\frac{352}{3}\right)\zeta_{3}+\frac{784}{3}\zeta_{4}+\frac{4352}{3}\right)\ln x+\frac{8}{15}\ln^{5}x+\frac{130}{9}\ln^{4}x$
$+$ $\left(-\frac{1088}{3}\right)\zeta_{3}+\left(-\frac{608}{3}\right)\zeta_{5}+\frac{296}{3}\zeta_{4}+\left(\frac{224}{3}\zeta_{3}-\frac{1840}{3}\right)\zeta_{2}+\frac{5344}{3}\bigg]\,.$ (160)
The small-$z$ expansion of coefficient functions $\Delta\mathscr{C}_{ij}(z)$ for helicity TMD FFs is provided in the ancillary files.
5 Conclusion
We have calculated, for the first time, the complete set of N3LO twist-2 matching coefficients for helicity TMDs in both HVBM and $\overline{\text{MS}}$ schemes. For the latter, the only missing ingredient is the three-loop pure-singlet transformation factor $z^{(3)}_{\text{ps}}$, whose numerical impact is expected to be negligible in practice as $\lim_{x\to 1}z^{(3)}_{\text{ps}}(x)=0$. Our computation reproduces the NNLO space-like helicity splitting functions (up to a subtlety in the cubic color structure compared to Ref. Moch:2015usa ) and yields entirely new NNLO time-like results. With these polarized matching coefficients, we predict the spectrum of transverse momentum imbalance to N3LL accuracy in the current fragmentation region. These developments on the perturbative regime, together with first-principles Lattice QCD calculation in the nonperturbative regime Bollweg:2025ecn , provide unprecedented theoretical input for unraveling how quark and gluon spin contribute to the proton’s longitudinal spin, a central physics goal of the EIC. Looking ahead, extending these results to the small-$x$ regime, with systematic exploration of BFKL dynamics for the resummation of small-$x$ logarithms, promises new insights into spin-dependent Sudakov processes at high-energy limit of QCD.
Acknowledgements.
I thank Hua Xing Zhu for encouraging me to look into this problem and Tong-Zhi Yang for collaboration in the early stages of the project.
Appendix A Flavor decomposition
In this section, we fix our notations for flavor group decomposition of the coefficient functions, the PDFs (FFs) and the splitting functions, our definitions are in consistence with those in Vogt:2004ns . We begin with the non-singlet sectors, denoting $q$ to be a quark, $\bar{q}$ an anti-quark and $q^{\prime}$ another flavor of quark, the ‘valence’ quark contributions are defined as ($f$ can be identified as partonic PDFs, the coefficient functions or the splitting functions of some flavor)
$f^{\text{v}}_{qq}=f_{qq}-f_{qq^{\prime}}\,,\quad f^{\text{v}}_{q\bar{q}}=f_{q\bar{q}}-f_{q\bar{q}^{\prime}}\,.$ (161)
By linear combinations, we define three independently evolving types of non-singlets
$f^{+}_{\text{ns}}=f^{\text{v}}_{qq}+f^{\text{v}}_{q\bar{q}}\,,\quad f^{-}_{\text{ns}}=f^{\text{v}}_{qq}-f^{\text{v}}_{q\bar{q}}\,,\quad f^{\text{v}}_{\text{ns}}=f^{-}_{\text{ns}}+N_{f}(f_{qq^{\prime}}-f_{q\bar{q}^{\prime}})\,$, (162)
note that the ‘sea’ quark difference $f_{qq^{\prime}}-f_{q\bar{q}^{\prime}}$ vanishes up to $\alpha_{s}^{2}$, and is non-vanishing from the third order. In particular, we denote the ‘sea’ quark difference as
$f_{d_{abc}^{2}}=N_{f}(f_{qq^{\prime}}-f_{q\bar{q}^{\prime}})\,$, (163)
since at 3-loop, they are proportional to the cubic color structure $d_{abc}^{2}$. For $i=\pm,\text{v}$, the evolutions of the non-singlets are governed by
$\frac{d}{d\ln\mu^{2}}f^{i}_{\text{ns}}=\Delta P^{i}_{\text{ns}}\otimes f^{i}_{\text{ns}}(\mu)\,.$ (164)
For the singlet sector, we define
$f^{\text{s}}_{qq}=$ $f^{+}_{\text{ns}}+N_{f}(f_{qq^{\prime}}+f_{q\bar{q}^{\prime}})\equiv f^{+}_{\text{ns}}+f^{\text{ps}}_{qq}\,$,
$f^{\text{s}}_{qg}=$ $2N_{f}f_{qg}\,,\quad f^{\text{s}}_{gq}=f_{gq}\,,\quad f^{\text{s}}_{gg}=f_{gg}\,.$ (165)
It proves useful to use a matrix notation for the partonic PDFs (FFs)
${\bf f}^{\text{s}}=\left(\begin{array}[]{cc}f^{\text{s}}_{qq}&f^{\text{s}}_{qg}\\ f^{\text{s}}_{gq}&f^{\text{s}}_{gg}\\ \end{array}\right)\,$, (168)
and the corresponding $2\times 2$ evolution system reads
$\frac{d}{d\ln\mu^{2}}{\bf f}^{s}(\mu)={\bf P}^{\text{s}}\otimes{\bf f}^{\text{s}}(\mu)\,.$ (169)
Appendix B Renormalization Group Consistency
From RG invariance of the cross section
$\frac{\mathrm{d}}{\mathrm{d}\ln\mu}\left[H(\alpha_{s}(\mu),L_{Q})\otimes f_{1}^{q}(x,\alpha_{s}(\mu),L_{b},L_{\xi_{n}})\otimes D_{1}^{q}(z,\alpha_{s}(\mu),L_{b},L_{\xi_{\bar{n}}})\right]\equiv 0\,$, (170)
we must have
$\Gamma_{H}(\alpha_{s}(\mu),L_{\xi_{n}}+L_{\xi_{\bar{n}}})=\Gamma_{f}(\alpha_{s}(\mu),L_{b},L_{\xi_{n}})+\Gamma_{D}(\alpha_{s}(\mu),L_{b},L_{\xi_{\bar{n}}}),$ (171)
where we have used physical TMDs for the factorization formula eq. (2), and the relation $2L_{Q}=L_{\xi_{n}}+L_{\xi_{\bar{n}}}$ for the Collins-Soper scales in eq. (48).
In RG relation eq. (171), we can set either $L_{\xi_{n}}$ or $L_{\xi_{\bar{n}}}$ to zero, this leads to the following relation for $\Gamma_{H}$
$\Gamma_{H}(\alpha_{s}(\mu),L_{\xi_{n}}+L_{\xi_{\bar{n}}})=\Gamma_{H}(\alpha_{s}(\mu),L_{\xi_{n}})+\Gamma_{H}(\alpha_{s}(\mu),L_{\xi_{\bar{n}}})-\Gamma_{H}(\alpha_{s}(\mu),0)\,.$ (172)
We can perform a redefinition by a shift $\overline{\Gamma}_{H}:=\Gamma_{H}-\Gamma_{H}(0)$, then additive property follows
$\overline{\Gamma}_{H}(x+y)=\overline{\Gamma}_{H}(x)+\overline{\Gamma}_{H}(y)\,.$ (173)
The additive property guarantees that $\overline{\Gamma}_{H}(Nx)=N\times\overline{\Gamma}_{H}(x)\,,\forall N\in\mathbb{Z}$. We only need to show homogeneity holds for all $c\in\mathbb{R}$ (or $c\in\mathbb{C}$). This is true as long as $\overline{\Gamma}_{H}(c)=c\times\overline{\Gamma}_{H}(1)$ holds for all $c\in\mathbb{R}$ (or $c\in\mathbb{C}$). Since $\overline{\Gamma}_{H}$ is assumed to be a continuous function of its argument and the real (complex) numbers are nothing but the Cauchy sequences of (pairs of) the rationals, it suffice to prove it for the rationals. Since we already have additivity for the integers, it suffice to prove it for unit fractions. Indeed, the following is true
$\overline{\Gamma}_{H}(1/M)=1/M\times M\times\overline{\Gamma}_{H}(1/M)=1/M\times\overline{\Gamma}_{H}(1)\,,\quad\forall M\in\mathbb{Z}\,$, (174)
which complete the proof that $\overline{\Gamma}_{H}$ is a linear function. From this we conclude that the anomalous dimensions associated with $\mu$-evolutions is linear in $\ln\mu$ to all-loop orders, e.g. ,
$\frac{\mathrm{d}}{\mathrm{d}\ln\mu^{2}}\ln f_{i}(x,\alpha_{s}(\mu),L_{b},L_{\xi_{i}})=-\gamma^{i}_{\rm cusp}(\alpha_{s}(\mu))L_{\xi_{i}}-\gamma^{i}(\alpha_{s}(\mu),\dots)\,.$ (175)
On the other hand, the rapidity evolution with respect to the Collins-Soper scale $\xi_{i}$ is
$\Gamma_{R}^{i}(L_{\xi_{i}},L_{b})\equiv\frac{\mathrm{d}}{\mathrm{d}L_{\xi_{i}}}\ln f_{i}(x,\alpha_{s}(\mu),L_{b},L_{\xi_{i}})\,$, (176)
and $\mu$-evolution of the Collins-Soper kernel is controlled by the cusp anomalous dimension
$\frac{\mathrm{d}}{\mathrm{d}\ln\mu^{2}}\Gamma_{R}^{i}(L_{\xi_{i}},L_{b})=\frac{\mathrm{d}}{\mathrm{d}L_{\xi_{i}}}\frac{\mathrm{d}}{\mathrm{d}\ln\mu^{2}}\ln f_{i}(x,\alpha_{s}(\mu),L_{b},L_{\xi_{i}})=-\gamma^{i}_{\rm cusp}(\alpha_{s}(\mu))\,$, (177)
thus the Collins-Soper kernel takes the form
$\Gamma_{R}^{i}(L_{\xi_{i}},L_{b})=-\int^{\mu^{2}}_{\mu_{b}^{2}}\frac{d\bar{\mu}^{2}}{\bar{\mu}^{2}}\gamma^{i}_{\rm cusp}[\alpha_{s}(\bar{\mu})]+\gamma^{i}_{R}[\alpha_{s}(\mu_{b})]\,.$ (178)
Appendix C QCD Beta Function
The QCD beta function is defined as
$\frac{d\alpha_{s}}{d\ln\mu}=\beta(\alpha_{s})=-2\alpha_{s}\sum_{n=0}^{\infty}\left(\frac{\alpha_{s}}{4\pi}\right)^{n+1}\,\beta_{n}\,$, (179)
with Baikov:2016tgj
$\beta_{0}$ $=\frac{11}{3}C_{A}-\frac{4}{3}T_{F}N_{f}\,$,
$\beta_{1}$ $=\frac{34}{3}C_{A}^{2}-\frac{20}{3}C_{A}T_{F}N_{f}-4C_{F}T_{F}N_{f}\,$,
$\beta_{2}$ $=\left(\frac{158C_{A}}{27}+\frac{44C_{F}}{9}\right)N_{f}^{2}T_{F}^{2}+\left(-\frac{205C_{A}C_{F}}{9}-\frac{1415C_{A}^{2}}{27}+2C_{F}^{2}\right)N_{f}T_{F}+\frac{2857C_{A}^{3}}{54}\,.$ (180)
Appendix D Anomalous dimensions
For all the anomalous dimensions entering the renormalization group equations of various TMD functions, we define the perturbative expansion in $\alpha_{s}$ according to
$\gamma(\alpha_{s})=\sum_{n=0}^{\infty}\left(\frac{\alpha_{s}}{4\pi}\right)^{n+1}\,\gamma_{n}\,$, (181)
where the coefficients for quark are given by
$\Gamma^{\text{cusp}}_{0}=$ $4C_{F}\,$,
$\Gamma^{\text{cusp}}_{1}=$ $\left(\frac{268}{9}-8\zeta_{2}\right)C_{A}C_{F}-\frac{80C_{F}T_{F}N_{f}}{9}\,$,
$\Gamma^{\text{cusp}}_{2}=$ $\bigg[\left(\frac{320\zeta_{2}}{9}-\frac{224\zeta_{3}}{3}-\frac{1672}{27}\right)C_{A}C_{F}+\left(64\zeta_{3}-\frac{220}{3}\right)C_{F}^{2}\bigg]N_{f}T_{F}$
$+$ $\left(-\frac{1072\zeta_{2}}{9}+\frac{88\zeta_{3}}{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{2}C_{F}-\frac{64}{27}C_{F}N_{f}^{2}T_{F}^{2}\,$,
$\gamma^{S}_{0}=$ $0\,$,
$\gamma^{S}_{1}=$ $\left[\left(-\frac{404}{27}+\frac{11\zeta_{2}}{3}+14\zeta_{3}\right)C_{A}+\left(\frac{112}{27}-\frac{4\zeta_{2}}{3}\right)T_{F}N_{f}\right]C_{F}\,$,
$\gamma^{S}_{2}=$ $\left(-\frac{88}{3}\zeta_{3}\zeta_{2}+\frac{6325\zeta_{2}}{81}+\frac{658\zeta_{3}}{3}-88\zeta_{4}-96\zeta_{5}-\frac{136781}{1458}\right)C_{A}^{2}C_{F}+\left(\frac{80\zeta_{2}}{27}-\frac{224\zeta_{3}}{27}\right$.
$+$ $\left.\frac{4160}{729}\right)C_{F}N_{f}^{2}T_{F}^{2}+\left(-\frac{2828\zeta_{2}}{81}-\frac{728\zeta_{3}}{27}+48\zeta_{4}+\frac{11842}{729}\right)C_{A}C_{F}N_{f}T_{F}$
$+$ $\left(-4\zeta_{2}-\frac{304\zeta_{3}}{9}-16\zeta_{4}+\frac{1711}{27}\right)C_{F}^{2}N_{f}T_{F}\,.$
$\gamma^{R}_{0}=$ $0\,$,
$\gamma^{R}_{1}=$ $\left[\left(-\frac{404}{27}+14\zeta_{3}\right)C_{A}+\frac{112}{27}T_{F}N_{f}\right]C_{F}\,$,
$\gamma^{R}_{2}=$ $\bigg[\left(-\frac{824\zeta_{2}}{81}-\frac{904\zeta_{3}}{27}+\frac{20\zeta_{4}}{3}+\frac{62626}{729}\right)C_{A}N_{f}T_{F}+\left(-\frac{88}{3}\zeta_{3}\zeta_{2}+\frac{3196\zeta_{2}}{81}+\frac{6164\zeta_{3}}{27}\right$.
$+$ $\left.\frac{77\zeta_{4}}{3}-96\zeta_{5}-\frac{297029}{1458}\right)C_{A}^{2}+\left(-\frac{304\zeta_{3}}{9}-16\zeta_{4}+\frac{1711}{27}\right)C_{F}N_{f}T_{F}+\left(-\frac{64\zeta_{3}}{9}\right$.
- $\left.\frac{3712}{729}\right)N_{f}^{2}T_{F}^{2}\bigg]C_{F}\,.$ (182)
Since cusp and soft and rapidity anomalous dimensions exhibit Casimir scaling, the corresponding anomalous dimensions for gluon could be obtained by multiplying from above by $C_{A}/C_{F}$.
The beam anomalous dimensions do not exhibits Casimir scaling, thus should be list separately. The beam anomalous dimensions for quark are
$\gamma^{B}_{0}=$ $3C_{F}\,$,
$\gamma^{B}_{1}=$ $\left[\left(\frac{3}{2}-12\zeta_{2}+24\zeta_{3}\right)C_{F}+\left(\frac{17}{6}+\frac{44\zeta_{2}}{3}-12\zeta_{3}\right)C_{A}+\left(-\frac{2}{3}-\frac{16\zeta_{2}}{3}\right)T_{F}N_{f}\right]C_{F},$
$\gamma^{B}_{2}=$ $\bigg[\left(-\frac{2672\zeta_{2}}{27}+\frac{400\zeta_{3}}{9}+4\zeta_{4}+40\right)C_{A}C_{F}+\left(\frac{40\zeta_{2}}{3}-\frac{272\zeta_{3}}{3}+\frac{232\zeta_{4}}{3}-46\right)C_{F}^{2}\bigg]N_{f}T_{F}$
$+$ $\left(16\zeta_{3}\zeta_{2}-\frac{410\zeta_{2}}{3}+\frac{844\zeta_{3}}{3}-\frac{494\zeta_{4}}{3}+120\zeta_{5}+\frac{151}{4}\right)C_{A}C_{F}^{2}+\left(\frac{320\zeta_{2}}{27}-\frac{64\zeta_{3}}{9}-\frac{68}{9}\right)$
$\times$ $C_{F}N_{f}^{2}T_{F}^{2}+\left(\frac{4496\zeta_{2}}{27}-\frac{1552\zeta_{3}}{9}-5\zeta_{4}+40\zeta_{5}-\frac{1657}{36}\right)C_{A}^{2}C_{F}+\bigg(-32\zeta_{3}\zeta_{2}+18\zeta_{2}$
$+$ $68\zeta_{3}+144\zeta_{4}-240\zeta_{5}+\frac{29}{2}\bigg)C_{F}^{3}\,.$ (183)
The beam anomalous dimensions for gluon are
$\gamma^{B}_{0}=$ $\frac{11}{3}C_{A}-\frac{4}{3}T_{F}N_{f}\,$,
$\gamma^{B}_{1}=$ $C_{A}^{2}\left(\frac{32}{3}+12\zeta_{3}\right)+\left(-\frac{16}{3}C_{A}-4C_{F}\right)N_{f}T_{F}\,$,
$\gamma^{B}_{2}=$ $C_{A}^{3}\left(-80\zeta_{5}-16\zeta_{3}\zeta_{2}+\frac{55}{3}\zeta_{4}+\frac{536}{3}\zeta_{3}+\frac{8}{3}\zeta_{2}+\frac{79}{2}\right)$
$+$ $C_{A}^{2}N_{f}T_{F}\left(-\frac{20}{3}\zeta_{4}-\frac{160}{3}\zeta_{3}-\frac{16}{3}\zeta_{2}-\frac{233}{9}\right)+\frac{58}{9}C_{A}N_{f}^{2}T_{F}^{2}-\frac{241}{9}C_{A}C_{F}N_{f}T_{F}$
$+$ $2C_{F}^{2}N_{f}T_{F}+\frac{44}{9}C_{F}N_{f}^{2}T_{F}^{2}\,.$ (184)
The cusp anomalous dimension $\Gamma^{\text{cusp}}$ can be found in Moch:2004pa . The beam anomalous dimension $\gamma^{B}$ is related to the soft anomalous dimension $\gamma^{S}$ Li:2014afw and the hard anomalous dimensions $\gamma^{H}$ Moch:2005tm ; Gehrmann:2010ue ; Becher:2009qa by renormalization group invariance condition $\gamma^{B}=\gamma^{S}-\gamma^{H}$. The rapidity anomalous dimension $\gamma^{R}$ can be found in Li:2016ctv ; Vladimirov:2016dll . Note that the normalization here differ from those in Li:2016ctv by a factor of $1/2$.
Appendix E Renormalization Constants
The following constants are needed for the renormalization of zero-bin subtracted Manohar:2006nz TMD PDFs through N3LO, see e.g. Ref. Luo:2019hmp ; Luo:2019bmw . The first three-order corrections to $Z^{B}$ and $Z^{S}$ are
$Z^{B}_{1}=$ $\frac{1}{2\epsilon}\left(2\gamma^{B}_{0}-\Gamma_{0}^{\text{cusp}}L_{Q}\right)\,$,
$Z^{B}_{2}=$ $\frac{1}{8\epsilon^{2}}\bigg((\Gamma_{0}^{\text{cusp}}L_{Q}-2\gamma^{B}_{0})^{2}+2\beta_{0}(\Gamma_{0}^{\text{cusp}}L_{Q}-2\gamma^{B}_{0})\bigg)+\frac{1}{4\epsilon}\left(2\gamma^{B}_{1}-\Gamma_{1}^{\text{cusp}}L_{Q}\right)\,$,
$Z^{B}_{3}=$ $\frac{1}{48\epsilon^{3}}\left(2\gamma^{B}_{0}-\Gamma^{\text{cusp}}_{0}L_{Q}\right)\biggl(8\beta_{0}^{2}+6\beta_{0}\left(-2\gamma^{B}_{0}+\Gamma^{\text{cusp}}_{0}L_{Q}\right)+\left(-2\gamma^{B}_{0}+\Gamma^{\text{cusp}}_{0}L_{Q}\right)^{2}\biggl)$
$+$ $\frac{1}{24\epsilon^{2}}\biggl(\beta_{1}\left(-8\gamma^{B}_{0}+4\Gamma^{\text{cusp}}_{0}L_{Q}\right)+\left(4\beta_{0}-6\gamma^{B}_{0}+3\Gamma^{\text{cusp}}_{0}L_{Q}\right)\left(-2\gamma^{B}_{1}+\Gamma^{\text{cusp}}_{1}L_{Q}\right)\biggl)$
$+$ $\frac{1}{6\epsilon}\biggl(2\gamma^{B}_{2}-\Gamma^{\text{cusp}}_{2}L_{Q}\biggl)$
$Z^{S}_{1}=$ $\frac{1}{\epsilon^{2}}\Gamma^{\text{cusp}}_{0}+\frac{1}{\epsilon}\left(-2\gamma^{S}_{0}-\Gamma_{0}^{\text{cusp}}L_{\nu}\right)\,$,
$Z^{S}_{2}=$ $\frac{1}{2\epsilon^{4}}(\Gamma^{\text{cusp}}_{0})^{2}-\frac{1}{4\epsilon^{3}}\bigg(\Gamma^{\text{cusp}}_{0}(3\beta_{0}+8\gamma^{S}_{0})+4(\Gamma^{\text{cusp}}_{0})^{2}L_{\nu}\bigg)-\frac{1}{2\epsilon}\left(2\gamma^{S}_{1}+\Gamma^{\text{cusp}}_{1}L_{\nu}\right)$
$+$ $\frac{1}{4\epsilon^{2}}\bigg(\Gamma^{\text{cusp}}_{1}+2(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu})(\beta_{0}+2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu})\bigg)\,$,
$Z^{S}_{3}=$ $\frac{1}{6\epsilon^{6}}\left(\Gamma^{\text{cusp}}_{0}\right)^{3}-\frac{1}{4\epsilon^{5}}\left(\Gamma^{\text{cusp}}_{0}\right)^{2}\left(3\beta_{0}+4\gamma^{S}_{0}+2\Gamma^{\text{cusp}}_{0}L_{\nu}\right)+\frac{1}{36\epsilon^{4}}\Gamma^{\text{cusp}}_{0}\bigg(22\beta_{0}^{2}+45\beta_{0}\left(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu}\right)$
$+$ $9\left(\Gamma^{\text{cusp}}_{1}+2\left(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu}\right)^{2}\right)\biggl)+\frac{1}{36\epsilon^{3}}\biggl(-16\beta_{1}\Gamma^{\text{cusp}}_{0}-12\beta_{0}^{2}\left(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu}\right)$
- $2\beta_{0}\left(5\Gamma^{\text{cusp}}_{1}+9\left(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu}\right)^{2}\right)-3\bigg[\Gamma^{\text{cusp}}_{1}\left(6\gamma^{S}_{0}+9\Gamma^{\text{cusp}}_{0}L_{\nu}\right)$
$+$ $2\left(8\left(\gamma^{S}_{0}\right)^{3}+6\Gamma^{\text{cusp}}_{0}\gamma^{S}_{1}+12\Gamma^{\text{cusp}}_{0}\left(\gamma^{S}_{0}\right)^{2}L_{\nu}+6\left(\Gamma^{\text{cusp}}_{0}\right)^{2}\gamma^{S}_{0}L_{\nu}^{2}+\left(\Gamma^{\text{cusp}}_{0}\right)^{3}L_{\nu}^{3}\right)\bigg]\biggl)$
$+$ $\frac{1}{18\epsilon^{2}}\biggl(2\Gamma^{\text{cusp}}_{2}+3\left(2\beta_{1}\left(2\gamma^{S}_{0}+\Gamma^{\text{cusp}}_{0}L_{\nu}\right)+\left(2\beta_{0}+6\gamma^{S}_{0}+3\Gamma^{\text{cusp}}_{0}L_{\nu}\right)\left(2\gamma^{S}_{1}+\Gamma^{\text{cusp}}_{1}L_{\nu}\right)\right)\biggl)$
- $\frac{2\gamma^{S}_{2}+\Gamma^{\text{cusp}}_{2}L_{\nu}}{3\epsilon}\,.$ (185)
Keep in mind that the anomalous dimensions appeared above depends on the flavor, they should be replaced by the corresponding values in Sec D. We also remind the reader that the renormalization constants are formally identical for TMD PDFs and TMD FFs, the logarithms appeared above should be replaced by their corresponding values in each case, and we have
$L_{\perp}=\ln\frac{b_{T}^{2}\mu^{2}}{b_{0}^{2}},\quad L_{\nu}=\ln\frac{\nu^{2}}{\mu^{2}}\,$, (186)
with $b_{0}=2\,e^{-\gamma_{E}}$ for both TMD PDFs and TMD FFs. On the other hand, we have for TMD PDFs
$L_{Q}=2\ln\frac{x\,P_{+}}{\nu}\,$, (187)
while for TMD FFs,
$L_{Q}=2\ln\frac{P_{+}}{z\,\nu}.$ (188)

### v4

@@
 To the IR, long-distance physics is parametrized while short-distance dynamics are computed from Feynman amplitudes, again thanks to the asymptotic freedom.
-UV renormalization and IR factorization scheme ambiguities drop out in the relations among physical quantities.
+UV and IR cutoff dependence drop out in the relations among physical quantities.
 In the context of TMDs, factorization provides the gateway that bridges partonic high-energy interactions to what is observed at the detectors: the confined motion can never been probed directly, it manifests through correlations with the distribution of final-state gauge bosons, leptons, and hadrons, owing to the unification of Standard Model gauge group ${\text{SU}}(3)\times{\text{SU}}(2)\times{\text{U}}(1)_{\text{Y}}~({\text{SU}}(2)\times{\text{U}}(1)_{\text{Y}}\to\text{U}(1)_{\text{EM}}$.
@@
 […]
-Once the hard and soft-collinear degrees of freedom are disentangled, the next step is to enumerate all allowed hard events, and within each hard label, zoom in to the soft-collinear subprocesses and count over particles in a proper way: the classes of momenta entering into the soft-collinear phase space differ only by their typical size of energy or rapidity, as a result, the phase space should be integrated over with a rapidity cut-off $\nu$, here we employ exponential cut-off scheme Li:2016axz .
+Once the hard and soft-collinear degrees of freedom are disentangled, the next step is to enumerate all allowed hard events, and within each hard label, zoom in to the soft-collinear subprocesses and count over particles in a proper way: the classes of momenta entering into the soft-collinear phase space differ only by their typical size of energy or rapidity, as a result, the phase space should be integrated over with a rapidity cutoff $\nu$, here we employ exponential cutoff scheme Li:2016axz .
 Additionally, a zero-bin subtraction is performed to remove double-counting between collinear and soft sectors, yielding genuine collinear beam functions Manohar:2006nz .
@@
 (3) was pioneered by Collins, Soper, and Sterman (CSS) Collins:1981uk ; Collins:1981uw ; Collins:1984kg and expanded in Collins:2011zzd ; Ji:2004wu ; deFlorian:2001zd ; Catani:2000vq ; Catani:2010pd ; Aybat:2011zv ; Ji:2005nu , and later reformulated in the framework of Soft Collinear Effective Theory Becher:2010tm ; Becher:2011xn ; Becher:2012yn ; Echevarria:2011epo ; Echevarria:2012js ; Echevarria:2014rua ; Ebert:2019tvc ; Collins:2017oxh .
+The N3LL predictions for eq.
+(3) is shown in fig.
+(1,2) for both unpolarized and polarized distributions, where we incorporate the non-perturbative model given in Ref Sun:2014dqm
 Figure 1: Resummed $q_{\ast}$ distributions at EIC energies.
@@
 Here, all the results are implemented with $b_{\ast}$ prescription, for otherwise the integration diverges.
-The N3LL predictions for eq. (3) is shown in fig. (1,2) for both unpolarized and polarized distributions, where we incorporate the non-perturbative model given in Ref Sun:2014dqm
 $S_{\text{NP}}^{\text{SIDIS}}=\left({g_{1}\over 2}+{g_{h}\over z_{h}^{2}}\right)b^{2}+g_{2}\ln\left(\frac{b}{b_{\ast}}\right)\ln\left(\frac{Q}{Q_{0}}\right)\,$,
@@
 Furthermore, the integral in Eq.
-(63) is properly defined only with a rapidity cutoff implemented through the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$.
+(63) requires a rapidity cutoff to be well defined.
+In this work, we implement such a cutoff via the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$.
 The TMD FFs are defined as crossings of TMD beam functions, and according to parton-hardron frame relation given in eq.
@@
 4.4.1 Space-like results in $\overline{\text{MS}}$}
-The space-like helicity dependent splitting functions were obtained in Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , by comparison we find all agreement except for the term with cubic color structure $d_{abc}^{2}$ for ‘sea’ quark difference $N_{f}(\Delta P_{qq^{\prime}}-\Delta P_{q\bar{q}^{\prime}})$:  [^1] Our results agree at transcendental weight-4, but discrepancies appear for the weight-3 terms, indicating differences of $\gamma_{5}$ prescriptions between the one used in Ref Moch:2015usa and our tensor reduction prescription following Ravindran:2003gi .
-$\Delta P_{d_{abc}^{2}}^{S,(2)}-\Delta P_{d_{abc}^{2}}^{\text{Moch:2015}}=\frac{16N_{f}}{3}\frac{d_{abc}^{2}}{N_{c}}\bigg[-18\zeta_{3}+3H_{0}+84H_{2}+36H_{3}+12H_{-2,0}+(1-z)\times\left(174H_{1}-12H_{1,0,0}-84\zeta_{2}\right)+(1+z)\times\left(36H_{-1,0}-48H_{-1,2}-24H_{-1,0,0}+48H_{-1}\zeta_{2}-36H_{0}\zeta_{2}\right)+z\times\left(-174H_{0}+90H_{2}+12H_{3}-24H_{-2,0}+54H_{0,0}+24H_{0,0,0}-54\zeta_{2}-72\zeta_{3}\right)\bigg]\,.$ (122)
+The space-like helicity dependent splitting functions were obtained in Moch:2014sna ; Moch:2015usa ; Blumlein:2021enk ; Blumlein:2021ryt ; Blumlein:2022gpp , by comparison we find all agreement except for the term with cubic color structure $d_{abc}^{2}$ for ‘sea’ quark difference $N_{f}(\Delta P_{qq^{\prime}}-\Delta P_{q\bar{q}^{\prime}})$:  [^1] Agreement is found at transcendental weight four, whereas discrepancies arise in the weight-three terms.
+The author is also grateful to the authors of Ref.
+Behring:2025avs for pointing out a typo in the $\zeta_{2}$ term in the output of this formula.
+Once corrected, the expression is in full agreement with the file dPSLarin.m accompanying the present work and with their independent calculations.
+$\Delta P_{d_{abc}^{2}}^{S,(2)}-\Delta P_{d_{abc}^{2}}^{\text{Moch:2015}}=\frac{16N_{f}}{3}\frac{d_{abc}^{2}}{N_{c}}\bigg[-84\zeta_{2}-18\zeta_{3}+3H_{0}+84H_{2}+36H_{3}+12H_{-2,0}+(1-z)\times\left(174H_{1}-12H_{1,0,0}\right)+(1+z)\times\left(36H_{-1,0}-48H_{-1,2}-24H_{-1,0,0}+48H_{-1}\zeta_{2}-36H_{0}\zeta_{2}\right)+z\times\left(-174H_{0}+90H_{2}+12H_{3}-24H_{-2,0}+54H_{0,0}+24H_{0,0,0}-54\zeta_{2}-72\zeta_{3}\right)\bigg]\,.$ (122)
 It’s interesting to investigate the first moment of the helicity splitting functions, up to $\text{N}^{3}$LO we found
@@
 $\Delta P_{qq}^{T,\text{s}(1)}=\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\frac{268}{9}-8\zeta_{2}\right)C_{A}C_{F}-\frac{40C_{F}N_{f}}{9}\bigg]+C_{A}C_{F}\bigg[\frac{z^{2}}{z+1}\left(-8H_{-1,0}-\frac{446}{9}\right)+\left(-\frac{8z^{2}}{(z-1)(z+1)}\right)H_{0,0}+\frac{1}{z+1}\left(\frac{178}{9}-8H_{-1,0}\right)+\left(-\frac{8}{(z-1)(z+1)}\right)H_{0,0}+\frac{2z^{2}}{3(z-1)}H_{0}+\left(-\frac{46}{3(z-1)}\right)H_{0}+\frac{z\left(8\zeta_{2}-\frac{268}{9}\right)}{z+1}\bigg]+C_{F}N_{f}\bigg[z\left(8H_{0,0}+\frac{440}{9}\right)+8H_{0,0}+\left(-\frac{80z^{2}}{3(z-1)}\right)H_{0}+\frac{40}{3(z-1)}H_{0}+\frac{16z}{z-1}H_{0}-\frac{400}{9}\bigg]+C_{F}^{2}\bigg[\frac{12z^{3}}{(z-1)(z+1)}H_{0,0}+\frac{z^{2}}{z+1}\left(16H_{-1,0}+8\zeta_{2}+36\right)+\frac{z^{2}}{z-1}\left(8H_{0,1}+8H_{1,0}-20H_{0}\right)+\frac{28z^{2}}{(z-1)(z+1)}H_{0,0}+\frac{1}{z+1}\left(16H_{-1,0}+8\zeta_{2}-36\right)+\frac{1}{z-1}\left(8H_{0,1}+8H_{1,0}+16H_{0}\right)+\frac{20}{(z-1)(z+1)}H_{0,0}+\frac{4z}{(z-1)(z+1)}H_{0,0}+\left(-\frac{8z}{z-1}\right)H_{0}\bigg]+\delta(1-z)\bigg[\left(\frac{44}{3}\zeta_{2}-12\zeta_{3}+\frac{17}{6}\right)C_{A}C_{F}+\left(\left(-\frac{8}{3}\right)\zeta_{2}-\frac{1}{3}\right)C_{F}N_{f}+\left(-12\zeta_{2}+24\zeta_{3}+\frac{3}{2}\right)C_{F}^{2}\bigg]\,$, (127)
-$\Delta P_{qq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\frac{1}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}-56H_{0,1,0}+96H_{1,0,0}-208H_{0,0,0,1}-64H_{0,0,1,1}-48H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z}{z-1}\left(-192\zeta_{2}H_{1}-800H_{1}+64H_{0,1,0}\right)+\frac{z}{z-1}\left(-144\zeta_{2}H_{0}-72H_{0}-12\zeta_{2}-192\zeta_{3}-324\zeta_{4}-352\zeta_{2}H_{0,-1}+144\zeta_{2}H_{0,0}+176H_{0,0}-96H_{0,0,0}-64H_{0,-1,-1,0}+128H_{0,-1,0,0}-96H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(-16\zeta_{2}H_{0}+128\zeta_{3}H_{0}+10H_{0}+204\zeta_{2}+48\zeta_{3}+252\zeta_{4}+160\zeta_{2}H_{0,-1}+208\zeta_{2}H_{0,0}+172H_{0,0}+528H_{0,0,0}-320H_{0,-1,-1,0}+192H_{0,-1,0,0}+288H_{0,0,-1,0}-592H_{0,0,0,0}-192H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}+88H_{0,1,0}+96H_{1,0,0}-240H_{0,0,0,1}-64H_{0,0,1,1}-144H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z^{3}}{z-1}\left(432\zeta_{2}H_{0}+114H_{0}+12\zeta_{2}+336\zeta_{3}-28\zeta_{4}-352\zeta_{2}H_{0,-1}+48\zeta_{2}H_{0,0}-228H_{0,0}+240H_{0,0,0}-64H_{0,-1,-1,0}-160H_{0,0,-1,0}-112H_{0,0,0,0}-128H_{0,0,1,0}+\frac{1}{z+1}+302\right)+\frac{1}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-12H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-64H_{0,-1,0}+64H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(16\zeta_{2}H_{0}-64\zeta_{3}H_{0}-64H_{0}-204\zeta_{2}-192\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}-224H_{0,0}-96H_{0,0,0}-320H_{0,-1,-1,0}+64H_{0,-1,0,0}+96H_{0,0,-1,0}-176H_{0,0,0,0}-128H_{0,0,1,0}+302\right)+\frac{z}{z+1}\left(-576H_{-1}\zeta_{2}-192H_{-1,0}-312H_{0,1}-64H_{-1,0,0}+768H_{-1,0,1}-320H_{0,-1,0}-96H_{0,0,1}\right)+\frac{z^{2}}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-300H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-160H_{0,-1,0}-256H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+144H_{1,0}+192H_{-1,-1,0}+z\left(192H_{-1,-1,0}-144H_{1,0}\right)\bigg]+C_{F}^{2}N_{f}\bigg[\frac{668}{3}H_{1}+\frac{1}{z-1}\left(\left(-\frac{2104}{9}\right)H_{0,1}+\left(-\frac{464}{3}\right)H_{0,1,0}+\left(-\frac{580}{9}\right)H_{1,0}-160\zeta_{2}H_{0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{5048}{9}\right)H_{0,0}+\left(-\frac{496}{3}\right)H_{0,0,1}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{4772}{9}\zeta_{2}+\frac{7868}{9}H_{0}+\left(-\frac{112\zeta_{2}}{3}\right)H_{0}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-152H_{0,0,0}+\frac{1}{z+1}-\frac{1727}{3}\right)+\frac{z}{z-1}\left(\frac{220}{3}H_{0,1}+\frac{280}{3}H_{1,0}+144H_{0,1,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{580}{9}\right)H_{1,0}+\left(-\frac{32}{3}\right)H_{0,1,0}+\frac{1124}{9}H_{0,1}+160\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{80}{3}\right)\zeta_{3}+\frac{20}{9}H_{0}+\frac{56}{3}H_{0,0,0}+\frac{760}{9}H_{0,0}+\frac{1672}{9}\zeta_{2}-64H_{0}\zeta_{2}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+64H_{0,0,1}+\frac{1}{z+1}-539\right)$ (128)
-$\frac{z^{3}}{z-1}\left(\left(-\frac{7664}{9}\right)H_{0}+\left(-\frac{5092}{9}\right)\zeta_{2}+\frac{112}{3}\zeta_{3}+\frac{368}{3}H_{0,0,1}+\frac{1624}{3}H_{0,0}+\frac{176\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+120H_{0,0,0}+\frac{1}{z+1}+\frac{1727}{3}\right)+\frac{1}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{1576}{9}\right)H_{0,0}+\left(-\frac{1352}{9}\right)\zeta_{2}+\left(-\frac{280}{3}\right)H_{0,0,0}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{184}{9}H_{0}+64H_{0}\zeta_{2}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-64H_{0,0,1}+539\right)+\left(-\frac{128z}{3(z+1)}\right)H_{-1,0}+\frac{z^{2}}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)+400H_{1}\zeta_{2}-52H_{1,1}+48H_{0,1,1}-440H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}+z\left(\left(-\frac{668}{3}\right)H_{1}-400H_{1}\zeta_{2}+52H_{1,1}-32H_{0,1,1}+440H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)-32H_{0,1,1,1}\bigg]+C_{A}^{2}C_{F}\bigg[\frac{1}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+36H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1184}{3}\right)H_{1}-64H_{1}\zeta_{2}-64H_{1,0,0}+32H_{0,-1,-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1744}{9}\right)\zeta_{2}+\left(-\frac{722}{9}\right)H_{0,0}+\left(-\frac{1258}{27}\right)H_{0}+\frac{160}{3}\zeta_{3}+\frac{344}{3}H_{0,-1,0}-100H_{0}\zeta_{2}+7\zeta_{4}+60\zeta_{2}H_{0,0}+100H_{0,0,1}+72H_{0,-1,0,0}+96H_{0,-1,0,1}-8H_{0,0,-1,0}-60H_{0,0,0,1}+\frac{1}{z+1}+\frac{5333}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{416}{3}\right)H_{0,0,0}+\left(-\frac{1378}{27}\right)H_{0}+\left(-\frac{4}{3}\right)H_{0,0,1}+\frac{328}{3}H_{0,-1,0}+\frac{1504}{9}\zeta_{2}+\frac{332\zeta_{2}}{3}H_{0}-144H_{0}\zeta_{3}+112\zeta_{3}-121\zeta_{4}-68\zeta_{2}H_{0,0}-130H_{0,0}-56H_{0,-1,0,0}-160H_{0,-1,0,1}-40H_{0,0,-1,0}+40H_{0,0,0,0}+28H_{0,0,0,1}+\frac{1}{z+1}+\frac{923}{27}\right)+\frac{z^{2}}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+28H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{376}{3}\right)H_{0,-1,0}+\left(-\frac{2266}{27}\right)H_{0}+\frac{238}{9}H_{0,0}+\frac{224}{3}\zeta_{2}+188H_{0}\zeta_{2}+48\zeta_{3}-31\zeta_{4}+36\zeta_{2}H_{0,0}-188H_{0,0,1}+88H_{0,-1,0,0}+96H_{0,-1,0,1}+8H_{0,0,-1,0}-36H_{0,0,0,1}+\frac{1}{z+1}-\frac{5333}{27}\right)+\frac{1}{z+1}\left(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-120\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{560}{3}\right)\zeta_{3}+\left(-\frac{3442}{27}\right)H_{0}+\left(-\frac{80}{3}\right)H_{0,0,0}+\left(-\frac{8}{3}\right)H_{0,-1,0}+\frac{92}{3}H_{0,0,1}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}-112H_{0}\zeta_{3}-63\zeta_{4}-28\zeta_{2}H_{0,0}-162H_{0,0}-40H_{0,-1,0,0}-160H_{0,-1,0,1}-24H_{0,0,-1,0}+24H_{0,0,0,0}+4H_{0,0,0,1}-\frac{923}{27}\right)+\frac{z}{z+1}\left(\left(-\frac{448}{3}\right)H_{0,1}+\left(-\frac{160}{3}\right)H_{-1,0}-256H_{-1}\zeta_{2}-16\zeta_{2}H_{0,1}+96H_{-1,0,0}+320H_{-1,0,1}\right)$ (129)
-$\frac{z^{2}}{z+1}\left(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-136\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\right)+64zH_{-1,-1,0}+64H_{-1,-1,0}\bigg]+C_{F}N_{f}^{2}\bigg[\left(-\frac{416}{9}\right)H_{1}+\left(-\frac{152}{9}\right)H_{0,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,1}+\frac{152}{9}\zeta_{2}+\frac{64}{3}\zeta_{3}+\frac{1}{z-1}\left(\frac{8}{3}H_{0,0}+\frac{112}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}\right)+\frac{z}{z-1}\left(\left(-\frac{64}{3}\right)H_{0,0}+\frac{320}{9}H_{0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{128}{3}\right)H_{0}+\frac{152}{9}H_{0,0}+\left(-\frac{16\zeta_{2}}{3}\right)H_{0}\right)+40H_{1,0}+16H_{0,1,0}+z\left(\left(-\frac{296}{9}\right)H_{0,1}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{64}{3}\zeta_{3}+\frac{296}{9}\zeta_{2}+\frac{416}{9}H_{1}-40H_{1,0}+16H_{0,1,0}+\frac{1952}{27}\right)-\frac{1936}{27}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[C_{F}\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{2}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)C_{F}N_{f}C_{A}-\frac{16}{27}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)\bigg]+\delta(1-z)\bigg[\left(-32\zeta_{3}\zeta_{2}+18\zeta_{2}+68\zeta_{3}+144\zeta_{4}-240\zeta_{5}+\frac{29}{2}\right)C_{F}^{3}+\left(\left(-\frac{136}{3}\right)\zeta_{3}+\frac{20}{3}\zeta_{2}+\frac{116}{3}\zeta_{4}-23\right)N_{f}C_{F}^{2}+\left(\left(-\frac{16}{9}\right)\zeta_{3}+\frac{80}{27}\zeta_{2}-\frac{17}{9}\right)N_{f}^{2}C_{F}+C_{A}^{2}\left(\left(-\frac{1552}{9}\right)\zeta_{3}+\frac{4496}{27}\zeta_{2}-5\zeta_{4}+40\zeta_{5}-\frac{1657}{36}\right)C_{F}+C_{A}\left(\left(\left(-\frac{494}{3}\right)\zeta_{4}+\left(-\frac{410}{3}\right)\zeta_{2}+\frac{844}{3}\zeta_{3}+16\zeta_{2}\zeta_{3}+120\zeta_{5}+\frac{151}{4}\right)C_{F}^{2}+N_{f}\left(\left(-\frac{1336}{27}\right)\zeta_{2}+\frac{200}{9}\zeta_{3}+2\zeta_{4}+20\right)C_{F}\right)\bigg]+C_{A}C_{F}^{2}\bigg[\frac{1}{z-1}\left(\left(-\frac{2008}{3}\right)H_{1}+\frac{272}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-80H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{832}{3}\right)H_{1,0}+\frac{4016}{3}H_{1}+224H_{1}\zeta_{2}+128H_{1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1072}{3}\right)H_{0,-1,0}+\left(-\frac{2648}{9}\right)H_{0,1}+\left(-\frac{536}{3}\right)H_{0,0,1}+\left(-\frac{832}{9}\right)H_{0,0}+\frac{56}{3}\zeta_{3}+\frac{244}{9}H_{0}+\frac{2296}{9}\zeta_{2}+\frac{904\zeta_{2}}{3}H_{0}+268\zeta_{4}+416\zeta_{2}H_{0,-1}-192\zeta_{2}H_{0,0}+64\zeta_{2}H_{0,1}+140H_{0,0,0}+128H_{0,-1,-1,0}-208H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+192H_{0,0,0,1}+\frac{1}{z+1}+\frac{1516}{3}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3880}{9}\right)\zeta_{2}+\left(-\frac{832}{3}\right)\zeta_{3}+\left(-\frac{416}{3}\right)H_{0,-1,0}+\left(-\frac{350}{9}\right)H_{0}+\frac{1280}{9}H_{0,0}+\frac{652}{3}H_{0,0,0}+\frac{2296}{9}H_{0,1}-184H_{0}\zeta_{2}+224H_{0}\zeta_{3}+28\zeta_{4}-352\zeta_{2}H_{0,-1}+32\zeta_{2}H_{0,0}-128\zeta_{2}H_{0,1}+40H_{0,0,1}+256H_{0,-1,-1,0}+16H_{0,-1,0,0}+480H_{0,-1,0,1}-64H_{0,0,-1,0}+160H_{0,0,0,0}-32H_{0,0,0,1}+32H_{0,0,1,0}+\frac{1}{z+1}+\frac{1516}{3}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{2008}{3}\right)H_{1}+\frac{80}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-16H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{880}{3}\right)\zeta_{3}+\left(-\frac{1798}{9}\right)H_{0}+\left(-\frac{152}{9}\right)\zeta_{2}+\frac{776}{3}H_{0,0}+\frac{992}{3}H_{0,-1,0}+\frac{4792}{9}H_{0,1}+\frac{1672}{3}H_{0,0,1}+\left(-\frac{1688\zeta_{2}}{3}\right)H_{0}-12\zeta_{4}+448\zeta_{2}H_{0,-1}-96\zeta_{2}H_{0,0}+96\zeta_{2}H_{0,1}+36H_{0,0,0}+192H_{0,-1,-1,0}-176H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+96H_{0,0,0,1}+\frac{1}{z+1}-\frac{1516}{3}\right)$ (130)
-$\frac{1}{z+1}\left(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{80}{3}\right)H_{0,-1,0}+\left(-\frac{152}{9}\right)H_{0,1}+\frac{92}{9}H_{0}+\frac{436}{3}H_{0,0,0}+\frac{1736}{9}\zeta_{2}+\frac{1496}{3}\zeta_{3}+\frac{4504}{9}H_{0,0}+88H_{0}\zeta_{2}+256H_{0}\zeta_{3}+268\zeta_{4}-320\zeta_{2}H_{0,-1}-160\zeta_{2}H_{0,1}+8H_{0,0,1}+320H_{0,-1,-1,0}+48H_{0,-1,0,0}+480H_{0,-1,0,1}+32H_{0,0,0,0}+64H_{0,0,0,1}+32H_{0,0,1,0}-\frac{1516}{3}\right)+\frac{z}{z+1}\left(\frac{608}{3}H_{-1,0}+800H_{-1}\zeta_{2}-160H_{-1,0,0}-1024H_{-1,0,1}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\right)-224zH_{-1,-1,0}-224H_{-1,-1,0}\bigg]+N_{f}C_{A}C_{F}\bigg[\left(-\frac{1960}{9}\right)H_{1}+\left(-\frac{844}{9}\right)H_{0,1}+\left(-\frac{88}{3}\right)H_{0,1,1}+\frac{260}{3}H_{1,1}+\left(-\frac{356}{z-1}\right)H_{1,0,0}+\frac{z}{z-1}\left(\left(-\frac{544}{3}\right)H_{0}+\left(-\frac{164}{3}\right)H_{0,0,1}+\frac{472}{3}H_{0,0,0}+\frac{4700}{9}H_{0,0}+\left(-\frac{220\zeta_{2}}{3}\right)H_{0}-248\zeta_{2}+160H_{0}\zeta_{3}-168\zeta_{3}+426\zeta_{4}+8\zeta_{2}H_{0,0}+\frac{1}{z+1}-\frac{2950}{9}\right)+\frac{696z}{z-1}H_{1,0,0}+\left(-\frac{356z^{2}}{z-1}\right)H_{1,0,0}+\frac{z^{2}}{z-1}\left(\left(-\frac{440}{3}\right)H_{0,0,0}+\left(-\frac{1664}{27}\right)H_{0}+\frac{448}{9}H_{0,0}+\frac{332}{3}\zeta_{2}+\frac{656}{3}\zeta_{3}+\frac{548\zeta_{2}}{3}H_{0}-32H_{0}\zeta_{3}-406\zeta_{4}+104\zeta_{2}H_{0,0}-52H_{0,0,1}+\frac{1}{z+1}-\frac{8014}{27}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{1508}{3}\right)H_{0,0}+\left(-\frac{472}{3}\right)H_{0,0,0}+\frac{212}{3}H_{0,0,1}+\frac{488}{3}\zeta_{3}+\frac{6232}{27}H_{0}+\frac{2392}{9}\zeta_{2}+\frac{172\zeta_{2}}{3}H_{0}-160H_{0}\zeta_{3}-426\zeta_{4}-8\zeta_{2}H_{0,0}+\frac{1}{z+1}+\frac{2950}{9}\right)+\frac{1}{z+1}\left(\left(-\frac{128}{3}\right)H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{280}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{1156}{9}\right)\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{140}{3}H_{0,0,1}+\frac{1000}{9}H_{0}-188H_{0}\zeta_{2}+32H_{0}\zeta_{3}-192\zeta_{3}+406\zeta_{4}-104\zeta_{2}H_{0,0}+168H_{0,0,0}+\frac{8014}{27}\right)+\frac{z}{z+1}\left(\frac{112}{3}H_{-1,0}-16H_{-1}\zeta_{2}+368H_{-1,0,0}-64H_{-1,0,1}+240H_{0,-1,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{128}{3}\right)H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{472}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\right)-360H_{1}\zeta_{2}-16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}+268H_{1,0}-80H_{-1,-1,0}-64H_{0,1,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-32H_{0,-1,-1,0}+80H_{0,-1,0,0}+112H_{0,0,-1,0}-304H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+z\left(\left(-\frac{2536}{9}\right)H_{0,1}+\left(-\frac{260}{3}\right)H_{1,1}+\frac{152}{3}H_{0,1,1}+\frac{1960}{9}H_{1}+360H_{1}\zeta_{2}+16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}-268H_{1,0}-80H_{-1,-1,0}+216H_{0,1,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}+32H_{0,-1,-1,0}-80H_{0,-1,0,0}-48H_{0,0,-1,0}+64H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}\right)\bigg]\,.$ (131)
+$\Delta P_{qq}^{T,\text{s}(2)}=C_{F}N_{f}^{2}\bigg[\left(-\frac{416}{9}\right)H_{1}+\left(-\frac{152}{9}\right)H_{0,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,1}+\frac{152}{9}\zeta_{2}+\frac{64}{3}\zeta_{3}+\frac{1}{z-1}\left(\frac{8}{3}H_{0,0}+\frac{112}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}\right)+\frac{z}{z-1}\left(\left(-\frac{64}{3}\right)H_{0,0}+\frac{320}{9}H_{0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{128}{3}\right)H_{0}+\frac{152}{9}H_{0,0}+\left(-\frac{16\zeta_{2}}{3}\right)H_{0}\right)+40H_{1,0}+16H_{0,1,0}+z\left(\left(-\frac{296}{9}\right)H_{0,1}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{16}{3}H_{0,0,1}+\frac{16}{3}H_{0,1,1}+\frac{64}{3}\zeta_{3}+\frac{296}{9}\zeta_{2}+\frac{416}{9}H_{1}-40H_{1,0}+16H_{0,1,0}+\frac{1952}{27}\right)-\frac{1936}{27}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[C_{F}\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{2}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)C_{F}N_{f}C_{A}-\frac{16}{27}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)\bigg]+\delta(1-z)\bigg[\left(-32\zeta_{3}\zeta_{2}+18\zeta_{2}+68\zeta_{3}+144\zeta_{4}-240\zeta_{5}+\frac{29}{2}\right)C_{F}^{3}+\left(\left(-\frac{136}{3}\right)\zeta_{3}+\frac{20}{3}\zeta_{2}+\frac{116}{3}\zeta_{4}-23\right)N_{f}C_{F}^{2}+\left(\left(-\frac{16}{9}\right)\zeta_{3}+\frac{80}{27}\zeta_{2}-\frac{17}{9}\right)N_{f}^{2}C_{F}+C_{A}^{2}\left(\left(-\frac{1552}{9}\right)\zeta_{3}+\frac{4496}{27}\zeta_{2}-5\zeta_{4}+40\zeta_{5}-\frac{1657}{36}\right)C_{F}+C_{A}\left(\left(\left(-\frac{494}{3}\right)\zeta_{4}+\left(-\frac{410}{3}\right)\zeta_{2}+\frac{844}{3}\zeta_{3}+16\zeta_{2}\zeta_{3}+120\zeta_{5}+\frac{151}{4}\right)C_{F}^{2}+N_{f}\left(\left(-\frac{1336}{27}\right)\zeta_{2}+\frac{200}{9}\zeta_{3}+2\zeta_{4}+20\right)C_{F}\right)\bigg]$
+$C_{F}^{3}\left[\frac{1}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}-56H_{0,1,0}+96H_{1,0,0}-208H_{0,0,0,1}-64H_{0,0,1,1}-48H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z}{z-1}\left(-192\zeta_{2}H_{1}-800H_{1}+64H_{0,1,0}\right)+\frac{z}{z-1}\left(-144\zeta_{2}H_{0}-72H_{0}-12\zeta_{2}-192\zeta_{3}-324\zeta_{4}-352\zeta_{2}H_{0,-1}+144\zeta_{2}H_{0,0}+176H_{0,0}-96H_{0,0,0}-64H_{0,-1,-1,0}+128H_{0,-1,0,0}-96H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(-16\zeta_{2}H_{0}+128\zeta_{3}H_{0}+10H_{0}+204\zeta_{2}+48\zeta_{3}+252\zeta_{4}+160\zeta_{2}H_{0,-1}+208\zeta_{2}H_{0,0}+172H_{0,0}+528H_{0,0,0}-320H_{0,-1,-1,0}+192H_{0,-1,0,0}+288H_{0,0,-1,0}-592H_{0,0,0,0}-192H_{0,0,1,0}+\frac{1}{z+1}-302\right)+\frac{z^{2}}{z-1}\left(96\zeta_{2}H_{1}-192\zeta_{3}H_{1}+400H_{1}+88H_{0,1,0}+96H_{1,0,0}-240H_{0,0,0,1}-64H_{0,0,1,1}-144H_{0,1,0,0}-64H_{0,1,0,1}-64H_{0,1,1,0}-128H_{1,0,-1,0}-128H_{1,0,0,0}-128H_{1,0,0,1}-64H_{1,0,1,0}\right)+\frac{z^{3}}{z-1}\left(432\zeta_{2}H_{0}+114H_{0}+12\zeta_{2}+336\zeta_{3}-28\zeta_{4}-352\zeta_{2}H_{0,-1}+48\zeta_{2}H_{0,0}-228H_{0,0}+240H_{0,0,0}-64H_{0,-1,-1,0}-160H_{0,0,-1,0}-112H_{0,0,0,0}-128H_{0,0,1,0}+\frac{1}{z+1}+302\right)+\frac{1}{z+1}\left(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-12H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-64H_{0,-1,0}+64H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\right)+\frac{1}{(z-1)(z+1)}\bigg(16\zeta_{2}H_{0}-64\zeta_{3}H_{0}-64H_{0}-204\zeta_{2}-192\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}-224H_{0,0}-96H_{0,0,0}-320H_{0,-1,-1,0}+64H_{0,-1,0,0}+96H_{0,0,-1,0}-176H_{0,0,0,0}-128H_{0,0,1,0}+302\bigg)+\frac{z}{z+1}\left(-576H_{-1}\zeta_{2}-192H_{-1,0}-312H_{0,1}-64H_{-1,0,0}+768H_{-1,0,1}-320H_{0,-1,0}-96H_{0,0,1}\right)+\frac{z^{2}}{z+1}\bigg(-384H_{-1}\zeta_{2}+768H_{-1,-1}\zeta_{2}-384H_{-1,0}\zeta_{2}-96H_{0,1}\zeta_{2}-576H_{-1}\zeta_{3}-96H_{-1,0}-300H_{0,1}+112H_{-1,0,0}+480H_{-1,0,1}-160H_{0,-1,0}-256H_{0,0,1}-128H_{-1,-1,0,0}-768H_{-1,-1,0,1}+128H_{-1,0,-1,0}-224H_{-1,0,0,0}+256H_{-1,0,0,1}-64H_{-1,0,1,0}+320H_{0,-1,0,1}\bigg)+144H_{1,0}+192H_{-1,-1,0}+z\left(192H_{-1,-1,0}-144H_{1,0}\right)\right]+C_{F}^{2}N_{f}\bigg[\frac{668}{3}H_{1}+\frac{1}{z-1}\left(\left(-\frac{2104}{9}\right)H_{0,1}+\left(-\frac{464}{3}\right)H_{0,1,0}+\left(-\frac{580}{9}\right)H_{1,0}-160\zeta_{2}H_{0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{5048}{9}\right)H_{0,0}+\left(-\frac{496}{3}\right)H_{0,0,1}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{4772}{9}\zeta_{2}+\frac{7868}{9}H_{0}+\left(-\frac{112\zeta_{2}}{3}\right)H_{0}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-152H_{0,0,0}+\frac{1}{z+1}-\frac{1727}{3}\right)+\frac{z}{z-1}\left(\frac{220}{3}H_{0,1}+\frac{280}{3}H_{1,0}+144H_{0,1,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{580}{9}\right)H_{1,0}+\left(-\frac{32}{3}\right)H_{0,1,0}+\frac{1124}{9}H_{0,1}+160\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\bigg(\left(-\frac{80}{3}\right)\zeta_{3}+\frac{20}{9}H_{0}+\frac{56}{3}H_{0,0,0}+\frac{760}{9}H_{0,0}+\frac{1672}{9}\zeta_{2}-64H_{0}\zeta_{2}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+64H_{0,0,1}+\frac{1}{z+1}-539\bigg)+\frac{z^{3}}{z-1}\bigg(\left(-\frac{7664}{9}\right)H_{0}+\left(-\frac{5092}{9}\right)\zeta_{2}+\frac{112}{3}\zeta_{3}+\frac{368}{3}H_{0,0,1}+\frac{1624}{3}H_{0,0}+\frac{176\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+492\zeta_{4}-16\zeta_{2}H_{0,0}+120H_{0,0,0}+\frac{1}{z+1}+\frac{1727}{3}\bigg)$
+$\left.\frac{1}{z+1}\left(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\right)\right.+\frac{1}{(z-1)(z+1)}\bigg(\left(-\frac{1576}{9}\right)H_{0,0}+\left(-\frac{1352}{9}\right)\zeta_{2}+\left(-\frac{280}{3}\right)H_{0,0,0}+\left(-\frac{80}{3}\right)\zeta_{3}+\frac{184}{9}H_{0}+64H_{0}\zeta_{2}-160H_{0}\zeta_{3}-492\zeta_{4}+16\zeta_{2}H_{0,0}-64H_{0,0,1}+539\bigg)+\left(-\frac{128z}{3(z+1)}\right)H_{-1,0}+\frac{z^{2}}{z+1}\bigg(\left(-\frac{512}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,0,1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{-1}\bigg)+400H_{1}\zeta_{2}-52H_{1,1}+48H_{0,1,1}-440H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}+z\bigg(\left(-\frac{668}{3}\right)H_{1}-400H_{1}\zeta_{2}+52H_{1,1}-32H_{0,1,1}+440H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-64H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-176H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\bigg)-32H_{0,1,1,1}\bigg]+C_{A}^{2}C_{F}\left[\frac{1}{z-1}\left(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+36H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\right)+\frac{z}{z-1}\left(\left(-\frac{1184}{3}\right)H_{1}-64H_{1}\zeta_{2}-64H_{1,0,0}+32H_{0,-1,-1,0}\right)+\frac{z}{z-1}\bigg(\left(-\frac{1744}{9}\right)\zeta_{2}+\left(-\frac{722}{9}\right)H_{0,0}+\left(-\frac{1258}{27}\right)H_{0}+\frac{160}{3}\zeta_{3}+\frac{344}{3}H_{0,-1,0}-100H_{0}\zeta_{2}+7\zeta_{4}+60\zeta_{2}H_{0,0}+100H_{0,0,1}+72H_{0,-1,0,0}+96H_{0,-1,0,1}-8H_{0,0,-1,0}-60H_{0,0,0,1}+\frac{1}{z+1}+\frac{5333}{27}\bigg)+\frac{z^{2}}{z-1}\bigg(\left(-\frac{416}{3}\right)H_{0,0,0}+\left(-\frac{1378}{27}\right)H_{0}+\left(-\frac{4}{3}\right)H_{0,0,1}+\frac{328}{3}H_{0,-1,0}+\frac{1504}{9}\zeta_{2}+\frac{332\zeta_{2}}{3}H_{0}-144H_{0}\zeta_{3}+112\zeta_{3}-121\zeta_{4}-68\zeta_{2}H_{0,0}-130H_{0,0}-56H_{0,-1,0,0}-160H_{0,-1,0,1}-40H_{0,0,-1,0}+40H_{0,0,0,0}+28H_{0,0,0,1}+\frac{1}{z+1}+\frac{923}{27}\bigg)+\frac{z^{2}}{z-1}\bigg(\frac{592}{3}H_{1}+32H_{1}\zeta_{2}-144H_{1}\zeta_{3}+16\zeta_{2}H_{1,0}+76H_{1,0,0}-80H_{0,-1,-1,0}+28H_{0,1,0,0}-96H_{1,0,-1,0}+48H_{1,0,0,0}-64H_{1,0,0,1}+64H_{1,1,0,0}\bigg)+\frac{z^{3}}{z-1}\bigg(\left(-\frac{376}{3}\right)H_{0,-1,0}+\left(-\frac{2266}{27}\right)H_{0}+\frac{238}{9}H_{0,0}+\frac{224}{3}\zeta_{2}+188H_{0}\zeta_{2}+48\zeta_{3}-31\zeta_{4}+36\zeta_{2}H_{0,0}-188H_{0,0,1}+88H_{0,-1,0,0}+96H_{0,-1,0,1}+8H_{0,0,-1,0}-36H_{0,0,0,1}+\frac{1}{z+1}-\frac{5333}{27}\bigg)+\frac{1}{(z-1)(z+1)}\bigg(\left(-\frac{560}{3}\right)\zeta_{3}+\left(-\frac{3442}{27}\right)H_{0}+\left(-\frac{80}{3}\right)H_{0,0,0}+\left(-\frac{8}{3}\right)H_{0,-1,0}+\frac{92}{3}H_{0,0,1}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}-112H_{0}\zeta_{3}-63\zeta_{4}-28\zeta_{2}H_{0,0}-162H_{0,0}-40H_{0,-1,0,0}-160H_{0,-1,0,1}-24H_{0,0,-1,0}+24H_{0,0,0,0}+4H_{0,0,0,1}-\frac{923}{27}\bigg)+\frac{z}{z+1}\bigg(\left(-\frac{448}{3}\right)H_{0,1}+\left(-\frac{160}{3}\right)H_{-1,0}-256H_{-1}\zeta_{2}-16\zeta_{2}H_{0,1}+96H_{-1,0,0}+320H_{-1,0,1}\bigg)+\frac{z^{2}}{z+1}\bigg(-\frac{1312}{9}H_{-1,0}-\frac{224}{3}H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}-\frac{560\zeta_{2}}{3}H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-136\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\bigg)+\frac{1}{z+1}\bigg(\left(-\frac{1312}{9}\right)H_{-1,0}+\left(-\frac{224}{3}\right)H_{0,1}+\frac{56}{3}H_{-1,0,0}+\frac{656}{3}H_{-1,0,1}+\left(-\frac{560\zeta_{2}}{3}\right)H_{-1}-192H_{-1}\zeta_{3}+256\zeta_{2}H_{-1,-1}-176\zeta_{2}H_{-1,0}-120\zeta_{2}H_{0,-1}-40\zeta_{2}H_{0,1}-128H_{-1,-1,0,0}-256H_{-1,-1,0,1}+48H_{-1,0,0,0}+128H_{-1,0,0,1}\bigg)+64zH_{-1,-1,0}+64H_{-1,-1,0}\right]$
+$C_{A}C_{F}^{2}\left[\frac{1}{z-1}\bigg(\left(-\frac{2008}{3}\right)H_{1}+\frac{272}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-80H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\bigg)+\frac{z}{z-1}\bigg(\left(-\frac{832}{3}\right)H_{1,0}+\frac{4016}{3}H_{1}+224H_{1}\zeta_{2}+128H_{1,0,0}\bigg)+\frac{z}{z-1}\bigg(\left(-\frac{1072}{3}\right)H_{0,-1,0}+\left(-\frac{2648}{9}\right)H_{0,1}+\left(-\frac{536}{3}\right)H_{0,0,1}+\left(-\frac{832}{9}\right)H_{0,0}+\frac{56}{3}\zeta_{3}+\frac{244}{9}H_{0}+\frac{2296}{9}\zeta_{2}+\frac{904\zeta_{2}}{3}H_{0}+268\zeta_{4}+416\zeta_{2}H_{0,-1}-192\zeta_{2}H_{0,0}+64\zeta_{2}H_{0,1}+140H_{0,0,0}+128H_{0,-1,-1,0}-208H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+192H_{0,0,0,1}+\frac{1}{z+1}+\frac{1516}{3}\bigg)+\frac{z^{2}}{z-1}\bigg(\left(-\frac{3880}{9}\right)\zeta_{2}+\left(-\frac{832}{3}\right)\zeta_{3}+\left(-\frac{416}{3}\right)H_{0,-1,0}+\left(-\frac{350}{9}\right)H_{0}+\frac{1280}{9}H_{0,0}+\frac{652}{3}H_{0,0,0}+\frac{2296}{9}H_{0,1}-184H_{0}\zeta_{2}+224H_{0}\zeta_{3}+28\zeta_{4}-352\zeta_{2}H_{0,-1}+32\zeta_{2}H_{0,0}-128\zeta_{2}H_{0,1}+40H_{0,0,1}+256H_{0,-1,-1,0}+16H_{0,-1,0,0}+480H_{0,-1,0,1}-64H_{0,0,-1,0}+160H_{0,0,0,0}-32H_{0,0,0,1}+32H_{0,0,1,0}+\frac{1}{z+1}+\frac{1516}{3}\bigg)+\frac{z^{2}}{z-1}\bigg(\left(-\frac{2008}{3}\right)H_{1}+\frac{80}{3}H_{0,1,0}+\frac{2320}{9}H_{1,0}-112H_{1}\zeta_{2}+384H_{1}\zeta_{3}-32\zeta_{2}H_{1,0}-112H_{1,0,0}-16H_{0,1,0,0}+256H_{1,0,-1,0}-80H_{1,0,0,0}+128H_{1,0,0,1}-128H_{1,1,0,0}\bigg)+\frac{z^{3}}{z-1}\bigg(\left(-\frac{880}{3}\right)\zeta_{3}+\left(-\frac{1798}{9}\right)H_{0}+\left(-\frac{152}{9}\right)\zeta_{2}+\frac{776}{3}H_{0,0}+\frac{992}{3}H_{0,-1,0}+\frac{4792}{9}H_{0,1}+\frac{1672}{3}H_{0,0,1}+\left(-\frac{1688\zeta_{2}}{3}\right)H_{0}-12\zeta_{4}+448\zeta_{2}H_{0,-1}-96\zeta_{2}H_{0,0}+96\zeta_{2}H_{0,1}+36H_{0,0,0}+192H_{0,-1,-1,0}-176H_{0,-1,0,0}-352H_{0,-1,0,1}+64H_{0,0,-1,0}+96H_{0,0,0,1}+\frac{1}{z+1}-\frac{1516}{3}\bigg)+\frac{1}{z+1}\bigg(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\bigg)+\frac{1}{(z-1)(z+1)}\bigg(\left(-\frac{80}{3}\right)H_{0,-1,0}+\left(-\frac{152}{9}\right)H_{0,1}+\frac{92}{9}H_{0}+\frac{436}{3}H_{0,0,0}+\frac{1736}{9}\zeta_{2}+\frac{1496}{3}\zeta_{3}+\frac{4504}{9}H_{0,0}+88H_{0}\zeta_{2}+256H_{0}\zeta_{3}+268\zeta_{4}-320\zeta_{2}H_{0,-1}-160\zeta_{2}H_{0,1}+8H_{0,0,1}+320H_{0,-1,-1,0}+48H_{0,-1,0,0}+480H_{0,-1,0,1}+32H_{0,0,0,0}+64H_{0,0,0,1}+32H_{0,0,1,0}-\frac{1516}{3}\bigg)+\frac{z}{z+1}\left(\frac{608}{3}H_{-1,0}+800H_{-1}\zeta_{2}-160H_{-1,0,0}-1024H_{-1,0,1}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{2032}{3}\right)H_{-1,0,1}+\left(-\frac{280}{3}\right)H_{-1,0,0}+\frac{3056}{9}H_{-1,0}+\frac{1696\zeta_{2}}{3}H_{-1}+672H_{-1}\zeta_{3}-896\zeta_{2}H_{-1,-1}+544\zeta_{2}H_{-1,0}+320H_{-1,-1,0,0}+896H_{-1,-1,0,1}-64H_{-1,0,-1,0}+16H_{-1,0,0,0}-384H_{-1,0,0,1}+32H_{-1,0,1,0}\right)-224zH_{-1,-1,0}-224H_{-1,-1,0}\right]+N_{f}C_{A}C_{F}\bigg[-\frac{1960}{9}H_{1}-\frac{844}{9}H_{0,1}-\frac{88}{3}H_{0,1,1}+\frac{260}{3}H_{1,1}-\frac{356}{z-1}H_{1,0,0}+\frac{z}{z-1}\bigg(\left(-\frac{544}{3}\right)H_{0}+\left(-\frac{164}{3}\right)H_{0,0,1}+\frac{472}{3}H_{0,0,0}+\frac{4700}{9}H_{0,0}+\left(-\frac{220\zeta_{2}}{3}\right)H_{0}-248\zeta_{2}+160H_{0}\zeta_{3}-168\zeta_{3}+426\zeta_{4}+8\zeta_{2}H_{0,0}+\frac{1}{z+1}-\frac{2950}{9}\bigg)+\frac{696z}{z-1}H_{1,0,0}+\left(-\frac{356z^{2}}{z-1}\right)H_{1,0,0}+\frac{z^{2}}{z-1}\bigg(\left(-\frac{440}{3}\right)H_{0,0,0}+\left(-\frac{1664}{27}\right)H_{0}+\frac{448}{9}H_{0,0}+\frac{332}{3}\zeta_{2}+\frac{656}{3}\zeta_{3}+\frac{548\zeta_{2}}{3}H_{0}-32H_{0}\zeta_{3}-406\zeta_{4}+104\zeta_{2}H_{0,0}-52H_{0,0,1}+\frac{1}{z+1}-\frac{8014}{27}\bigg)$
+$\frac{z^{3}}{z-1}\left(-\frac{1508}{3}H_{0,0}-\frac{472}{3}H_{0,0,0}+\frac{212}{3}H_{0,0,1}+\frac{488}{3}\zeta_{3}+\frac{6232}{27}H_{0}+\frac{2392}{9}\zeta_{2}+\frac{172\zeta_{2}}{3}H_{0}-160H_{0}\zeta_{3}-426\zeta_{4}-8\zeta_{2}H_{0,0}+\frac{1}{z+1}+\frac{2950}{9}\right)+\frac{1}{z+1}\bigg(-\frac{128}{3}H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{280}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\bigg)+\frac{1}{(z-1)(z+1)}\bigg(-\frac{1156}{9}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{140}{3}H_{0,0,1}+\frac{1000}{9}H_{0}-188H_{0}\zeta_{2}+32H_{0}\zeta_{3}-192\zeta_{3}+406\zeta_{4}-104\zeta_{2}H_{0,0}+168H_{0,0,0}+\frac{8014}{27}\bigg)+\frac{z}{z+1}\left(\frac{112}{3}H_{-1,0}-16H_{-1}\zeta_{2}+368H_{-1,0,0}-64H_{-1,0,1}+240H_{0,-1,0}\right)+\frac{z^{2}}{z+1}\bigg(-\frac{128}{3}H_{-1,0,1}+\frac{328}{9}H_{-1,0}+\frac{472}{3}H_{0,-1,0}+\frac{568}{3}H_{-1,0,0}+\frac{8\zeta_{2}}{3}H_{-1}\bigg)-360H_{1}\zeta_{2}-16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}+268H_{1,0}-80H_{-1,-1,0}-64H_{0,1,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}-32H_{0,-1,-1,0}+80H_{0,-1,0,0}+112H_{0,0,-1,0}-304H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+z\bigg(-\frac{2536}{9}H_{0,1}-\frac{260}{3}H_{1,1}+\frac{152}{3}H_{0,1,1}+\frac{1960}{9}H_{1}+360H_{1}\zeta_{2}+16\zeta_{2}H_{0,-1}-144\zeta_{2}H_{0,1}-268H_{1,0}-80H_{-1,-1,0}+216H_{0,1,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}+32H_{0,-1,-1,0}-80H_{0,-1,0,0}-48H_{0,0,-1,0}+64H_{0,0,0,0}+8H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+136H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}\bigg)\bigg]\,.$ (128)
 $\Delta P_{gq}^{T,\text{s}(0)}=(4-2z)C_{F}\,$, (129)
 $\Delta P_{gq}^{T,\text{s}(1)}=C_{A}C_{F}\left(z\left(8H_{-1,0}-8H_{0,0}-8H_{0,1}+24H_{1,0}+8H_{1,1}+44H_{0}+16H_{1}+8\zeta_{2}-16\right)+16H_{-1,0}-48H_{0,0}+16H_{0,1}-48H_{1,0}-16H_{1,1}-16H_{0}-16H_{1}+20\right)+C_{F}^{2}\left(z\left(4H_{0,0}+16H_{0,1}-16H_{1,0}-8H_{1,1}-18H_{0}-16H_{1}-28\right)-8H_{0,0}-32H_{0,1}+32H_{1,0}+16H_{1,1}+32H_{0}+16H_{1}+38\right)\,$, (130)
-$\Delta P_{gq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\left(-\frac{48}{z}\right)H_{-1,0}+447H_{0}+626H_{1}+32H_{-1}\zeta_{2}+24H_{0}\zeta_{2}-112H_{1}\zeta_{2}-424\zeta_{2}-400H_{0}\zeta_{3}+64H_{1}\zeta_{3}-224\zeta_{3}+84\zeta_{4}+48H_{-1,0}+192\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+174H_{0,0}+400\zeta_{2}H_{0,1}+40H_{0,1}+128\zeta_{2}H_{1,0}+304H_{1,0}+96\zeta_{2}H_{1,1}+240H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-168H_{0,0,0}-344H_{0,0,1}+40H_{0,1,0}-40H_{0,1,1}+80H_{1,0,0}-16H_{1,0,1}+80H_{1,1,0}+96H_{1,1,1}-128H_{-1,-1,0,0}-64H_{-1,0,-1,0}+32H_{-1,0,0,0}+384H_{0,-1,-1,0}-128H_{0,-1,0,0}-192H_{0,0,-1,0}+192H_{0,0,0,0}+128H_{0,0,0,1}+176H_{0,0,1,0}+240H_{0,0,1,1}+80H_{0,1,0,0}-144H_{0,1,0,1}-80H_{0,1,1,0}-208H_{0,1,1,1}-64H_{1,0,-1,0}-192H_{1,0,0,0}-256H_{1,0,0,1}+32H_{1,0,1,0}-128H_{1,0,1,1}-192H_{1,1,0,0}-96H_{1,1,0,1}+32H_{1,1,1,0}+z\left(-60\zeta_{2}H_{0}-40\zeta_{3}H_{0}-221H_{0}-592H_{1}+32H_{-1}\zeta_{2}+76H_{1}\zeta_{2}+44\zeta_{2}-32H_{1}\zeta_{3}+76\zeta_{3}-194\zeta_{4}+64H_{-1,0}+64\zeta_{2}H_{0,-1}-32\zeta_{2}H_{0,0}-73H_{0,0}-232\zeta_{2}H_{0,1}+244H_{0,1}-64\zeta_{2}H_{1,0}-248H_{1,0}-48\zeta_{2}H_{1,1}-184H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-144H_{0,-1,0}+108H_{0,0,0}+140H_{0,0,1}+56H_{0,1,0}+112H_{0,1,1}+28H_{1,0,0}+52H_{1,0,1}-44H_{1,1,0}-84H_{1,1,1}-64H_{-1,-1,0,0}-32H_{-1,0,-1,0}+16H_{-1,0,0,0}+128H_{0,-1,-1,0}-32H_{0,-1,0,0}-32H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,0,1}-88H_{0,0,1,0}-120H_{0,0,1,1}-8H_{0,1,0,0}+72H_{0,1,0,1}+40H_{0,1,1,0}+104H_{0,1,1,1}+32H_{1,0,-1,0}+96H_{1,0,0,0}+128H_{1,0,0,1}-16H_{1,0,1,0}+64H_{1,0,1,1}+96H_{1,1,0,0}+48H_{1,1,0,1}-16H_{1,1,1,0}-32H_{1,1,1,1}-\frac{3133}{4}\right)+64H_{1,1,1,1}+\frac{1569}{2}\bigg]+N_{f}C_{F}^{2}\bigg[\left(-\frac{80}{3}\right)H_{1,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{112}{9}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,1,1}+\left(-\frac{16}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{1,0,1}+\left(-\frac{16}{3}\right)H_{1,1,0}+\frac{76}{27}H_{1}+\frac{16}{3}H_{1,1,1}+\frac{64}{3}H_{0,0,1}+\frac{224}{9}H_{0,1}+\frac{380}{3}H_{0,0}+\frac{4334}{27}H_{0}-16H_{1}\zeta_{2}-64H_{0,0,0,0}+z\left(\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{2038}{27}\right)H_{0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{64}{9}\right)H_{0,1}+\left(-\frac{8}{3}\right)H_{1,1,1}+\frac{8}{3}H_{1,0,1}+\frac{8}{3}H_{1,1,0}+\frac{32}{9}H_{1,1}+\frac{112}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,0,0}+8H_{1}\zeta_{2}+154H_{0,0}+32H_{0,0,0,0}-\frac{1567}{18}\right)+\frac{767}{9}\bigg]+C_{A}^{2}C_{F}\bigg[\left(-\frac{7520}{9}\right)\zeta_{2}+\left(-\frac{7376}{9}\right)H_{1,0}+\left(-\frac{1384}{3}\right)H_{-1,0,0}+\left(-\frac{904}{3}\right)H_{0,-1,0}+\left(-\frac{2248}{9}\right)H_{1,1}+\left(-\frac{2216}{9}\right)H_{-1,0}+\left(-\frac{268}{3}\right)H_{1,0,0}+\frac{8}{3}H_{0,1,1}+\frac{88}{3}H_{1,1,1}+\frac{112}{3}H_{1,1,0}+\frac{932}{3}H_{0,0,1}+\frac{5584}{9}H_{0,0}+\frac{2216}{3}H_{0,0,0}+\frac{7624}{9}H_{0,1}+\frac{28256}{27}H_{0}+\frac{32362}{27}H_{1}+\left(-\frac{24}{z}\right)H_{-1,0}+\left(-\frac{1468\zeta_{2}}{3}\right)H_{0}+104H_{-1}\zeta_{2}+264H_{1}\zeta_{2}-16H_{-1}\zeta_{3}+480H_{0}\zeta_{3}+656H_{1}\zeta_{3}-712\zeta_{3}+1570\zeta_{4}-32\zeta_{2}H_{-1,-1}$ (134)
-$160\zeta_{2}H_{-1,0}-16\zeta_{2}H_{0,-1}-312\zeta_{2}H_{0,0}+496\zeta_{2}H_{0,1}-128\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}+288H_{0,1,0}-96H_{1,0,1}-192H_{-1,-1,-1,0}+192H_{-1,-1,0,0}-64H_{-1,-1,0,1}+352H_{-1,0,-1,0}-336H_{-1,0,0,0}+64H_{-1,0,0,1}-192H_{-1,0,1,0}-64H_{-1,0,1,1}+96H_{0,-1,-1,0}-272H_{0,-1,0,0}+64H_{0,-1,0,1}-496H_{0,0,-1,0}+1264H_{0,0,0,0}-184H_{0,0,0,1}+960H_{0,0,1,0}+320H_{0,0,1,1}-152H_{0,1,0,0}-352H_{0,1,0,1}-224H_{0,1,1,0}-160H_{0,1,1,1}-96H_{1,0,-1,0}+560H_{1,0,0,0}+160H_{1,0,0,1}+640H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}+128H_{1,1,0,1}+192H_{1,1,1,0}+z\left(\left(-\frac{31412}{27}\right)H_{1}+\left(-\frac{20644}{27}\right)H_{0}+\left(-\frac{1472}{3}\right)H_{-1,0,0}+\left(-\frac{992}{3}\right)H_{0,-1,0}+\left(-\frac{2752}{9}\right)H_{-1,0}+\left(-\frac{580}{3}\right)H_{0,0,1}+\left(-\frac{484}{3}\right)H_{0,1,1}+\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{200}{3}\right)H_{1,1,0}+\left(-\frac{172}{3}\right)H_{1,0,0}+\left(-\frac{44}{3}\right)H_{1,1,1}+\frac{1712}{9}H_{1,1}+\frac{5380}{9}H_{0,1}+\frac{5944}{9}H_{1,0}+\frac{10172}{9}H_{0,0}+104H_{-1}\zeta_{2}-156H_{0}\zeta_{2}-352H_{1}\zeta_{2}-796\zeta_{2}-8H_{-1}\zeta_{3}+176H_{0}\zeta_{3}-328H_{1}\zeta_{3}-464\zeta_{3}+143\zeta_{4}-16\zeta_{2}H_{-1,-1}+80\zeta_{2}H_{-1,0}-72\zeta_{2}H_{0,-1}+76\zeta_{2}H_{0,0}+24\zeta_{2}H_{0,1}+64\zeta_{2}H_{1,0}-48\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}-848H_{0,1,0}+96H_{1,0,1}-96H_{-1,-1,-1,0}+96H_{-1,-1,0,0}-32H_{-1,-1,0,1}+176H_{-1,0,-1,0}-168H_{-1,0,0,0}+32H_{-1,0,0,1}-96H_{-1,0,1,0}-32H_{-1,0,1,1}-80H_{0,-1,-1,0}+120H_{0,-1,0,0}+32H_{0,-1,0,1}+56H_{0,0,-1,0}-64H_{0,0,0,0}-76H_{0,0,0,1}+192H_{0,0,1,0}+64H_{0,0,1,1}+116H_{0,1,0,0}+32H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+48H_{1,0,-1,0}-280H_{1,0,0,0}-80H_{1,0,0,1}-320H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}-64H_{1,1,0,1}-96H_{1,1,1,0}-32H_{1,1,1,1}-\frac{13346}{9}\right)+64H_{1,1,1,1}+\frac{13840}{9}\bigg]+C_{A}C_{F}^{2}\bigg[\left(-\frac{52072}{27}\right)H_{1}+\left(-\frac{10868}{9}\right)H_{0,1}+\left(-\frac{1466}{3}\right)H_{0,0}+\left(-\frac{1136}{3}\right)H_{0,0,0}+\left(-\frac{3719}{27}\right)H_{0}+\left(-\frac{376}{3}\right)H_{1,1,1}+\left(-\frac{368}{3}\right)H_{1,1,0}+\left(-\frac{184}{3}\right)H_{0,0,1}+\left(-\frac{344}{9}\right)H_{1,1}+\left(-\frac{88}{3}\right)H_{1,0}+\frac{56}{3}H_{0,1,1}+\frac{272}{3}H_{1,0,0}+\frac{352}{3}H_{1,0,1}+\frac{1}{z}\left(-72H_{-1}\zeta_{2}-24H_{1}\zeta_{2}+72H_{-1,0}-48H_{-1,-1,0}+48H_{-1,0,0}+48H_{-1,0,1}\right)-256H_{-1}\zeta_{2}+296H_{0}\zeta_{2}-240H_{1}\zeta_{2}+716\zeta_{2}-176H_{-1}\zeta_{3}-752H_{0}\zeta_{3}-816H_{1}\zeta_{3}+728\zeta_{3}-2144\zeta_{4}+288\zeta_{2}H_{-1,-1}-192\zeta_{2}H_{-1,0}+136H_{-1,0}-32\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-912\zeta_{2}H_{0,1}+32\zeta_{2}H_{1,0}-192\zeta_{2}H_{1,1}-256H_{-1,-1,0}+144H_{-1,0,0}+128H_{-1,0,1}+96H_{0,-1,0}-472H_{0,1,0}+192H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-192H_{-1,-1,0,1}-160H_{-1,0,-1,0}-16H_{-1,0,0,0}-64H_{-1,0,0,1}+128H_{-1,0,1,0}+64H_{-1,0,1,1}-192H_{0,-1,-1,0}-64H_{0,-1,0,0}-64H_{0,-1,0,1}+96H_{0,0,-1,0}+128H_{0,0,0,0}+752H_{0,0,0,1}-688H_{0,0,1,0}-432H_{0,0,1,1}+416H_{0,1,0,0}+624H_{0,1,0,1}+432H_{0,1,1,0}+368H_{0,1,1,1}+96H_{1,0,-1,0}-48H_{1,0,0,0}+224H_{1,0,0,1}-544H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}-32H_{1,1,0,1}-224H_{1,1,1,0}-128H_{1,1,1,1}+z\left(\left(-\frac{6224}{9}\right)H_{0,1}+\left(-\frac{688}{3}\right)H_{0,0,1}+\left(-\frac{452}{3}\right)H_{1,0,1}+\left(-\frac{448}{3}\right)H_{1,0,0}+\frac{56}{3}H_{1,0}+\frac{68}{3}H_{0,1,1}+\frac{376}{9}H_{1,1}+\frac{296}{3}H_{1,1,1}+\frac{340}{3}H_{1,1,0}+\frac{364}{3}H_{0,0,0}+\frac{50168}{27}H_{1}+\frac{57247}{27}H_{0}-184H_{-1}\zeta_{2}+72H_{0}\zeta_{2}+380H_{1}\zeta_{2}+1176\zeta_{2}-88H_{-1}\zeta_{3}-136H_{0}\zeta_{3}+408H_{1}\zeta_{3}+368\zeta_{3}-152\zeta_{4}+144\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}+64H_{-1,0}-64\zeta_{2}H_{0,-1}-8\zeta_{2}H_{0,0}-889H_{0,0}+136\zeta_{2}H_{0,1}-16\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}-208H_{-1,-1,0}+24H_{-1,0,0}+80H_{-1,0,1}+264H_{0,-1,0}+528H_{0,1,0}+96H_{-1,-1,-1,0}+32H_{-1,-1,0,0}-96H_{-1,-1,0,1}-80H_{-1,0,-1,0}-8H_{-1,0,0,0}-32H_{-1,0,0,1}+64H_{-1,0,1,0}+32H_{-1,0,1,1}-64H_{0,-1,-1,0}-16H_{0,-1,0,0}+32H_{0,-1,0,1}-16H_{0,0,-1,0}+32H_{0,0,0,0}+168H_{0,0,0,1}-104H_{0,0,1,0}-8H_{0,0,1,1}-160H_{0,1,0,0}-168H_{0,1,0,1}-168H_{0,1,1,0}-136H_{0,1,1,1}-48H_{1,0,-1,0}+24H_{1,0,0,0}-112H_{1,0,0,1}+272H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}+16H_{1,1,0,1}+112H_{1,1,1,0}+64H_{1,1,1,1}-\frac{2737}{36}\right)+\frac{1577}{18}\bigg]$ (135)
-$N_{f}C_{A}C_{F}\bigg[\left(-\frac{728}{27}\right)H_{0}+\left(-\frac{112}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{88}{9}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,1,1}+\frac{8}{9}\zeta_{2}+\frac{32}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{32}{3}H_{1,1,0}+\frac{112}{9}H_{1,1}+\frac{368}{9}H_{1,0}+\frac{160}{3}H_{1,0,0}+\frac{584}{9}H_{0,0}+\frac{256}{3}H_{0,0,0}+\left(-\frac{32\zeta_{2}}{3}\right)H_{0}+32H_{1}\zeta_{2}-80\zeta_{3}+z\left(\left(-\frac{356}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{208}{9}\right)H_{1,0}+\left(-\frac{518}{27}\right)H_{0}+\left(-\frac{184}{27}\right)H_{1}+\left(-\frac{16}{3}\right)H_{-1,0,0}+\left(-\frac{16}{3}\right)H_{0,-1,0}+\left(-\frac{16}{3}\right)H_{1,1,0}+\left(-\frac{32}{9}\right)H_{-1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\left(-\frac{8}{3}\right)H_{0,1,1}+\frac{8}{3}H_{1,1,1}+\frac{32}{9}H_{0,1}+\frac{16}{3}H_{0,0,0}+\frac{16}{3}H_{0,0,1}-16H_{1}\zeta_{2}-16\zeta_{2}+32\zeta_{3}+\frac{416}{9}\right)-\frac{526}{9}\bigg]\,$, (136)
+$\Delta P_{gq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\left(-\frac{48}{z}\right)H_{-1,0}+447H_{0}+626H_{1}+32H_{-1}\zeta_{2}+24H_{0}\zeta_{2}-112H_{1}\zeta_{2}-424\zeta_{2}-400H_{0}\zeta_{3}+64H_{1}\zeta_{3}-224\zeta_{3}+84\zeta_{4}+48H_{-1,0}+192\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+174H_{0,0}+400\zeta_{2}H_{0,1}+40H_{0,1}+128\zeta_{2}H_{1,0}+304H_{1,0}+96\zeta_{2}H_{1,1}+240H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-168H_{0,0,0}-344H_{0,0,1}+40H_{0,1,0}-40H_{0,1,1}+80H_{1,0,0}-16H_{1,0,1}+80H_{1,1,0}+96H_{1,1,1}-128H_{-1,-1,0,0}-64H_{-1,0,-1,0}+32H_{-1,0,0,0}+384H_{0,-1,-1,0}-128H_{0,-1,0,0}-192H_{0,0,-1,0}+192H_{0,0,0,0}+128H_{0,0,0,1}+176H_{0,0,1,0}+240H_{0,0,1,1}+80H_{0,1,0,0}-144H_{0,1,0,1}-80H_{0,1,1,0}-208H_{0,1,1,1}-64H_{1,0,-1,0}-192H_{1,0,0,0}-256H_{1,0,0,1}+32H_{1,0,1,0}-128H_{1,0,1,1}-192H_{1,1,0,0}-96H_{1,1,0,1}+32H_{1,1,1,0}+z\left(-60\zeta_{2}H_{0}-40\zeta_{3}H_{0}-221H_{0}-592H_{1}+32H_{-1}\zeta_{2}+76H_{1}\zeta_{2}+44\zeta_{2}-32H_{1}\zeta_{3}+76\zeta_{3}-194\zeta_{4}+64H_{-1,0}+64\zeta_{2}H_{0,-1}-32\zeta_{2}H_{0,0}-73H_{0,0}-232\zeta_{2}H_{0,1}+244H_{0,1}-64\zeta_{2}H_{1,0}-248H_{1,0}-48\zeta_{2}H_{1,1}-184H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-144H_{0,-1,0}+108H_{0,0,0}+140H_{0,0,1}+56H_{0,1,0}+112H_{0,1,1}+28H_{1,0,0}+52H_{1,0,1}-44H_{1,1,0}-84H_{1,1,1}-64H_{-1,-1,0,0}-32H_{-1,0,-1,0}+16H_{-1,0,0,0}+128H_{0,-1,-1,0}-32H_{0,-1,0,0}-32H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,0,1}-88H_{0,0,1,0}-120H_{0,0,1,1}-8H_{0,1,0,0}+72H_{0,1,0,1}+40H_{0,1,1,0}+104H_{0,1,1,1}+32H_{1,0,-1,0}+96H_{1,0,0,0}+128H_{1,0,0,1}-16H_{1,0,1,0}+64H_{1,0,1,1}+96H_{1,1,0,0}+48H_{1,1,0,1}-16H_{1,1,1,0}-32H_{1,1,1,1}-\frac{3133}{4}\right)+64H_{1,1,1,1}+\frac{1569}{2}\bigg]+N_{f}C_{F}^{2}\bigg[\left(-\frac{80}{3}\right)H_{1,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{112}{9}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,1,1}+\left(-\frac{16}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{1,0,1}+\left(-\frac{16}{3}\right)H_{1,1,0}+\frac{76}{27}H_{1}+\frac{16}{3}H_{1,1,1}+\frac{64}{3}H_{0,0,1}+\frac{224}{9}H_{0,1}+\frac{380}{3}H_{0,0}+\frac{4334}{27}H_{0}-16H_{1}\zeta_{2}-64H_{0,0,0,0}+z\left(\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{2038}{27}\right)H_{0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{64}{9}\right)H_{0,1}+\left(-\frac{8}{3}\right)H_{1,1,1}+\frac{8}{3}H_{1,0,1}+\frac{8}{3}H_{1,1,0}+\frac{32}{9}H_{1,1}+\frac{112}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,0,0}+8H_{1}\zeta_{2}+154H_{0,0}+32H_{0,0,0,0}-\frac{1567}{18}\right)+\frac{767}{9}\bigg]+C_{A}^{2}C_{F}\bigg[\left(-\frac{7520}{9}\right)\zeta_{2}+\left(-\frac{7376}{9}\right)H_{1,0}+\left(-\frac{1384}{3}\right)H_{-1,0,0}+\left(-\frac{904}{3}\right)H_{0,-1,0}+\left(-\frac{2248}{9}\right)H_{1,1}+\left(-\frac{2216}{9}\right)H_{-1,0}+\left(-\frac{268}{3}\right)H_{1,0,0}+\frac{8}{3}H_{0,1,1}+\frac{88}{3}H_{1,1,1}+\frac{112}{3}H_{1,1,0}+\frac{932}{3}H_{0,0,1}+\frac{5584}{9}H_{0,0}+\frac{2216}{3}H_{0,0,0}+\frac{7624}{9}H_{0,1}+\frac{28256}{27}H_{0}+\frac{32362}{27}H_{1}+\left(-\frac{24}{z}\right)H_{-1,0}+\left(-\frac{1468\zeta_{2}}{3}\right)H_{0}+104H_{-1}\zeta_{2}+264H_{1}\zeta_{2}-16H_{-1}\zeta_{3}+480H_{0}\zeta_{3}+656H_{1}\zeta_{3}-712\zeta_{3}+1570\zeta_{4}-32\zeta_{2}H_{-1,-1}$
+$160\zeta_{2}H_{-1,0}-16\zeta_{2}H_{0,-1}-312\zeta_{2}H_{0,0}+496\zeta_{2}H_{0,1}-128\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}+288H_{0,1,0}-96H_{1,0,1}-192H_{-1,-1,-1,0}+192H_{-1,-1,0,0}-64H_{-1,-1,0,1}+352H_{-1,0,-1,0}-336H_{-1,0,0,0}+64H_{-1,0,0,1}-192H_{-1,0,1,0}-64H_{-1,0,1,1}+96H_{0,-1,-1,0}-272H_{0,-1,0,0}+64H_{0,-1,0,1}-496H_{0,0,-1,0}+1264H_{0,0,0,0}-184H_{0,0,0,1}+960H_{0,0,1,0}+320H_{0,0,1,1}-152H_{0,1,0,0}-352H_{0,1,0,1}-224H_{0,1,1,0}-160H_{0,1,1,1}-96H_{1,0,-1,0}+560H_{1,0,0,0}+160H_{1,0,0,1}+640H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}+128H_{1,1,0,1}+192H_{1,1,1,0}+z\left(\left(-\frac{31412}{27}\right)H_{1}+\left(-\frac{20644}{27}\right)H_{0}+\left(-\frac{1472}{3}\right)H_{-1,0,0}+\left(-\frac{992}{3}\right)H_{0,-1,0}+\left(-\frac{2752}{9}\right)H_{-1,0}+\left(-\frac{580}{3}\right)H_{0,0,1}+\left(-\frac{484}{3}\right)H_{0,1,1}+\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{200}{3}\right)H_{1,1,0}+\left(-\frac{172}{3}\right)H_{1,0,0}+\left(-\frac{44}{3}\right)H_{1,1,1}+\frac{1712}{9}H_{1,1}+\frac{5380}{9}H_{0,1}+\frac{5944}{9}H_{1,0}+\frac{10172}{9}H_{0,0}+104H_{-1}\zeta_{2}-156H_{0}\zeta_{2}-352H_{1}\zeta_{2}-796\zeta_{2}-8H_{-1}\zeta_{3}+176H_{0}\zeta_{3}-328H_{1}\zeta_{3}-464\zeta_{3}+143\zeta_{4}-16\zeta_{2}H_{-1,-1}+80\zeta_{2}H_{-1,0}-72\zeta_{2}H_{0,-1}+76\zeta_{2}H_{0,0}+24\zeta_{2}H_{0,1}+64\zeta_{2}H_{1,0}-48\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}-848H_{0,1,0}+96H_{1,0,1}-96H_{-1,-1,-1,0}+96H_{-1,-1,0,0}-32H_{-1,-1,0,1}+176H_{-1,0,-1,0}-168H_{-1,0,0,0}+32H_{-1,0,0,1}-96H_{-1,0,1,0}-32H_{-1,0,1,1}-80H_{0,-1,-1,0}+120H_{0,-1,0,0}+32H_{0,-1,0,1}+56H_{0,0,-1,0}-64H_{0,0,0,0}-76H_{0,0,0,1}+192H_{0,0,1,0}+64H_{0,0,1,1}+116H_{0,1,0,0}+32H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+48H_{1,0,-1,0}-280H_{1,0,0,0}-80H_{1,0,0,1}-320H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}-64H_{1,1,0,1}-96H_{1,1,1,0}-32H_{1,1,1,1}-\frac{13346}{9}\right)+64H_{1,1,1,1}+\frac{13840}{9}\bigg]+C_{A}C_{F}^{2}\bigg[\left(-\frac{52072}{27}\right)H_{1}+\left(-\frac{10868}{9}\right)H_{0,1}+\left(-\frac{1466}{3}\right)H_{0,0}+\left(-\frac{1136}{3}\right)H_{0,0,0}+\left(-\frac{3719}{27}\right)H_{0}+\left(-\frac{376}{3}\right)H_{1,1,1}+\left(-\frac{368}{3}\right)H_{1,1,0}+\left(-\frac{184}{3}\right)H_{0,0,1}+\left(-\frac{344}{9}\right)H_{1,1}+\left(-\frac{88}{3}\right)H_{1,0}+\frac{56}{3}H_{0,1,1}+\frac{272}{3}H_{1,0,0}+\frac{352}{3}H_{1,0,1}+\frac{1}{z}\left(-72H_{-1}\zeta_{2}-24H_{1}\zeta_{2}+72H_{-1,0}-48H_{-1,-1,0}+48H_{-1,0,0}+48H_{-1,0,1}\right)-256H_{-1}\zeta_{2}+296H_{0}\zeta_{2}-240H_{1}\zeta_{2}+716\zeta_{2}-176H_{-1}\zeta_{3}-752H_{0}\zeta_{3}-816H_{1}\zeta_{3}+728\zeta_{3}-2144\zeta_{4}+288\zeta_{2}H_{-1,-1}-192\zeta_{2}H_{-1,0}+136H_{-1,0}-32\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-912\zeta_{2}H_{0,1}+32\zeta_{2}H_{1,0}-192\zeta_{2}H_{1,1}-256H_{-1,-1,0}+144H_{-1,0,0}+128H_{-1,0,1}+96H_{0,-1,0}-472H_{0,1,0}+192H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-192H_{-1,-1,0,1}-160H_{-1,0,-1,0}-16H_{-1,0,0,0}-64H_{-1,0,0,1}+128H_{-1,0,1,0}+64H_{-1,0,1,1}-192H_{0,-1,-1,0}-64H_{0,-1,0,0}-64H_{0,-1,0,1}+96H_{0,0,-1,0}+128H_{0,0,0,0}+752H_{0,0,0,1}-688H_{0,0,1,0}-432H_{0,0,1,1}+416H_{0,1,0,0}+624H_{0,1,0,1}+432H_{0,1,1,0}+368H_{0,1,1,1}+96H_{1,0,-1,0}-48H_{1,0,0,0}+224H_{1,0,0,1}-544H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}-32H_{1,1,0,1}-224H_{1,1,1,0}-128H_{1,1,1,1}+z\left(\left(-\frac{6224}{9}\right)H_{0,1}+\left(-\frac{688}{3}\right)H_{0,0,1}+\left(-\frac{452}{3}\right)H_{1,0,1}+\left(-\frac{448}{3}\right)H_{1,0,0}+\frac{56}{3}H_{1,0}+\frac{68}{3}H_{0,1,1}+\frac{376}{9}H_{1,1}+\frac{296}{3}H_{1,1,1}+\frac{340}{3}H_{1,1,0}+\frac{364}{3}H_{0,0,0}+\frac{50168}{27}H_{1}+\frac{57247}{27}H_{0}-184H_{-1}\zeta_{2}+72H_{0}\zeta_{2}+380H_{1}\zeta_{2}+1176\zeta_{2}-88H_{-1}\zeta_{3}-136H_{0}\zeta_{3}+408H_{1}\zeta_{3}+368\zeta_{3}-152\zeta_{4}+144\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}+64H_{-1,0}-64\zeta_{2}H_{0,-1}-8\zeta_{2}H_{0,0}-889H_{0,0}+136\zeta_{2}H_{0,1}-16\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}-208H_{-1,-1,0}+24H_{-1,0,0}+80H_{-1,0,1}+264H_{0,-1,0}+528H_{0,1,0}+96H_{-1,-1,-1,0}+32H_{-1,-1,0,0}-96H_{-1,-1,0,1}-80H_{-1,0,-1,0}-8H_{-1,0,0,0}-32H_{-1,0,0,1}+64H_{-1,0,1,0}+32H_{-1,0,1,1}-64H_{0,-1,-1,0}-16H_{0,-1,0,0}+32H_{0,-1,0,1}-16H_{0,0,-1,0}+32H_{0,0,0,0}+168H_{0,0,0,1}-104H_{0,0,1,0}-8H_{0,0,1,1}-160H_{0,1,0,0}-168H_{0,1,0,1}-168H_{0,1,1,0}-136H_{0,1,1,1}-48H_{1,0,-1,0}+24H_{1,0,0,0}-112H_{1,0,0,1}+272H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}+16H_{1,1,0,1}+112H_{1,1,1,0}+64H_{1,1,1,1}-\frac{2737}{36}\right)+\frac{1577}{18}\bigg]$
+$N_{f}C_{A}C_{F}\bigg[\left(-\frac{728}{27}\right)H_{0}+\left(-\frac{112}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{88}{9}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,1,1}+\frac{8}{9}\zeta_{2}+\frac{32}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{32}{3}H_{1,1,0}+\frac{112}{9}H_{1,1}+\frac{368}{9}H_{1,0}+\frac{160}{3}H_{1,0,0}+\frac{584}{9}H_{0,0}+\frac{256}{3}H_{0,0,0}+\left(-\frac{32\zeta_{2}}{3}\right)H_{0}+32H_{1}\zeta_{2}-80\zeta_{3}+z\left(\left(-\frac{356}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{208}{9}\right)H_{1,0}+\left(-\frac{518}{27}\right)H_{0}+\left(-\frac{184}{27}\right)H_{1}+\left(-\frac{16}{3}\right)H_{-1,0,0}+\left(-\frac{16}{3}\right)H_{0,-1,0}+\left(-\frac{16}{3}\right)H_{1,1,0}+\left(-\frac{32}{9}\right)H_{-1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\left(-\frac{8}{3}\right)H_{0,1,1}+\frac{8}{3}H_{1,1,1}+\frac{32}{9}H_{0,1}+\frac{16}{3}H_{0,0,0}+\frac{16}{3}H_{0,0,1}-16H_{1}\zeta_{2}-16\zeta_{2}+32\zeta_{3}+\frac{416}{9}\right)-\frac{526}{9}\bigg]\,.$ (131)
 $\Delta P_{qg}^{T,\text{s}(0)}=(4z-2)N_{f}\,$, (132)
 $\Delta P_{qg}^{T,\text{s}(1)}=C_{A}N_{f}\bigg[z\left(-16H_{-1,0}+16H_{0,0}-32H_{0,1}+32H_{1,0}+16H_{1,1}+\left(-\frac{40}{3}\right)H_{1}+\left(-\frac{8}{3}\right)H_{0}+\frac{740}{9}\right)-8H_{-1,0}+24H_{0,0}+16H_{0,1}-16H_{1,0}-8H_{1,1}+\left(-\frac{32}{3}\right)H_{0}+\left(-\frac{4}{3}\right)H_{1}-8\zeta_{2}-\frac{436}{9}\bigg]+C_{F}N_{f}\bigg[z\left(-8H_{0,0}+16H_{0,1}-48H_{1,0}-16H_{1,1}+8H_{1}-16\zeta_{2}-50\right)+4H_{0,0}-8H_{0,1}+24H_{1,0}+8H_{1,1}+22H_{0}+4H_{1}+8\zeta_{2}+32\bigg]+N_{f}^{2}\bigg[\left(\left(-\frac{16}{3}\right)H_{0}+\frac{16}{3}H_{1}-\frac{32}{9}\right)z+\left(-\frac{8}{3}\right)H_{1}+\frac{8}{3}H_{0}-\frac{8}{9}\bigg]\,$, (133)
-$\Delta P_{qg}^{T,\text{s}(2)}=N_{f}^{3}\bigg[\left(-\frac{8}{3}\right)\zeta_{2}+\left(-\frac{8}{3}\right)H_{0,0}+\left(-\frac{8}{3}\right)H_{1,1}+\left(-\frac{8}{9}\right)H_{1}+\frac{8}{9}H_{0}+\frac{8}{3}H_{0,1}+\frac{8}{3}H_{1,0}+z\left(\left(-\frac{16}{3}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1}+\frac{32}{9}H_{0}+\frac{16}{3}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{16}{3}H_{1,1}-\frac{16}{3}\right)+\frac{40}{9}\bigg]+C_{F}N_{f}^{2}\bigg[\left(16\zeta_{2}+16H_{-1,0}-16H_{0,0}\right)z^{2}+\left(\left(-\frac{128}{3}\right)H_{0,0,1}+\left(-\frac{112}{3}\right)H_{1,1,0}+\left(-\frac{280}{9}\right)H_{0,1}+\left(-\frac{260}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,1}+\left(-\frac{80}{3}\right)H_{1,1,1}+\left(-\frac{680}{27}\right)H_{1}+\frac{280}{9}\zeta_{2}+\frac{328}{9}H_{1,1}+\frac{128}{3}H_{0,1,1}+\frac{160}{3}H_{0,1,0}+\frac{232}{3}H_{1,0}+\frac{320}{3}H_{0,0,0}+\frac{5090}{27}H_{0}+\frac{608}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}-16H_{1}\zeta_{2}-64H_{-1,0}+64H_{0,-1,0}+64H_{1,0,0}-64H_{0,0,0,0}-\frac{1157}{3}\right)z$ (139)
-$\left(-\frac{304}{3}\right)\zeta_{3}+\left(-\frac{848}{9}\right)\zeta_{2}+\left(-\frac{80}{3}\right)H_{0,1,0}+\left(-\frac{64}{3}\right)H_{0,1,1}+\left(-\frac{56}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\frac{184}{27}H_{1}+\frac{40}{3}H_{1,0,1}+\frac{40}{3}H_{1,1,1}+\frac{56}{3}H_{1,1,0}+\frac{64}{3}H_{0,0,1}+\frac{272}{9}H_{0,1}+\frac{898}{9}H_{0,0}+\frac{320}{3}H_{0,0,0}+\frac{3830}{27}H_{0}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}+8H_{1}\zeta_{2}-80H_{-1,0}-32H_{0,-1,0}-32H_{1,0,0}+32H_{0,0,0,0}+\frac{6587}{18}\bigg]+N_{f}C_{A}^{2}\bigg[\left(72H_{-1}\zeta_{2}-96H_{0}\zeta_{2}-24H_{1}\zeta_{2}+24\zeta_{2}-120\zeta_{3}+24H_{-1,0}-24H_{0,0}+48H_{-1,-1,0}-48H_{-1,0,0}-48H_{-1,0,1}-48H_{0,-1,0}+48H_{0,0,0}+48H_{0,0,1}\right)z^{2}+\left(\left(-\frac{8836}{9}\right)H_{0,1}+\left(-\frac{976}{3}\right)H_{0,0,0}+\left(-\frac{848}{3}\right)H_{-1,-1,0}+\left(-\frac{2348}{9}\right)H_{0,0}+\left(-\frac{592}{3}\right)H_{1,0,1}+\left(-\frac{592}{3}\right)H_{1,1,0}+\left(-\frac{440}{3}\right)H_{1,1,1}+\left(-\frac{3796}{27}\right)H_{0}+\left(-\frac{404}{3}\right)H_{1,0,0}+\left(-\frac{320}{3}\right)H_{-1,0,1}+\left(-\frac{308}{3}\right)\zeta_{2}+\left(-\frac{304}{3}\right)H_{-1,0}+\frac{284}{3}H_{0,0,1}+\frac{536}{3}H_{-1,0,0}+\frac{5464}{27}H_{1}+\frac{1828}{9}H_{1,1}+\frac{2068}{9}H_{1,0}+\frac{824}{3}H_{0,-1,0}+\left(-\frac{104\zeta_{2}}{3}\right)H_{-1}+\frac{472\zeta_{2}}{3}H_{1}+60H_{0}\zeta_{2}+272H_{-1}\zeta_{3}+32H_{0}\zeta_{3}+496H_{1}\zeta_{3}-184\zeta_{3}+170\zeta_{4}-416\zeta_{2}H_{-1,-1}+224\zeta_{2}H_{-1,0}+144\zeta_{2}H_{0,-1}-56\zeta_{2}H_{0,0}+336\zeta_{2}H_{0,1}+96\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}-24H_{0,1,0}+192H_{0,1,1}-192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}+320H_{-1,-1,0,1}+224H_{-1,0,-1,0}+16H_{-1,0,0,0}-128H_{-1,0,1,0}-64H_{-1,0,1,1}+32H_{0,-1,-1,0}-208H_{0,-1,0,0}-128H_{0,-1,0,1}-112H_{0,0,-1,0}+128H_{0,0,0,0}-200H_{0,0,0,1}+128H_{0,0,1,0}+352H_{0,0,1,1}+184H_{0,1,0,0}-160H_{0,1,1,1}-32H_{1,0,-1,0}-240H_{1,0,0,0}-256H_{1,0,0,1}-64H_{1,0,1,1}-320H_{1,1,0,0}-64H_{1,1,0,1}-64H_{1,1,1,0}+64H_{1,1,1,1}+\frac{5408}{3}\right)z+\left(-\frac{19762}{27}\right)H_{0}+\left(-\frac{9620}{27}\right)H_{1}+\left(-\frac{928}{3}\right)H_{0,0,0}+\left(-\frac{640}{3}\right)H_{-1,-1,0}+\left(-\frac{1502}{9}\right)H_{0,0}+\left(-\frac{352}{3}\right)H_{-1,0,1}+\left(-\frac{566}{9}\right)H_{1,1}+\left(-\frac{434}{9}\right)H_{1,0}+\frac{220}{3}H_{1,1,1}+\frac{232}{3}H_{-1,0}+\frac{304}{3}\zeta_{3}+\frac{680}{3}H_{1,0,1}+\frac{680}{3}H_{1,1,0}+\frac{776}{3}H_{0,-1,0}+\frac{2390}{9}H_{0,1}+\frac{856}{3}H_{-1,0,0}+\frac{1240}{3}H_{1,0,0}+\frac{1342}{3}\zeta_{2}+\left(-\frac{224\zeta_{2}}{3}\right)H_{1}+\frac{32\zeta_{2}}{3}H_{-1}+\frac{856\zeta_{2}}{3}H_{0}+136H_{-1}\zeta_{3}+480H_{0}\zeta_{3}-248H_{1}\zeta_{3}+63\zeta_{4}-208\zeta_{2}H_{-1,-1}+112\zeta_{2}H_{-1,0}-88\zeta_{2}H_{0,-1}+316\zeta_{2}H_{0,0}-216\zeta_{2}H_{0,1}-48\zeta_{2}H_{1,0}+16\zeta_{2}H_{1,1}-40H_{0,0,1}+4H_{0,1,0}-56H_{0,1,1}-96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}+160H_{-1,-1,0,1}+112H_{-1,0,-1,0}+8H_{-1,0,0,0}-64H_{-1,0,1,0}-32H_{-1,0,1,1}-176H_{0,-1,-1,0}+216H_{0,-1,0,0}+248H_{0,0,-1,0}-632H_{0,0,0,0}-388H_{0,0,0,1}-48H_{0,0,1,0}-96H_{0,0,1,1}+204H_{0,1,0,0}+144H_{0,1,0,1}+144H_{0,1,1,0}+128H_{0,1,1,1}+16H_{1,0,-1,0}+120H_{1,0,0,0}+128H_{1,0,0,1}+32H_{1,0,1,1}+160H_{1,1,0,0}+32H_{1,1,0,1}+32H_{1,1,1,0}-32H_{1,1,1,1}-\frac{4670}{3}\bigg]+N_{f}C_{F}^{2}\bigg[\left(48\zeta_{2}+48H_{-1,0}-48H_{0,0}\right)z^{2}+\left(-80\zeta_{2}H_{0}+80\zeta_{3}H_{0}+289H_{0}+600H_{1}-64H_{-1}\zeta_{2}-176H_{1}\zeta_{2}+612\zeta_{2}+992H_{1}\zeta_{3}-368\zeta_{3}+280\zeta_{4}-192\zeta_{2}H_{-1,0}+304H_{-1,0}+128\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-350H_{0,0}+208\zeta_{2}H_{0,1}-340H_{0,1}-224\zeta_{2}H_{1,0}+592H_{1,0}-32\zeta_{2}H_{1,1}+32H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+56H_{0,0,0}+16H_{0,0,1}-16H_{0,1,0}+32H_{0,1,1}-64H_{1,0,0}-48H_{1,0,1}-144H_{1,1,0}-48H_{1,1,1}-384H_{-1,-1,0,0}-192H_{-1,0,-1,0}+352H_{-1,0,0,0}+256H_{0,-1,-1,0}+64H_{0,-1,0,0}-192H_{0,0,-1,0}+32H_{0,0,0,0}+16H_{0,0,0,1}+192H_{0,0,1,0}+128H_{0,0,1,1}-352H_{0,1,0,0}-208H_{0,1,0,1}-240H_{0,1,1,0}-112H_{0,1,1,1}-192H_{1,0,-1,0}+672H_{1,0,0,0}+160H_{1,0,0,1}+608H_{1,0,1,0}+128H_{1,0,1,1}+32H_{1,1,0,0}+160H_{1,1,0,1}+96H_{1,1,1,0}+64H_{1,1,1,1}+\frac{3747}{2}\right)z$ (140)
-$1505H_{0}-582H_{1}-64H_{-1}\zeta_{2}+164H_{0}\zeta_{2}+140H_{1}\zeta_{2}+394\zeta_{2}-152H_{0}\zeta_{3}-496H_{1}\zeta_{3}-548\zeta_{3}-476\zeta_{4}-96\zeta_{2}H_{-1,0}+288H_{-1,0}+32\zeta_{2}H_{0,-1}+72\zeta_{2}H_{0,0}-947H_{0,0}-72\zeta_{2}H_{0,1}-242H_{0,1}+112\zeta_{2}H_{1,0}-390H_{1,0}+16\zeta_{2}H_{1,1}-34H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+144H_{0,-1,0}-476H_{0,0,0}-68H_{0,0,1}-304H_{0,1,0}-80H_{0,1,1}-32H_{1,0,0}-60H_{1,0,1}+12H_{1,1,0}-12H_{1,1,1}-192H_{-1,-1,0,0}-96H_{-1,0,-1,0}+176H_{-1,0,0,0}+64H_{0,-1,-1,0}+64H_{0,-1,0,0}+32H_{0,0,-1,0}-128H_{0,0,0,0}-8H_{0,0,0,1}-96H_{0,0,1,0}-64H_{0,0,1,1}+144H_{0,1,0,0}+104H_{0,1,0,1}+120H_{0,1,1,0}+56H_{0,1,1,1}+96H_{1,0,-1,0}-336H_{1,0,0,0}-80H_{1,0,0,1}-304H_{1,0,1,0}-64H_{1,0,1,1}-16H_{1,1,0,0}-80H_{1,1,0,1}-48H_{1,1,1,0}-32H_{1,1,1,1}-\frac{7379}{4}\bigg]+C_{A}N_{f}^{2}\bigg[\left(-16\zeta_{2}-16H_{-1,0}+16H_{0,0}\right)z^{2}+\left(\left(-\frac{1000}{9}\right)H_{0,0}+\left(-\frac{304}{3}\right)H_{1,0,0}+\left(-\frac{2576}{27}\right)H_{0}+\left(-\frac{784}{9}\right)H_{1,0}+\left(-\frac{640}{9}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,-1,0}+\left(-\frac{128}{3}\right)H_{0,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{32}{3}\right)H_{1,0,1}+\left(-\frac{32}{3}\right)H_{1,1,0}+\frac{16}{3}\zeta_{2}+\frac{64}{3}H_{-1,0,0}+\frac{80}{3}H_{1,1,1}+\frac{128}{3}H_{-1,-1,0}+\frac{160}{3}H_{-1,0}+\frac{592}{9}H_{0,1}+\frac{208}{3}H_{0,0,1}+\frac{2456}{27}H_{1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{1}+\frac{128\zeta_{2}}{3}H_{-1}-16H_{0}\zeta_{2}-80\zeta_{3}+16H_{0,1,0}-64H_{0,1,1}+70\right)z+\left(-\frac{320}{3}\right)H_{0,0,0}+\left(-\frac{748}{9}\right)H_{0,0}+\left(-\frac{902}{27}\right)H_{0}+\left(-\frac{284}{9}\right)H_{0,1}+\left(-\frac{784}{27}\right)H_{1}+\left(-\frac{40}{3}\right)H_{1,1,1}+\left(-\frac{32}{3}\right)H_{-1,0,1}+\frac{16}{3}H_{1,0,1}+\frac{16}{3}H_{1,1,0}+\frac{32}{3}H_{-1,0,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,-1,0}+\frac{296}{9}H_{1,1}+\frac{128}{3}H_{-1,0}+\frac{152}{3}H_{1,0,0}+\frac{176}{3}\zeta_{3}+\frac{196}{3}\zeta_{2}+\frac{728}{9}H_{1,0}+\frac{8\zeta_{2}}{3}H_{0}+\frac{32\zeta_{2}}{3}H_{1}+\frac{64\zeta_{2}}{3}H_{-1}-8H_{0,0,1}+40H_{0,1,0}+48H_{0,1,1}-\frac{988}{9}\bigg]+C_{A}C_{F}N_{f}\bigg[\left(-72\zeta_{2}-72H_{-1,0}+72H_{0,0}\right)z^{2}+\left(\left(-\frac{3484}{3}\right)H_{1,0}+\left(-\frac{25828}{27}\right)H_{1}+\left(-\frac{25589}{27}\right)H_{0}+\left(-\frac{7276}{9}\right)\zeta_{2}+\left(-\frac{2284}{9}\right)H_{1,1}+\left(-\frac{536}{3}\right)H_{0,1,1}+\frac{128}{3}H_{0,0,1}+\frac{272}{3}H_{0,1,0}+\frac{280}{3}H_{0,0,0}+\frac{584}{3}H_{1,1,1}+\frac{784}{3}\zeta_{3}+\frac{848}{3}H_{1,0,1}+\frac{1168}{3}H_{1,1,0}+\frac{5774}{9}H_{0,0}+\frac{10156}{9}H_{0,1}+\frac{80\zeta_{2}}{3}H_{0}+176H_{-1}\zeta_{2}-16H_{1}\zeta_{2}-80H_{-1}\zeta_{3}-112H_{0}\zeta_{3}-1584H_{1}\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{-1,-1}+48H_{-1,0}-128\zeta_{2}H_{0,-1}+64\zeta_{2}H_{0,0}-400\zeta_{2}H_{0,1}+160\zeta_{2}H_{1,0}+64\zeta_{2}H_{1,1}+224H_{-1,-1,0}-288H_{-1,0,0}-64H_{-1,0,1}-176H_{0,-1,0}+344H_{1,0,0}+192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}-64H_{-1,-1,0,1}-160H_{-1,0,-1,0}-48H_{-1,0,0,0}+192H_{-1,0,1,0}+64H_{-1,0,1,1}-256H_{0,-1,-1,0}+288H_{0,0,-1,0}-64H_{0,0,0,0}+128H_{0,0,0,1}-320H_{0,0,1,0}-352H_{0,0,1,1}+272H_{0,1,0,0}+336H_{0,1,0,1}+368H_{0,1,1,0}+272H_{0,1,1,1}+160H_{1,0,-1,0}-112H_{1,0,0,0}+224H_{1,0,0,1}-480H_{1,0,1,0}-64H_{1,0,1,1}+416H_{1,1,0,0}-96H_{1,1,0,1}-32H_{1,1,1,0}-128H_{1,1,1,1}-\frac{9751}{6}\right)z+\left(-\frac{3016}{9}\right)\zeta_{2}+\left(-\frac{788}{3}\right)H_{1,1,0}+\left(-\frac{556}{3}\right)H_{1,0,1}+\left(-\frac{220}{3}\right)H_{0,0,1}+\left(-\frac{184}{3}\right)H_{1,1,1}+\frac{640}{9}H_{0,1}+\frac{244}{3}H_{0,0,0}+\frac{1064}{9}H_{1,1}+\frac{400}{3}H_{0,1,1}+\frac{716}{3}H_{0,1,0}+\frac{4607}{9}H_{0,0}+\frac{1904}{3}H_{1,0}+\frac{2872}{3}\zeta_{3}+\frac{28226}{27}H_{1}+\frac{31993}{27}H_{0}+\left(-\frac{364\zeta_{2}}{3}\right)H_{0}+80H_{-1}\zeta_{2}+12H_{1}\zeta_{2}-40H_{-1}\zeta_{3}+8H_{0}\zeta_{3}+792H_{1}\zeta_{3}+658\zeta_{4}+80\zeta_{2}H_{-1,-1}+8H_{-1,0}-16\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+296\zeta_{2}H_{0,1}-80\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}+128H_{-1,-1,0}-240H_{-1,0,0}-16H_{-1,0,1}+16H_{0,-1,0}-592H_{1,0,0}+96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-32H_{-1,-1,0,1}-80H_{-1,0,-1,0}-24H_{-1,0,0,0}+96H_{-1,0,1,0}+32H_{-1,0,1,1}-32H_{0,-1,-1,0}-48H_{0,-1,0,0}+16H_{0,0,-1,0}-32H_{0,0,0,0}+48H_{0,0,0,1}-80H_{0,0,1,0}+96H_{0,0,1,1}-520H_{0,1,0,0}-312H_{0,1,0,1}-328H_{0,1,1,0}-184H_{0,1,1,1}-80H_{1,0,-1,0}+56H_{1,0,0,0}-112H_{1,0,0,1}+240H_{1,0,1,0}+32H_{1,0,1,1}-208H_{1,1,0,0}+48H_{1,1,0,1}+16H_{1,1,1,0}+64H_{1,1,1,1}+\frac{52313}{36}\bigg]\,$, (141)
+$\Delta P_{qg}^{T,\text{s}(2)}=N_{f}^{3}\bigg[\left(-\frac{8}{3}\right)\zeta_{2}+\left(-\frac{8}{3}\right)H_{0,0}+\left(-\frac{8}{3}\right)H_{1,1}+\left(-\frac{8}{9}\right)H_{1}+\frac{8}{9}H_{0}+\frac{8}{3}H_{0,1}+\frac{8}{3}H_{1,0}+z\left(\left(-\frac{16}{3}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1}+\frac{32}{9}H_{0}+\frac{16}{3}\zeta_{2}+\frac{16}{3}H_{0,0}+\frac{16}{3}H_{1,1}-\frac{16}{3}\right)+\frac{40}{9}\bigg]+C_{F}N_{f}^{2}\bigg[\left(16\zeta_{2}+16H_{-1,0}-16H_{0,0}\right)z^{2}+\left(\left(-\frac{128}{3}\right)H_{0,0,1}+\left(-\frac{112}{3}\right)H_{1,1,0}+\left(-\frac{280}{9}\right)H_{0,1}+\left(-\frac{260}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,1}+\left(-\frac{80}{3}\right)H_{1,1,1}+\left(-\frac{680}{27}\right)H_{1}+\frac{280}{9}\zeta_{2}+\frac{328}{9}H_{1,1}+\frac{128}{3}H_{0,1,1}+\frac{160}{3}H_{0,1,0}+\frac{232}{3}H_{1,0}+\frac{320}{3}H_{0,0,0}+\frac{5090}{27}H_{0}+\frac{608}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}-16H_{1}\zeta_{2}-64H_{-1,0}+64H_{0,-1,0}+64H_{1,0,0}-64H_{0,0,0,0}-\frac{1157}{3}\right)z$
+$\left(-\frac{304}{3}\right)\zeta_{3}+\left(-\frac{848}{9}\right)\zeta_{2}+\left(-\frac{80}{3}\right)H_{0,1,0}+\left(-\frac{64}{3}\right)H_{0,1,1}+\left(-\frac{56}{3}\right)H_{1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\frac{184}{27}H_{1}+\frac{40}{3}H_{1,0,1}+\frac{40}{3}H_{1,1,1}+\frac{56}{3}H_{1,1,0}+\frac{64}{3}H_{0,0,1}+\frac{272}{9}H_{0,1}+\frac{898}{9}H_{0,0}+\frac{320}{3}H_{0,0,0}+\frac{3830}{27}H_{0}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}+8H_{1}\zeta_{2}-80H_{-1,0}-32H_{0,-1,0}-32H_{1,0,0}+32H_{0,0,0,0}+\frac{6587}{18}\bigg]+N_{f}C_{A}^{2}\bigg[\left(72H_{-1}\zeta_{2}-96H_{0}\zeta_{2}-24H_{1}\zeta_{2}+24\zeta_{2}-120\zeta_{3}+24H_{-1,0}-24H_{0,0}+48H_{-1,-1,0}-48H_{-1,0,0}-48H_{-1,0,1}-48H_{0,-1,0}+48H_{0,0,0}+48H_{0,0,1}\right)z^{2}+\left(\left(-\frac{8836}{9}\right)H_{0,1}+\left(-\frac{976}{3}\right)H_{0,0,0}+\left(-\frac{848}{3}\right)H_{-1,-1,0}+\left(-\frac{2348}{9}\right)H_{0,0}+\left(-\frac{592}{3}\right)H_{1,0,1}+\left(-\frac{592}{3}\right)H_{1,1,0}+\left(-\frac{440}{3}\right)H_{1,1,1}+\left(-\frac{3796}{27}\right)H_{0}+\left(-\frac{404}{3}\right)H_{1,0,0}+\left(-\frac{320}{3}\right)H_{-1,0,1}+\left(-\frac{308}{3}\right)\zeta_{2}+\left(-\frac{304}{3}\right)H_{-1,0}+\frac{284}{3}H_{0,0,1}+\frac{536}{3}H_{-1,0,0}+\frac{5464}{27}H_{1}+\frac{1828}{9}H_{1,1}+\frac{2068}{9}H_{1,0}+\frac{824}{3}H_{0,-1,0}+\left(-\frac{104\zeta_{2}}{3}\right)H_{-1}+\frac{472\zeta_{2}}{3}H_{1}+60H_{0}\zeta_{2}+272H_{-1}\zeta_{3}+32H_{0}\zeta_{3}+496H_{1}\zeta_{3}-184\zeta_{3}+170\zeta_{4}-416\zeta_{2}H_{-1,-1}+224\zeta_{2}H_{-1,0}+144\zeta_{2}H_{0,-1}-56\zeta_{2}H_{0,0}+336\zeta_{2}H_{0,1}+96\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}-24H_{0,1,0}+192H_{0,1,1}-192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}+320H_{-1,-1,0,1}+224H_{-1,0,-1,0}+16H_{-1,0,0,0}-128H_{-1,0,1,0}-64H_{-1,0,1,1}+32H_{0,-1,-1,0}-208H_{0,-1,0,0}-128H_{0,-1,0,1}-112H_{0,0,-1,0}+128H_{0,0,0,0}-200H_{0,0,0,1}+128H_{0,0,1,0}+352H_{0,0,1,1}+184H_{0,1,0,0}-160H_{0,1,1,1}-32H_{1,0,-1,0}-240H_{1,0,0,0}-256H_{1,0,0,1}-64H_{1,0,1,1}-320H_{1,1,0,0}-64H_{1,1,0,1}-64H_{1,1,1,0}+64H_{1,1,1,1}+\frac{5408}{3}\right)z+\left(-\frac{19762}{27}\right)H_{0}+\left(-\frac{9620}{27}\right)H_{1}+\left(-\frac{928}{3}\right)H_{0,0,0}+\left(-\frac{640}{3}\right)H_{-1,-1,0}+\left(-\frac{1502}{9}\right)H_{0,0}+\left(-\frac{352}{3}\right)H_{-1,0,1}+\left(-\frac{566}{9}\right)H_{1,1}+\left(-\frac{434}{9}\right)H_{1,0}+\frac{220}{3}H_{1,1,1}+\frac{232}{3}H_{-1,0}+\frac{304}{3}\zeta_{3}+\frac{680}{3}H_{1,0,1}+\frac{680}{3}H_{1,1,0}+\frac{776}{3}H_{0,-1,0}+\frac{2390}{9}H_{0,1}+\frac{856}{3}H_{-1,0,0}+\frac{1240}{3}H_{1,0,0}+\frac{1342}{3}\zeta_{2}+\left(-\frac{224\zeta_{2}}{3}\right)H_{1}+\frac{32\zeta_{2}}{3}H_{-1}+\frac{856\zeta_{2}}{3}H_{0}+136H_{-1}\zeta_{3}+480H_{0}\zeta_{3}-248H_{1}\zeta_{3}+63\zeta_{4}-208\zeta_{2}H_{-1,-1}+112\zeta_{2}H_{-1,0}-88\zeta_{2}H_{0,-1}+316\zeta_{2}H_{0,0}-216\zeta_{2}H_{0,1}-48\zeta_{2}H_{1,0}+16\zeta_{2}H_{1,1}-40H_{0,0,1}+4H_{0,1,0}-56H_{0,1,1}-96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}+160H_{-1,-1,0,1}+112H_{-1,0,-1,0}+8H_{-1,0,0,0}-64H_{-1,0,1,0}-32H_{-1,0,1,1}-176H_{0,-1,-1,0}+216H_{0,-1,0,0}+248H_{0,0,-1,0}-632H_{0,0,0,0}-388H_{0,0,0,1}-48H_{0,0,1,0}-96H_{0,0,1,1}+204H_{0,1,0,0}+144H_{0,1,0,1}+144H_{0,1,1,0}+128H_{0,1,1,1}+16H_{1,0,-1,0}+120H_{1,0,0,0}+128H_{1,0,0,1}+32H_{1,0,1,1}+160H_{1,1,0,0}+32H_{1,1,0,1}+32H_{1,1,1,0}-32H_{1,1,1,1}-\frac{4670}{3}\bigg]+N_{f}C_{F}^{2}\bigg[\left(48\zeta_{2}+48H_{-1,0}-48H_{0,0}\right)z^{2}+\left(-80\zeta_{2}H_{0}+80\zeta_{3}H_{0}+289H_{0}+600H_{1}-64H_{-1}\zeta_{2}-176H_{1}\zeta_{2}+612\zeta_{2}+992H_{1}\zeta_{3}-368\zeta_{3}+280\zeta_{4}-192\zeta_{2}H_{-1,0}+304H_{-1,0}+128\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-350H_{0,0}+208\zeta_{2}H_{0,1}-340H_{0,1}-224\zeta_{2}H_{1,0}+592H_{1,0}-32\zeta_{2}H_{1,1}+32H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+56H_{0,0,0}+16H_{0,0,1}-16H_{0,1,0}+32H_{0,1,1}-64H_{1,0,0}-48H_{1,0,1}-144H_{1,1,0}-48H_{1,1,1}-384H_{-1,-1,0,0}-192H_{-1,0,-1,0}+352H_{-1,0,0,0}+256H_{0,-1,-1,0}+64H_{0,-1,0,0}-192H_{0,0,-1,0}+32H_{0,0,0,0}+16H_{0,0,0,1}+192H_{0,0,1,0}+128H_{0,0,1,1}-352H_{0,1,0,0}-208H_{0,1,0,1}-240H_{0,1,1,0}-112H_{0,1,1,1}-192H_{1,0,-1,0}+672H_{1,0,0,0}+160H_{1,0,0,1}+608H_{1,0,1,0}+128H_{1,0,1,1}+32H_{1,1,0,0}+160H_{1,1,0,1}+96H_{1,1,1,0}+64H_{1,1,1,1}+\frac{3747}{2}\right)z$
+$1505H_{0}-582H_{1}-64H_{-1}\zeta_{2}+164H_{0}\zeta_{2}+140H_{1}\zeta_{2}+394\zeta_{2}-152H_{0}\zeta_{3}-496H_{1}\zeta_{3}-548\zeta_{3}-476\zeta_{4}-96\zeta_{2}H_{-1,0}+288H_{-1,0}+32\zeta_{2}H_{0,-1}+72\zeta_{2}H_{0,0}-947H_{0,0}-72\zeta_{2}H_{0,1}-242H_{0,1}+112\zeta_{2}H_{1,0}-390H_{1,0}+16\zeta_{2}H_{1,1}-34H_{1,1}-128H_{-1,-1,0}+352H_{-1,0,0}+144H_{0,-1,0}-476H_{0,0,0}-68H_{0,0,1}-304H_{0,1,0}-80H_{0,1,1}-32H_{1,0,0}-60H_{1,0,1}+12H_{1,1,0}-12H_{1,1,1}-192H_{-1,-1,0,0}-96H_{-1,0,-1,0}+176H_{-1,0,0,0}+64H_{0,-1,-1,0}+64H_{0,-1,0,0}+32H_{0,0,-1,0}-128H_{0,0,0,0}-8H_{0,0,0,1}-96H_{0,0,1,0}-64H_{0,0,1,1}+144H_{0,1,0,0}+104H_{0,1,0,1}+120H_{0,1,1,0}+56H_{0,1,1,1}+96H_{1,0,-1,0}-336H_{1,0,0,0}-80H_{1,0,0,1}-304H_{1,0,1,0}-64H_{1,0,1,1}-16H_{1,1,0,0}-80H_{1,1,0,1}-48H_{1,1,1,0}-32H_{1,1,1,1}-\frac{7379}{4}\bigg]+C_{A}N_{f}^{2}\bigg[\left(-16\zeta_{2}-16H_{-1,0}+16H_{0,0}\right)z^{2}+\left(\left(-\frac{1000}{9}\right)H_{0,0}+\left(-\frac{304}{3}\right)H_{1,0,0}+\left(-\frac{2576}{27}\right)H_{0}+\left(-\frac{784}{9}\right)H_{1,0}+\left(-\frac{640}{9}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,-1,0}+\left(-\frac{128}{3}\right)H_{0,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{32}{3}\right)H_{1,0,1}+\left(-\frac{32}{3}\right)H_{1,1,0}+\frac{16}{3}\zeta_{2}+\frac{64}{3}H_{-1,0,0}+\frac{80}{3}H_{1,1,1}+\frac{128}{3}H_{-1,-1,0}+\frac{160}{3}H_{-1,0}+\frac{592}{9}H_{0,1}+\frac{208}{3}H_{0,0,1}+\frac{2456}{27}H_{1}+\left(-\frac{64\zeta_{2}}{3}\right)H_{1}+\frac{128\zeta_{2}}{3}H_{-1}-16H_{0}\zeta_{2}-80\zeta_{3}+16H_{0,1,0}-64H_{0,1,1}+70\right)z+\left(-\frac{320}{3}\right)H_{0,0,0}+\left(-\frac{748}{9}\right)H_{0,0}+\left(-\frac{902}{27}\right)H_{0}+\left(-\frac{284}{9}\right)H_{0,1}+\left(-\frac{784}{27}\right)H_{1}+\left(-\frac{40}{3}\right)H_{1,1,1}+\left(-\frac{32}{3}\right)H_{-1,0,1}+\frac{16}{3}H_{1,0,1}+\frac{16}{3}H_{1,1,0}+\frac{32}{3}H_{-1,0,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,-1,0}+\frac{296}{9}H_{1,1}+\frac{128}{3}H_{-1,0}+\frac{152}{3}H_{1,0,0}+\frac{176}{3}\zeta_{3}+\frac{196}{3}\zeta_{2}+\frac{728}{9}H_{1,0}+\frac{8\zeta_{2}}{3}H_{0}+\frac{32\zeta_{2}}{3}H_{1}+\frac{64\zeta_{2}}{3}H_{-1}-8H_{0,0,1}+40H_{0,1,0}+48H_{0,1,1}-\frac{988}{9}\bigg]+C_{A}C_{F}N_{f}\bigg[\left(-72\zeta_{2}-72H_{-1,0}+72H_{0,0}\right)z^{2}+\left(\left(-\frac{3484}{3}\right)H_{1,0}+\left(-\frac{25828}{27}\right)H_{1}+\left(-\frac{25589}{27}\right)H_{0}+\left(-\frac{7276}{9}\right)\zeta_{2}+\left(-\frac{2284}{9}\right)H_{1,1}+\left(-\frac{536}{3}\right)H_{0,1,1}+\frac{128}{3}H_{0,0,1}+\frac{272}{3}H_{0,1,0}+\frac{280}{3}H_{0,0,0}+\frac{584}{3}H_{1,1,1}+\frac{784}{3}\zeta_{3}+\frac{848}{3}H_{1,0,1}+\frac{1168}{3}H_{1,1,0}+\frac{5774}{9}H_{0,0}+\frac{10156}{9}H_{0,1}+\frac{80\zeta_{2}}{3}H_{0}+176H_{-1}\zeta_{2}-16H_{1}\zeta_{2}-80H_{-1}\zeta_{3}-112H_{0}\zeta_{3}-1584H_{1}\zeta_{3}-44\zeta_{4}+160\zeta_{2}H_{-1,-1}+48H_{-1,0}-128\zeta_{2}H_{0,-1}+64\zeta_{2}H_{0,0}-400\zeta_{2}H_{0,1}+160\zeta_{2}H_{1,0}+64\zeta_{2}H_{1,1}+224H_{-1,-1,0}-288H_{-1,0,0}-64H_{-1,0,1}-176H_{0,-1,0}+344H_{1,0,0}+192H_{-1,-1,-1,0}+128H_{-1,-1,0,0}-64H_{-1,-1,0,1}-160H_{-1,0,-1,0}-48H_{-1,0,0,0}+192H_{-1,0,1,0}+64H_{-1,0,1,1}-256H_{0,-1,-1,0}+288H_{0,0,-1,0}-64H_{0,0,0,0}+128H_{0,0,0,1}-320H_{0,0,1,0}-352H_{0,0,1,1}+272H_{0,1,0,0}+336H_{0,1,0,1}+368H_{0,1,1,0}+272H_{0,1,1,1}+160H_{1,0,-1,0}-112H_{1,0,0,0}+224H_{1,0,0,1}-480H_{1,0,1,0}-64H_{1,0,1,1}+416H_{1,1,0,0}-96H_{1,1,0,1}-32H_{1,1,1,0}-128H_{1,1,1,1}-\frac{9751}{6}\right)z+\left(-\frac{3016}{9}\right)\zeta_{2}+\left(-\frac{788}{3}\right)H_{1,1,0}+\left(-\frac{556}{3}\right)H_{1,0,1}+\left(-\frac{220}{3}\right)H_{0,0,1}+\left(-\frac{184}{3}\right)H_{1,1,1}+\frac{640}{9}H_{0,1}+\frac{244}{3}H_{0,0,0}+\frac{1064}{9}H_{1,1}+\frac{400}{3}H_{0,1,1}+\frac{716}{3}H_{0,1,0}+\frac{4607}{9}H_{0,0}+\frac{1904}{3}H_{1,0}+\frac{2872}{3}\zeta_{3}+\frac{28226}{27}H_{1}+\frac{31993}{27}H_{0}+\left(-\frac{364\zeta_{2}}{3}\right)H_{0}+80H_{-1}\zeta_{2}+12H_{1}\zeta_{2}-40H_{-1}\zeta_{3}+8H_{0}\zeta_{3}+792H_{1}\zeta_{3}+658\zeta_{4}+80\zeta_{2}H_{-1,-1}+8H_{-1,0}-16\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+296\zeta_{2}H_{0,1}-80\zeta_{2}H_{1,0}-32\zeta_{2}H_{1,1}+128H_{-1,-1,0}-240H_{-1,0,0}-16H_{-1,0,1}+16H_{0,-1,0}-592H_{1,0,0}+96H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-32H_{-1,-1,0,1}-80H_{-1,0,-1,0}-24H_{-1,0,0,0}+96H_{-1,0,1,0}+32H_{-1,0,1,1}-32H_{0,-1,-1,0}-48H_{0,-1,0,0}+16H_{0,0,-1,0}-32H_{0,0,0,0}+48H_{0,0,0,1}-80H_{0,0,1,0}+96H_{0,0,1,1}-520H_{0,1,0,0}-312H_{0,1,0,1}-328H_{0,1,1,0}-184H_{0,1,1,1}-80H_{1,0,-1,0}+56H_{1,0,0,0}-112H_{1,0,0,1}+240H_{1,0,1,0}+32H_{1,0,1,1}-208H_{1,1,0,0}+48H_{1,1,0,1}+16H_{1,1,1,0}+64H_{1,1,1,1}+\frac{52313}{36}\bigg]\,.$ (134)
 $\Delta P_{gg}^{T,\text{s}(0)}=4C_{A}\left[\frac{1}{1-z}\right]_{+}+\delta(1-z)\left(\frac{11C_{A}}{3}-\frac{2N_{f}}{3}\right)+\left(-\frac{8z^{2}}{z-1}+\frac{12z}{z-1}-\frac{4}{z-1}\right)C_{A}\,$, (135)
 $\Delta P_{gg}^{T,\text{s}(1)}=\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\frac{268}{9}-8\zeta_{2}\right)C_{A}^{2}-\frac{40C_{A}N_{f}}{9}\bigg]+C_{A}^{2}\bigg[\frac{z^{2}}{z-1}\left(-64H_{0,0}+\frac{1}{z+1}-\frac{74}{9}\right)+\frac{z^{2}}{z-1}\left(32H_{0,1}+32H_{1,0}-20H_{0}\right)+\frac{32z^{2}}{z+1}H_{-1,0}+\frac{1}{z-1}\left(32H_{0,1}+32H_{1,0}+\frac{92}{3}H_{0}-8\zeta_{2}+\frac{268}{9}\right)+\frac{z}{z-1}\left(16H_{0,0}+\frac{1}{z+1}-16\zeta_{2}-\frac{74}{9}\right)+\frac{z}{z-1}\left(-48H_{0,1}-48H_{1,0}-40H_{0}\right)+\frac{32}{z+1}H_{-1,0}+\frac{1}{(z-1)(z+1)}\left(96H_{0,0}-\frac{194}{9}\right)+\frac{48z}{z+1}H_{-1,0}+\frac{z^{3}\left(32\zeta_{2}-\frac{194}{9}\right)}{(z-1)(z+1)}\bigg]+C_{F}N_{f}\bigg[z\left(8H_{0,0}+4H_{0}-28\right)+8H_{0,0}+28H_{0}+28\bigg]+\delta(1-z)\bigg[-\frac{8C_{A}N_{f}}{3}+\left(12\zeta_{3}+\frac{32}{3}\right)C_{A}^{2}-2C_{F}N_{f}\bigg]+C_{A}N_{f}\bigg[\frac{z^{2}}{z-1}\left(8H_{0}+\frac{152}{9}\right)+\frac{1}{z-1}\left(\frac{40}{3}H_{0}+\frac{152}{9}\right)+\frac{z}{z-1}\left(-16H_{0}-\frac{88}{3}\right)-\frac{40}{9(z-1)}\bigg]\,$, (136)
-$\Delta P_{gg}^{T,\text{s}(2)}=C_{A}^{3}\bigg[\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)z^{2}+\left(336H_{-1,-1,0}+368H_{-1,0,1}\right)z+\frac{1}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{1072}{9}\right)\zeta_{2}+\left(-\frac{8}{9}\right)H_{1}+\frac{88}{3}\zeta_{3}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}+88\zeta_{4}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-208H_{0,0,1}-144H_{0,1,0}-256H_{0,0,1,1}-608H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}+\frac{490}{3}\right)+\frac{1}{z}\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{48478}{27}\right)H_{0}+\left(-\frac{2968}{3}\right)\zeta_{3}+\left(-\frac{320}{3}\right)H_{0,1}+\left(-\frac{284}{3}\right)H_{0,0}+\left(-\frac{592}{9}\right)\zeta_{2}+\frac{320}{3}H_{0,-1,0}+\frac{280\zeta_{2}}{3}H_{0}+472\zeta_{4}-96\zeta_{2}H_{0,0}+160\zeta_{2}H_{0,1}-112H_{0,0,0}+64H_{0,-1,-1,0}-224H_{0,-1,0,0}-128H_{0,-1,0,1}-64H_{0,0,-1,0}+128H_{0,0,0,0}-160H_{0,0,0,1}-64H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)+\frac{z}{z-1}\left(\left(-\frac{1232}{3}\right)H_{1,0}+\frac{16}{9}H_{1}-336H_{1}\zeta_{2}-288H_{1}\zeta_{3}+416\zeta_{2}H_{0,-1}+96\zeta_{2}H_{1,0}+584H_{0,0,1}+320H_{0,1,0}+568H_{1,0,0}+384H_{0,0,1,1}+672H_{0,1,0,0}+384H_{0,1,0,1}+384H_{0,1,1,0}-192H_{1,0,-1,0}+960H_{1,0,0,0}+384H_{1,0,0,1}+384H_{1,0,1,0}+384H_{1,1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3232}{3}\right)H_{0,-1,0}+\left(-\frac{9136}{9}\right)\zeta_{2}+\left(-\frac{592}{9}\right)H_{0,1}+\frac{16964}{9}H_{0,0}+\frac{6736}{3}H_{0,0,0}+\frac{8990}{3}H_{0}+\left(-\frac{2024\zeta_{2}}{3}\right)H_{0}-800H_{0}\zeta_{3}-920\zeta_{3}-576\zeta_{4}-608\zeta_{2}H_{0,0}-32\zeta_{2}H_{0,1}+320H_{0,-1,-1,0}-544H_{0,-1,0,0}-704H_{0,0,-1,0}+1984H_{0,0,0,0}+800H_{0,0,0,1}+512H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)$ (144)
-$\frac{z^{2}}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{8}{9}\right)H_{1}+\frac{176}{3}H_{0,1,0}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-112H_{0,0,1}-256H_{0,0,1,1}-288H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,1}+\frac{88\zeta_{2}}{3}H_{1}\right)+\frac{z^{3}}{z-1}\left(\frac{1060}{9}H_{0,0}+\frac{3104}{9}H_{0,1}+\frac{2792}{3}\zeta_{3}+\frac{13402}{9}H_{0}+\frac{160\zeta_{2}}{3}H_{0}+304\zeta_{2}-680\zeta_{4}+96\zeta_{2}H_{0,0}-256\zeta_{2}H_{0,1}-48H_{0,-1,0}+464H_{0,0,0}+128H_{0,-1,0,0}+256H_{0,-1,0,1}-64H_{0,0,-1,0}-256H_{0,0,0,0}-96H_{0,0,0,1}-128H_{0,0,1,0}+\frac{1}{z+1}-\frac{47572}{9}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,0}+\frac{176}{3}H_{0,-1,0}+\frac{440}{3}\zeta_{3}+\frac{352\zeta_{2}}{3}H_{0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{89182}{27}\right)H_{0}+\left(-\frac{14612}{9}\right)H_{0,0}+\left(-\frac{4448}{3}\right)H_{0,0,0}+\frac{2144}{3}\zeta_{3}+\frac{2528}{3}H_{0,-1,0}+\frac{9136}{9}\zeta_{2}+528H_{0}\zeta_{2}+1088H_{0}\zeta_{3}+808\zeta_{4}+864\zeta_{2}H_{0,0}+304H_{0,1}-512H_{0,-1,-1,0}+768H_{0,-1,0,0}+960H_{0,0,-1,0}-2496H_{0,0,0,0}-1184H_{0,0,0,1}-768H_{0,0,1,0}-\frac{47572}{9}\right)+\frac{1}{z(z+1)}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{z}{z+1}\left(\left(-\frac{4736}{3}\right)H_{-1,0}-400H_{-1}\zeta_{2}-576H_{-1}\zeta_{3}+768\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}-1312H_{-1,0,0}+384H_{-1,-1,0,0}-768H_{-1,-1,0,1}+384H_{-1,0,-1,0}-960H_{-1,0,0,0}-192H_{-1,0,1,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{z^{3}}{z+1}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{88\zeta_{2}}{3(z-1)z}H_{1}+336H_{-1,-1,0}+368H_{-1,0,1}\bigg]+N_{f}C_{A}^{2}\bigg[\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\frac{16\zeta_{2}}{3}H_{1}-16H_{-1}\zeta_{2}\right)z^{2}+32H_{-1,0,1}+96H_{0,-1,-1,0}+\left(\left(-\frac{2174}{9}\right)H_{1}+56H_{-1}\zeta_{2}+88H_{1}\zeta_{2}-48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+32H_{-1,0,1}-96H_{0,-1,-1,0}+48H_{0,-1,0,0}+80H_{0,0,-1,0}+24H_{0,0,0,1}+24H_{0,1,0,0}\right)z+\frac{2174}{9}H_{1}-48H_{0,-1,0,0}-16H_{0,0,-1,0}+\frac{1}{z-1}\left(\left(-\frac{1844}{9}\right)H_{0,1}+\left(-\frac{356}{3}\right)H_{1,0,0}+\left(-\frac{928}{9}\right)H_{1,0}+\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}+48\zeta_{2}H_{0,1}-124H_{0,0,1}-96H_{0,1,0}-\frac{836}{27}\right)+16H_{0,0,0,0}+24H_{0,0,0,1}+24H_{0,1,0,0}+\frac{1}{z}\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\left(-\frac{16\zeta_{2}}{3}\right)H_{1}-16H_{-1}\zeta_{2}\right)+\frac{z}{z-1}\left(\left(-\frac{1396}{9}\right)H_{0,0}+\left(-\frac{368}{3}\right)\zeta_{3}+\left(-\frac{1138}{27}\right)H_{0}+\frac{376}{3}H_{0,-1,0}+\frac{508}{3}\zeta_{2}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-138\zeta_{4}+24\zeta_{2}H_{0,0}-32H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+56H_{-1}\zeta_{2}-88H_{1}\zeta_{2}+48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}$ (145)
-$\frac{z}{z-1}\left(\frac{512}{3}H_{0,1}+\frac{512}{3}H_{1,0}+200H_{0,0,1}+128H_{0,1,0}+200H_{1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{356}{3}\right)H_{1,0,0}+\left(-\frac{928}{9}\right)H_{1,0}+\left(-\frac{308}{3}\right)H_{0,0,1}+\left(-\frac{224}{3}\right)H_{0,1,0}+\left(-\frac{4}{3}\right)H_{0,1}-48\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{832}{3}\right)\zeta_{3}+\left(-\frac{2140}{9}\right)\zeta_{2}+\frac{1880}{3}H_{0,0}+\frac{6308}{9}H_{0}-116H_{0}\zeta_{2}-96H_{0}\zeta_{3}+78\zeta_{4}-40\zeta_{2}H_{0,0}-104H_{0,-1,0}+392H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+\left(-\frac{32z^{3}}{3(z-1)}\right)H_{0,0,1}+\frac{z^{3}}{z-1}\left(\left(-\frac{1844}{9}\right)\zeta_{2}+\frac{1246}{9}H_{0}+\frac{496}{3}\zeta_{3}+\frac{1780}{9}H_{0,0}+\frac{20\zeta_{2}}{3}H_{0}+138\zeta_{4}-24\zeta_{2}H_{0,0}-136H_{0,-1,0}-32H_{0,0,0}+\frac{1}{z+1}-\frac{22904}{27}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{32}{3}H_{0,-1,0}+\frac{80}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}+16\zeta_{2}-16H_{0,0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{16324}{27}\right)H_{0}+\left(-\frac{5432}{9}\right)H_{0,0}+\left(-\frac{1528}{3}\right)H_{0,0,0}+\frac{376}{3}H_{0,-1,0}+\frac{1996}{9}\zeta_{2}+\frac{880}{3}\zeta_{3}+100H_{0}\zeta_{2}+96H_{0}\zeta_{3}-78\zeta_{4}+40\zeta_{2}H_{0,0}-\frac{22904}{27}\right)+\frac{1}{z(z+1)}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)+\frac{z}{z+1}\left(\left(-\frac{1232}{3}\right)H_{-1,0}-272H_{-1,0,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{z^{3}}{z+1}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)\bigg]+C_{A}N_{f}^{2}\bigg[\left(-\frac{82}{9}\right)H_{1}+\left(-\frac{32}{9}\right)H_{0,1}+z\left(\left(-\frac{32}{9}\right)H_{0,1}+\frac{82}{9}H_{1}\right)+\frac{1}{z-1}\left(\left(-\frac{176}{9}\right)H_{0,0}+\left(-\frac{496}{27}\right)H_{0}+\left(-\frac{32}{9}\right)\zeta_{2}+\frac{230}{27}\right)+\frac{z}{z-1}\left(\frac{64}{3}H_{0,0}+30H_{0}-\frac{148}{9}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{158}{9}\right)H_{0}+\left(-\frac{80}{9}\right)H_{0,0}+\frac{32}{9}\zeta_{2}+\frac{230}{27}\right)-\frac{16}{27(z-1)}\bigg]+N_{f}C_{F}C_{A}\bigg[\left(\left(-\frac{64}{3}\right)H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{64}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,0,0}+\frac{64}{3}H_{0,0,1}+\left(-\frac{32\zeta_{2}}{3}\right)H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}+48H_{0,0}\right)z^{2}+\left(\left(-\frac{1756}{9}\right)H_{1}+\left(-\frac{488}{3}\right)H_{0,1,0}+\left(-\frac{296}{3}\right)H_{0,0,0}+\left(-\frac{152}{3}\right)H_{0,1,1}+\frac{260}{3}H_{1,1}+\frac{2932}{9}H_{0,1}+\frac{1004}{3}H_{1,0}+\frac{1352}{3}H_{-1,0}+\frac{2168}{3}H_{0,0}-176H_{-1}\zeta_{2}-640H_{1}\zeta_{2}+128\zeta_{2}H_{0,-1}-16\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+272H_{0,-1,0}+8H_{0,0,1}+712H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+256H_{0,-1,-1,0}-192H_{0,-1,0,0}-256H_{0,0,-1,0}+128H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)z+\left(-\frac{3632}{3}\right)H_{0,0}+\left(-\frac{3056}{3}\right)H_{0,0,0}+\left(-\frac{1004}{3}\right)H_{1,0}+\left(-\frac{260}{3}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,1,0}+\frac{40}{9}H_{0,1}+\frac{88}{3}H_{0,1,1}+\frac{1756}{9}H_{1}+\frac{1352}{3}H_{-1,0}+\frac{1}{z-1}\left(\left(-\frac{1936}{3}\right)\zeta_{3}+\left(-\frac{2672}{9}\right)\zeta_{2}+\frac{51500}{27}H_{0}+\left(-\frac{416\zeta_{2}}{3}\right)H_{0}-512H_{0}\zeta_{3}-468\zeta_{4}+\frac{88478}{27}\right)+\frac{1}{z}\left(\left(-\frac{64}{3}\right)H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\frac{64}{3}H_{-1,-1,0}+\frac{32\zeta_{2}}{3}H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{3148}{9}\right)H_{0}+\frac{1420}{3}\zeta_{2}+168H_{0}\zeta_{2}+352H_{0}\zeta_{3}+1240\zeta_{3}+160\zeta_{4}-\frac{177946}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{41624}{27}\right)H_{0}+\left(-\frac{1624}{3}\right)\zeta_{3}+\left(-\frac{1156}{9}\right)\zeta_{2}+\frac{40\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+308\zeta_{4}+\frac{89468}{27}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{160}{3}\right)\zeta_{3}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}\right)\bigg]$ (146)
-$176H_{-1}\zeta_{2}+640H_{1}\zeta_{2}-128\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+368H_{0,-1,0}-32H_{0,0,1}-712H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-256H_{0,-1,-1,0}+192H_{0,-1,0,0}+128H_{0,0,-1,0}-448H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}+\delta(1-z)\bigg[\left(\frac{8}{3}\zeta_{2}+\frac{55}{3}\zeta_{4}+\frac{536}{3}\zeta_{3}-16\zeta_{2}\zeta_{3}-80\zeta_{5}+\frac{79}{2}\right)C_{A}^{3}+\left(\left(-\frac{80}{3}\right)\zeta_{3}+\left(-\frac{10}{3}\right)\zeta_{4}+\left(-\frac{8}{3}\right)\zeta_{2}-\frac{233}{18}\right)N_{f}C_{A}^{2}+\left(\frac{29N_{f}^{2}}{18}-\frac{241C_{F}N_{f}}{18}\right)C_{A}+\frac{11}{9}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{3}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)N_{f}C_{A}^{2}+\left(C_{F}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)-\frac{16}{27}N_{f}^{2}\right)C_{A}\bigg]+C_{F}N_{f}^{2}\bigg[\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{40}{3}\right)H_{1,0}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{328}{27}\right)H_{0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{136}{9}\zeta_{2}+\frac{224}{9}H_{1}+\frac{344}{9}H_{0,1}+\frac{16\zeta_{2}}{3}H_{0}-32H_{0,0}+16H_{0,0,1}+z\left(\left(-\frac{488}{9}\right)\zeta_{2}+\left(-\frac{224}{9}\right)H_{1}+\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{8}{9}H_{0,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,1}+\frac{776}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}+16H_{0,0,1}-\frac{992}{27}\right)+\frac{992}{27}\bigg]+C_{F}^{2}N_{f}\bigg[\left(32\zeta_{2}+32H_{-1,0}-32H_{0,0}\right)z^{2}+\left(-48\zeta_{2}H_{0}-160\zeta_{3}H_{0}+1032H_{0}+580H_{1}+128H_{-1}\zeta_{2}+528H_{1}\zeta_{2}+580\zeta_{2}-48\zeta_{3}-380\zeta_{4}+80H_{-1,0}-64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}-664H_{0,0}-224\zeta_{2}H_{0,1}-404H_{0,1}-68H_{1,0}-52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-256H_{0,-1,0}+80H_{0,0,1}+112H_{0,1,0}+32H_{0,1,1}-568H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-128H_{0,-1,-1,0}+64H_{0,-1,0,0}+128H_{0,0,-1,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}-664\right)z+\frac{32}{z}H_{-1,0}+252H_{0}-580H_{1}+128H_{-1}\zeta_{2}+64H_{0}\zeta_{2}-528H_{1}\zeta_{2}+152\zeta_{2}-288H_{0}\zeta_{3}-272\zeta_{3}-460\zeta_{4}+80H_{-1,0}+64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}+16H_{0,0}-224\zeta_{2}H_{0,1}-248H_{0,1}+68H_{1,0}+52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-64H_{0,0,0}-160H_{0,0,1}-192H_{0,1,0}-48H_{0,1,1}+568H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+128H_{0,-1,-1,0}-64H_{0,-1,0,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+664\bigg]\,$, (147)
+$\Delta P_{gg}^{T,\text{s}(2)}=C_{A}^{3}\bigg[\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)z^{2}+\left(336H_{-1,-1,0}+368H_{-1,0,1}\right)z+\frac{1}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{1072}{9}\right)\zeta_{2}+\left(-\frac{8}{9}\right)H_{1}+\frac{88}{3}\zeta_{3}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}+88\zeta_{4}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-208H_{0,0,1}-144H_{0,1,0}-256H_{0,0,1,1}-608H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}+\frac{490}{3}\right)+\frac{1}{z}\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{48478}{27}\right)H_{0}+\left(-\frac{2968}{3}\right)\zeta_{3}+\left(-\frac{320}{3}\right)H_{0,1}+\left(-\frac{284}{3}\right)H_{0,0}+\left(-\frac{592}{9}\right)\zeta_{2}+\frac{320}{3}H_{0,-1,0}+\frac{280\zeta_{2}}{3}H_{0}+472\zeta_{4}-96\zeta_{2}H_{0,0}+160\zeta_{2}H_{0,1}-112H_{0,0,0}+64H_{0,-1,-1,0}-224H_{0,-1,0,0}-128H_{0,-1,0,1}-64H_{0,0,-1,0}+128H_{0,0,0,0}-160H_{0,0,0,1}-64H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)+\frac{z}{z-1}\left(\left(-\frac{1232}{3}\right)H_{1,0}+\frac{16}{9}H_{1}-336H_{1}\zeta_{2}-288H_{1}\zeta_{3}+416\zeta_{2}H_{0,-1}+96\zeta_{2}H_{1,0}+584H_{0,0,1}+320H_{0,1,0}+568H_{1,0,0}+384H_{0,0,1,1}+672H_{0,1,0,0}+384H_{0,1,0,1}+384H_{0,1,1,0}-192H_{1,0,-1,0}+960H_{1,0,0,0}+384H_{1,0,0,1}+384H_{1,0,1,0}+384H_{1,1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3232}{3}\right)H_{0,-1,0}+\left(-\frac{9136}{9}\right)\zeta_{2}+\left(-\frac{592}{9}\right)H_{0,1}+\frac{16964}{9}H_{0,0}+\frac{6736}{3}H_{0,0,0}+\frac{8990}{3}H_{0}+\left(-\frac{2024\zeta_{2}}{3}\right)H_{0}-800H_{0}\zeta_{3}-920\zeta_{3}-576\zeta_{4}-608\zeta_{2}H_{0,0}-32\zeta_{2}H_{0,1}+320H_{0,-1,-1,0}-544H_{0,-1,0,0}-704H_{0,0,-1,0}+1984H_{0,0,0,0}+800H_{0,0,0,1}+512H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)$
+$\left.+\frac{z^{4}}{z-1}\bigg(-\frac{176}{3}H_{0,0,0}+\frac{176}{3}H_{0,-1,0}+\frac{440}{3}\zeta_{3}+\frac{352\zeta_{2}}{3}H_{0}+\frac{1}{z+1}\bigg)\right.+\frac{z^{2}}{z-1}\bigg(-\frac{544}{3}H_{1,0,0}-\frac{8}{9}H_{1}+\frac{176}{3}H_{0,1,0}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-112H_{0,0,1}-256H_{0,0,1,1}-288H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}\bigg)+\frac{z^{3}}{z-1}\bigg(-\frac{176}{3}H_{0,0,1}+\frac{88\zeta_{2}}{3}H_{1}\bigg)+\frac{z^{3}}{z-1}\bigg(\frac{1060}{9}H_{0,0}+\frac{3104}{9}H_{0,1}+\frac{2792}{3}\zeta_{3}+\frac{13402}{9}H_{0}+\frac{160\zeta_{2}}{3}H_{0}+304\zeta_{2}-680\zeta_{4}+96\zeta_{2}H_{0,0}-256\zeta_{2}H_{0,1}-48H_{0,-1,0}+464H_{0,0,0}+128H_{0,-1,0,0}+256H_{0,-1,0,1}-64H_{0,0,-1,0}-256H_{0,0,0,0}-96H_{0,0,0,1}-128H_{0,0,1,0}+\frac{1}{z+1}-\frac{47572}{9}\bigg)+\frac{1}{z+1}\bigg(-\frac{6032}{9}H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\bigg)+\frac{1}{(z-1)(z+1)}\bigg(-\frac{89182}{27}H_{0}-\frac{14612}{9}H_{0,0}-\frac{4448}{3}H_{0,0,0}+\frac{2144}{3}\zeta_{3}+\frac{2528}{3}H_{0,-1,0}+\frac{9136}{9}\zeta_{2}+528H_{0}\zeta_{2}+1088H_{0}\zeta_{3}+808\zeta_{4}+864\zeta_{2}H_{0,0}+304H_{0,1}-512H_{0,-1,-1,0}+768H_{0,-1,0,0}+960H_{0,0,-1,0}-2496H_{0,0,0,0}-1184H_{0,0,0,1}-768H_{0,0,1,0}-\frac{47572}{9}\bigg)+\frac{1}{z(z+1)}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{z}{z+1}\bigg(-\frac{4736}{3}H_{-1,0}-400H_{-1}\zeta_{2}-576H_{-1}\zeta_{3}+768\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}-1312H_{-1,0,0}+384H_{-1,-1,0,0}-768H_{-1,-1,0,1}+384H_{-1,0,-1,0}-960H_{-1,0,0,0}-192H_{-1,0,1,0}\bigg)+\frac{z^{2}}{z+1}\bigg(-\frac{6032}{9}H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\bigg)+\frac{z^{3}}{z+1}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{88\zeta_{2}}{3(z-1)z}H_{1}+336H_{-1,-1,0}+368H_{-1,0,1}\bigg]+N_{f}C_{A}^{2}\bigg[\left(-\frac{32}{3}H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\frac{16\zeta_{2}}{3}H_{1}-16H_{-1}\zeta_{2}\right)z^{2}+32H_{-1,0,1}+96H_{0,-1,-1,0}+\left(-\frac{2174}{9}H_{1}+56H_{-1}\zeta_{2}+88H_{1}\zeta_{2}-48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+32H_{-1,0,1}-96H_{0,-1,-1,0}+48H_{0,-1,0,0}+80H_{0,0,-1,0}+24H_{0,0,0,1}+24H_{0,1,0,0}\right)z+\frac{2174}{9}H_{1}-48H_{0,-1,0,0}-16H_{0,0,-1,0}+\frac{1}{z-1}\bigg(-\frac{1844}{9}H_{0,1}-\frac{356}{3}H_{1,0,0}-\frac{928}{9}H_{1,0}-\frac{112}{3}\zeta_{3}+\frac{160}{9}\zeta_{2}+48\zeta_{2}H_{0,1}-124H_{0,0,1}-96H_{0,1,0}-\frac{836}{27}\bigg)+16H_{0,0,0,0}+24H_{0,0,0,1}+24H_{0,1,0,0}+\frac{1}{z}\bigg(-\frac{32}{3}H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\left(-\frac{16\zeta_{2}}{3}\right)H_{1}-16H_{-1}\zeta_{2}\bigg)+\frac{z}{z-1}\bigg(-\frac{1396}{9}H_{0,0}-\frac{368}{3}\zeta_{3}-\frac{1138}{27}H_{0}+\frac{376}{3}H_{0,-1,0}+\frac{508}{3}\zeta_{2}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-138\zeta_{4}+24\zeta_{2}H_{0,0}-32H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\bigg)+56H_{-1}\zeta_{2}-88H_{1}\zeta_{2}+48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+\frac{z}{z-1}\left(\frac{512}{3}H_{0,1}+\frac{512}{3}H_{1,0}+200H_{0,0,1}+128H_{0,1,0}+200H_{1,0,0}\right)$
+$\frac{z^{2}}{z-1}\left(-\frac{356}{3}H_{1,0,0}-\frac{928}{9}H_{1,0}+\left(-\frac{308}{3}\right)H_{0,0,1}+\left(-\frac{224}{3}\right)H_{0,1,0}+\left(-\frac{4}{3}\right)H_{0,1}-48\zeta_{2}H_{0,1}\right)+\frac{z^{2}}{z-1}\left(-\frac{832}{3}\zeta_{3}-\frac{2140}{9}\zeta_{2}+\frac{1880}{3}H_{0,0}+\frac{6308}{9}H_{0}-116H_{0}\zeta_{2}-96H_{0}\zeta_{3}+78\zeta_{4}-40\zeta_{2}H_{0,0}-104H_{0,-1,0}+392H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+\left(-\frac{32z^{3}}{3(z-1)}\right)H_{0,0,1}+\frac{z^{3}}{z-1}\left(\left(-\frac{1844}{9}\right)\zeta_{2}+\frac{1246}{9}H_{0}+\frac{496}{3}\zeta_{3}+\frac{1780}{9}H_{0,0}+\frac{20\zeta_{2}}{3}H_{0}+138\zeta_{4}-24\zeta_{2}H_{0,0}-136H_{0,-1,0}-32H_{0,0,0}+\frac{1}{z+1}-\frac{22904}{27}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{32}{3}\right)H_{0,0,0}+\frac{32}{3}H_{0,-1,0}+\frac{80}{3}\zeta_{3}+\frac{64\zeta_{2}}{3}H_{0}+16\zeta_{2}-16H_{0,0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{16324}{27}\right)H_{0}+\left(-\frac{5432}{9}\right)H_{0,0}+\left(-\frac{1528}{3}\right)H_{0,0,0}+\frac{376}{3}H_{0,-1,0}+\frac{1996}{9}\zeta_{2}+\frac{880}{3}\zeta_{3}+100H_{0}\zeta_{2}+96H_{0}\zeta_{3}-78\zeta_{4}+40\zeta_{2}H_{0,0}-\frac{22904}{27}\right)+\frac{1}{z(z+1)}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)+\frac{z}{z+1}\left(\left(-\frac{1232}{3}\right)H_{-1,0}-272H_{-1,0,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{1864}{9}\right)H_{-1,0}+\left(-\frac{440}{3}\right)H_{-1,0,0}\right)+\frac{z^{3}}{z+1}\left(\frac{32}{3}H_{-1,0,0}+16H_{-1,0}\right)\bigg]+C_{A}N_{f}^{2}\bigg[-\frac{82}{9}H_{1}-\frac{32}{9}H_{0,1}+z\left(-\frac{32}{9}H_{0,1}+\frac{82}{9}H_{1}\right)+\frac{1}{z-1}\left(\left(-\frac{176}{9}\right)H_{0,0}+\left(-\frac{496}{27}\right)H_{0}+\left(-\frac{32}{9}\right)\zeta_{2}+\frac{230}{27}\right)+\frac{z}{z-1}\left(\frac{64}{3}H_{0,0}+30H_{0}-\frac{148}{9}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{158}{9}\right)H_{0}+\left(-\frac{80}{9}\right)H_{0,0}+\frac{32}{9}\zeta_{2}+\frac{230}{27}\right)-\frac{16}{27(z-1)}\bigg]+N_{f}C_{F}C_{A}\bigg[\left(\left(-\frac{64}{3}\right)H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\left(-\frac{64}{3}\right)H_{0,-1,0}+\frac{64}{3}H_{-1,-1,0}+\frac{64}{3}H_{0,0,0}+\frac{64}{3}H_{0,0,1}+\left(-\frac{32\zeta_{2}}{3}\right)H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}+48H_{0,0}\right)z^{2}+\left(\left(-\frac{1756}{9}\right)H_{1}+\left(-\frac{488}{3}\right)H_{0,1,0}+\left(-\frac{296}{3}\right)H_{0,0,0}+\left(-\frac{152}{3}\right)H_{0,1,1}+\frac{260}{3}H_{1,1}+\frac{2932}{9}H_{0,1}+\frac{1004}{3}H_{1,0}+\frac{1352}{3}H_{-1,0}+\frac{2168}{3}H_{0,0}-176H_{-1}\zeta_{2}-640H_{1}\zeta_{2}+128\zeta_{2}H_{0,-1}-16\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+272H_{0,-1,0}+8H_{0,0,1}+712H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+256H_{0,-1,-1,0}-192H_{0,-1,0,0}-256H_{0,0,-1,0}+128H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}\right)z+\left(-\frac{3632}{3}\right)H_{0,0}+\left(-\frac{3056}{3}\right)H_{0,0,0}+\left(-\frac{1004}{3}\right)H_{1,0}+\left(-\frac{260}{3}\right)H_{1,1}+\left(-\frac{128}{3}\right)H_{0,1,0}+\frac{40}{9}H_{0,1}+\frac{88}{3}H_{0,1,1}+\frac{1756}{9}H_{1}+\frac{1352}{3}H_{-1,0}+\frac{1}{z-1}\left(\left(-\frac{1936}{3}\right)\zeta_{3}+\left(-\frac{2672}{9}\right)\zeta_{2}+\frac{51500}{27}H_{0}+\left(-\frac{416\zeta_{2}}{3}\right)H_{0}-512H_{0}\zeta_{3}-468\zeta_{4}+\frac{88478}{27}\right)+\frac{1}{z}\left(-\frac{64}{3}H_{-1,0,0}+\left(-\frac{64}{3}\right)H_{-1,0,1}+\frac{64}{3}H_{-1,-1,0}+\frac{32\zeta_{2}}{3}H_{1}+32H_{-1}\zeta_{2}-48H_{-1,0}\right)+\frac{z}{z-1}\left(\left(-\frac{3148}{9}\right)H_{0}+\frac{1420}{3}\zeta_{2}+168H_{0}\zeta_{2}+352H_{0}\zeta_{3}+1240\zeta_{3}+160\zeta_{4}-\frac{177946}{27}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{41624}{27}\right)H_{0}+\left(-\frac{1624}{3}\right)\zeta_{3}+\left(-\frac{1156}{9}\right)\zeta_{2}+\frac{40\zeta_{2}}{3}H_{0}+160H_{0}\zeta_{3}+308\zeta_{4}+\frac{89468}{27}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{160}{3}\right)\zeta_{3}+\left(-\frac{128\zeta_{2}}{3}\right)H_{0}-48\zeta_{2}\right)\bigg]$ (137)
+$176H_{-1}\zeta_{2}+640H_{1}\zeta_{2}-128\zeta_{2}H_{0,-1}+112\zeta_{2}H_{0,0}+288\zeta_{2}H_{0,1}-480H_{-1,-1,0}+368H_{-1,0,0}-64H_{-1,0,1}+368H_{0,-1,0}-32H_{0,0,1}-712H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-256H_{0,-1,-1,0}+192H_{0,-1,0,0}+128H_{0,0,-1,0}-448H_{0,0,0,0}-112H_{0,0,0,1}+80H_{0,0,1,0}+48H_{0,0,1,1}-304H_{0,1,0,0}-96H_{0,1,0,1}-64H_{0,1,1,0}-32H_{0,1,1,1}+\delta(1-z)\bigg[\left(\frac{8}{3}\zeta_{2}+\frac{55}{3}\zeta_{4}+\frac{536}{3}\zeta_{3}-16\zeta_{2}\zeta_{3}-80\zeta_{5}+\frac{79}{2}\right)C_{A}^{3}+\left(\left(-\frac{80}{3}\right)\zeta_{3}+\left(-\frac{10}{3}\right)\zeta_{4}+\left(-\frac{8}{3}\right)\zeta_{2}-\frac{233}{18}\right)N_{f}C_{A}^{2}+\left(\frac{29N_{f}^{2}}{18}-\frac{241C_{F}N_{f}}{18}\right)C_{A}+\frac{11}{9}C_{F}N_{f}^{2}+C_{F}^{2}N_{f}\bigg]+\left[\frac{1}{1-z}\right]_{+}\bigg[\left(\left(-\frac{1072}{9}\right)\zeta_{2}+\frac{88}{3}\zeta_{3}+88\zeta_{4}+\frac{490}{3}\right)C_{A}^{3}+\left(\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}-\frac{836}{27}\right)N_{f}C_{A}^{2}+\left(C_{F}N_{f}\left(32\zeta_{3}-\frac{110}{3}\right)-\frac{16}{27}N_{f}^{2}\right)C_{A}\bigg]+C_{F}N_{f}^{2}\bigg[\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{40}{3}\right)H_{1,0}+\left(-\frac{40}{3}\right)H_{1,1}+\left(-\frac{328}{27}\right)H_{0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{136}{9}\zeta_{2}+\frac{224}{9}H_{1}+\frac{344}{9}H_{0,1}+\frac{16\zeta_{2}}{3}H_{0}-32H_{0,0}+16H_{0,0,1}+z\left(\left(-\frac{488}{9}\right)\zeta_{2}+\left(-\frac{224}{9}\right)H_{1}+\left(-\frac{64}{3}\right)\zeta_{3}+\left(-\frac{64}{3}\right)H_{0,0}+\left(-\frac{64}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{0,1,0}+\left(-\frac{16}{3}\right)H_{0,1,1}+\frac{8}{9}H_{0,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,1}+\frac{776}{27}H_{0}+\frac{16\zeta_{2}}{3}H_{0}+16H_{0,0,1}-\frac{992}{27}\right)+\frac{992}{27}\bigg]+C_{F}^{2}N_{f}\bigg[\left(32\zeta_{2}+32H_{-1,0}-32H_{0,0}\right)z^{2}+\left(-48\zeta_{2}H_{0}-160\zeta_{3}H_{0}+1032H_{0}+580H_{1}+128H_{-1}\zeta_{2}+528H_{1}\zeta_{2}+580\zeta_{2}-48\zeta_{3}-380\zeta_{4}+80H_{-1,0}-64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}-664H_{0,0}-224\zeta_{2}H_{0,1}-404H_{0,1}-68H_{1,0}-52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-256H_{0,-1,0}+80H_{0,0,1}+112H_{0,1,0}+32H_{0,1,1}-568H_{1,0,0}-240H_{1,0,1}-160H_{1,1,0}-80H_{1,1,1}-128H_{0,-1,-1,0}+64H_{0,-1,0,0}+128H_{0,0,-1,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}-664\right)z+\frac{32}{z}H_{-1,0}+252H_{0}-580H_{1}+128H_{-1}\zeta_{2}+64H_{0}\zeta_{2}-528H_{1}\zeta_{2}+152\zeta_{2}-288H_{0}\zeta_{3}-272\zeta_{3}-460\zeta_{4}+80H_{-1,0}+64\zeta_{2}H_{0,-1}+16\zeta_{2}H_{0,0}+16H_{0,0}-224\zeta_{2}H_{0,1}-248H_{0,1}+68H_{1,0}+52H_{1,1}+256H_{-1,-1,0}-128H_{-1,0,0}-64H_{0,0,0}-160H_{0,0,1}-192H_{0,1,0}-48H_{0,1,1}+568H_{1,0,0}+240H_{1,0,1}+160H_{1,1,0}+80H_{1,1,1}+128H_{0,-1,-1,0}-64H_{0,-1,0,0}-32H_{0,0,0,0}-16H_{0,0,0,1}-144H_{0,0,1,0}-48H_{0,0,1,1}+240H_{0,1,0,0}+96H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+664\bigg]\,.$ (138)
 4.5 N3LO coefficient functions and small-x expansion
@@
 Below we collect the small-x data for the physical coefficient functions, we use superscripts ‘s’ to indicate the corresponding results are in the singlet sector
-$\Delta\mathscr{I}^{\text{s}(1)}_{qq}(x)=$ $2C_{F}\,$, (150)
-$\Delta\mathscr{I}^{\text{s}(2)}_{qq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{2}{3}\right)\ln^{3}x+\left(-\frac{11}{6}\right)\ln^{2}x+\left(-\frac{31}{9}\right)\ln x-12\zeta_{2}+4\zeta_{3}+\frac{397}{27}\right)$
+$\Delta\mathscr{I}^{\text{s}(1)}_{qq}(x)\simeq$ $2C_{F}\,$, (141)
+$\Delta\mathscr{I}^{\text{s}(2)}_{qq}(x)\simeq$ $C_{A}C_{F}\left(\left(-\frac{2}{3}\right)\ln^{3}x+\left(-\frac{11}{6}\right)\ln^{2}x+\left(-\frac{31}{9}\right)\ln x-12\zeta_{2}+4\zeta_{3}+\frac{397}{27}\right)$
 $+$ $C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+\frac{32}{3}\ln^{2}x+\frac{344}{9}\ln x-16\zeta_{2}+\frac{2452}{27}\right)$
-$+$ $C_{F}^{2}\left(\ln^{3}x+3\ln^{2}x+4\ln x+18\zeta_{2}+20\zeta_{3}-52\right)\,$, (151)
-$\Delta\mathscr{I}^{\text{s}(1)}_{qg}(x)=$ $8N_{f}T_{F}\,$, (152)
-$\Delta\mathscr{I}^{\text{s}(2)}_{qg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+22\ln^{2}x+72\ln x-32\zeta_{2}+8\zeta_{3}+140\right)$
+$+$ $C_{F}^{2}\left(\ln^{3}x+3\ln^{2}x+4\ln x+18\zeta_{2}+20\zeta_{3}-52\right)\,.$ (142)
+$\Delta\mathscr{I}^{\text{s}(1)}_{qg}(x)\simeq$ $8N_{f}T_{F}\,$, (143)
+$\Delta\mathscr{I}^{\text{s}(2)}_{qg}(x)\simeq$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+22\ln^{2}x+72\ln x-32\zeta_{2}+8\zeta_{3}+140\right)$
 $+$ $C_{F}N_{f}T_{F}\left(\frac{2}{3}\ln^{3}x+13\ln^{2}x+56\ln x-24\zeta_{2}-56\zeta_{3}+142\right)\,$, (144)
-$\Delta\mathscr{I}^{\text{s}(3)}_{qg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(\frac{32308}{81}-12\zeta_{2}\right)\ln^{3}x+\left(\left(-\frac{424}{3}\right)\zeta_{2}+\left(-\frac{328}{3}\right)\zeta_{3}+\frac{19276}{9}\right)\ln^{2}x$
+$\Delta\mathscr{I}^{\text{s}(3)}_{qg}(x)\simeq$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(\frac{32308}{81}-12\zeta_{2}\right)\ln^{3}x+\left(\left(-\frac{424}{3}\right)\zeta_{2}+\left(-\frac{328}{3}\right)\zeta_{3}+\frac{19276}{9}\right)\ln^{2}x$
 $+$ $\left(\left(-\frac{38}{3}\right)\zeta_{4}+\frac{376}{3}\zeta_{2}-1024\zeta_{3}+\frac{1376872}{243}\right)\ln x+2\ln^{5}x+\frac{410}{9}\ln^{4}x+\frac{11456}{9}\zeta_{4}$
@@
 $+$ $C_{F}N_{f}^{2}T_{F}^{2}\bigg[\left(32\zeta_{2}-\frac{127328}{81}\right)\ln^{2}x+\left(128\zeta_{2}+64\zeta_{3}-\frac{1108564}{243}\right)\ln x+\left(-\frac{16}{15}\right)\ln^{5}x$
-$+$ $\left(-\frac{764}{27}\right)\ln^{4}x+\left(-\frac{23300}{81}\right)\ln^{3}x+\left(-\frac{848}{9}\right)\zeta_{4}+\frac{2416}{81}\zeta_{2}+\frac{5984}{27}\zeta_{3}-\frac{423361}{81}\bigg]\,$, (154)
-$\Delta\mathscr{I}^{\text{s}(1)}_{gq}(x)=$ $-4C_{F}\,$, (155)
-$\Delta\mathscr{I}^{\text{s}(2)}_{gq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{4}{3}\right)\ln^{3}x-16\ln^{2}x-50\ln x+28\zeta_{2}+48\zeta_{3}-\frac{4454}{27}\right)$
+$+$ $\left(-\frac{764}{27}\right)\ln^{4}x+\left(-\frac{23300}{81}\right)\ln^{3}x+\left(-\frac{848}{9}\right)\zeta_{4}+\frac{2416}{81}\zeta_{2}+\frac{5984}{27}\zeta_{3}-\frac{423361}{81}\bigg]\,.$ (145)
+$\Delta\mathscr{I}^{\text{s}(1)}_{gq}(x)\simeq$ $-4C_{F}\,$, (146)
+$\Delta\mathscr{I}^{\text{s}(2)}_{gq}(x)\simeq$ $C_{A}C_{F}\left(\left(-\frac{4}{3}\right)\ln^{3}x-16\ln^{2}x-50\ln x+28\zeta_{2}+48\zeta_{3}-\frac{4454}{27}\right)$
 $+$ $\frac{544}{27}C_{F}N_{f}T_{F}+C_{F}^{2}\left(\left(-\frac{2}{3}\right)\ln^{3}x-10\ln^{2}x-19\ln x+23\right)\,$, (147)
-$\Delta\mathscr{I}^{\text{s}(3)}_{gq}(x)=$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\frac{64}{9}\zeta_{2}+\frac{19232}{81}\right)\ln^{2}x+\left(\frac{3904}{27}\zeta_{2}+\frac{1472}{9}\zeta_{3}+\frac{110612}{243}\right)\ln x$
+$\Delta\mathscr{I}^{\text{s}(3)}_{gq}(x)\simeq$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\frac{64}{9}\zeta_{2}+\frac{19232}{81}\right)\ln^{2}x+\left(\frac{3904}{27}\zeta_{2}+\frac{1472}{9}\zeta_{3}+\frac{110612}{243}\right)\ln x$
 $+$ $\frac{32}{27}\ln^{4}x+\frac{728}{27}\ln^{3}x+\left(-\frac{304}{9}\right)\zeta_{4}+\left(-\frac{3448}{81}\right)\zeta_{2}+\frac{9296}{27}\zeta_{3}+\frac{219356}{243}\bigg]$
@@
 $+$ $\left(\frac{220}{3}\zeta_{2}+\frac{1232}{3}\zeta_{3}-24\zeta_{4}+\frac{328}{3}\right)\ln x+\frac{16}{45}\ln^{5}x+\frac{28}{3}\ln^{4}x+\frac{1184}{3}\zeta_{4}+\frac{2516}{3}\zeta_{3}$
-$+$ $\zeta_{2}\left(144\zeta_{3}+284\right)-640\zeta_{5}-\frac{12365}{12}\bigg]\,$, (157)
-$\Delta\mathscr{I}^{\text{s}(1)}_{gg}(x)=$ $-8C_{A}\,$, (158)
-$\Delta\mathscr{I}^{\text{s}(2)}_{gg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{2}x+\frac{136}{9}\ln x+\frac{1882}{27}\right)+C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+18\ln^{2}x+84\ln x+120\right)$
+$+$ $\zeta_{2}\left(144\zeta_{3}+284\right)-640\zeta_{5}-\frac{12365}{12}\bigg]\,.$ (148)
+$\Delta\mathscr{I}^{\text{s}(1)}_{gg}(x)\simeq$ $-8C_{A}\,$, (149)
+$\Delta\mathscr{I}^{\text{s}(2)}_{gg}(x)\simeq$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{2}x+\frac{136}{9}\ln x+\frac{1882}{27}\right)+C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+18\ln^{2}x+84\ln x+120\right)$
 $+$ $C_{A}^{2}\left(\left(-\frac{8}{3}\right)\ln^{3}x+\left(-\frac{101}{3}\right)\ln^{2}x+\left(-\frac{1109}{9}\right)\ln x+56\zeta_{2}+96\zeta_{3}-\frac{9020}{27}\right)\,$, (150)
-$\Delta\mathscr{I}^{\text{s}(3)}_{gg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(8\zeta_{2}+\frac{3416}{81}\right)\ln^{3}x+\left(\frac{488}{9}\zeta_{2}+112\zeta_{3}+\frac{10708}{27}\right)\ln^{2}x$
+$\Delta\mathscr{I}^{\text{s}(3)}_{gg}(x)\simeq$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(8\zeta_{2}+\frac{3416}{81}\right)\ln^{3}x+\left(\frac{488}{9}\zeta_{2}+112\zeta_{3}+\frac{10708}{27}\right)\ln^{2}x$
 $+$ $\left(\frac{2992}{27}\zeta_{2}+\frac{452}{3}\zeta_{4}+\frac{6160}{9}\zeta_{3}+\frac{93668}{81}\right)\ln x+\left(-\frac{4}{15}\right)\ln^{5}x+\left(-\frac{34}{27}\right)\ln^{4}x$